### v3

# Robust Blockwise Random Pivoting:

Abstract
The interpolative decomposition (ID) aims to construct a low-rank approximation formed by a basis consisting of row/column skeletons in the original matrix and a corresponding interpolation matrix. This work explores fast and accurate ID algorithms from five essential perspectives for empirical performance: (a) skeleton complexity that measures the minimum possible ID rank for a given low-rank approximation error, (b) asymptotic complexity in FLOPs, (c) parallelizability of the computational bottleneck as matrix-matrix multiplications, (d) error-revealing property that enables automatic rank detection for given error tolerances without prior knowledge of target ranks, (e) ID-revealing property that ensures efficient construction of the optimal interpolation matrix after selecting the skeletons. While a broad spectrum of algorithms have been developed to optimize parts of the aforementioned perspectives, practical ID algorithms proficient in all perspectives remain absent. To fill in the gap, we introduce robust blockwise random pivoting (RBRP) that is parallelizable, error-revealing, and exactly ID-revealing, with comparable skeleton and asymptotic complexities to the best existing ID algorithms in practice. Through extensive numerical experiments on various synthetic and natural datasets, we demonstrate the appealing empirical performance of RBRP from the five perspectives above, as well as the robustness of RBRP to adversarial inputs.
keywords:
Randomized Numerical Linear Algebra, Interpolative Decomposition, Column Subset Selection, Adaptive Sampling, Parallelization
{MSCcodes}
65F55, 65Y05, 68T05
1 Introduction
The interpolative decomposition (ID) is a special type of low-rank approximation where a basis for the row/column space is formed by subselecting rows/columns from the target matrix. This allows the ID to capture the essential structures and relationships within the data. ID has found applications in various fields, including numerical analysis, data compression, and machine learning, where preserving the integrity of the original data is of paramount importance.
Given a data matrix $\mathbf{X}=\left[\mathbf{x}_{1},\cdots,\mathbf{x}_{n}\right]^{T}\in\mathbb{R}^{n\times d}$ consisting of $n$ data points $\left\{\mathbf{x}_{i}\in\mathbb{R}^{d}\right\}_{i\in[n]}$, along with a constant $\epsilon>0$ and a target rank $r\in\mathbb{N}$, we aim to construct an $(r,\epsilon)$-ID of $\mathbf{X}$:
(1) $\underset{n\times d}{\mathbf{X}}\approx\underset{n\times k}{\mathbf{W}}\ \underset{k\times d}{\mathbf{X}_{S}}\quad\emph{s.t.}\quad\left\|\mathbf{X}-\mathbf{W}\mathbf{X}_{S}\right\|_{F}^{2}\leq(1+\epsilon)\left\|\mathbf{X}-{\mathbf{X}}_{\langle r\rangle}\right\|_{F}^{2}$
where (i) $k$is the (unknown) target rank; (ii) ${\mathbf{X}}_{\langle r\rangle}$is the optimal rank-$r$ approximation of $\mathbf{X}$ (given by the singular value decomposition (SVD)); (iii) $S=\left\{s_{1},\cdots,s_{k}\right\}\subset[n]$are the skeleton indices corresponding to the row skeleton submatrix $\mathbf{X}_{S}=[\mathbf{x}_{s_{1}},\cdots,\mathbf{x}_{s_{k}}]^{\top}\in\mathbb{R}^{k\times d}$; and (iv) $\mathbf{W}\in\mathbb{R}^{n\times k}$is an interpolation matrix with (nearly) optimal error $\left\|\mathbf{X}-\mathbf{W}\mathbf{X}_{S}\right\|_{F}^{2}$ for the given $\mathbf{X}_{S}$. The construction of an ID can be divided into two stages:
  1. (i)
    Select a skeleton subset $S\subset[n]$ that achieves a small skeletonization error
    (2) $\mathcal{E}_{\mathbf{X}}\left(S\right)\triangleq\left\|\mathbf{X}-\mathbf{X}\mathbf{X}_{S}^{\dagger}\mathbf{X}_{S}\right\|_{F}^{2}=\min_{\mathbf{W}\in\mathbb{R}^{n\times\left|S\right|}}\left\|\mathbf{X}-\mathbf{W}\mathbf{X}_{S}\right\|_{F}^{2}.$
  2. (ii)
    Given $S$, compute an interpolation matrix $\mathbf{W}\in\mathbb{R}^{n\times\left|S\right|}$ efficiently111 It is worth clarifying that although Equation 2 directly provides $\mathbf{W}=\mathbf{X}\mathbf{X}_{S}^{\dagger}$ as the minimizer of $\min_{\mathbf{W}}\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)=\mathcal{E}_{\mathbf{X}}\left(S\right)$, the explicit computation of $\mathbf{X}\mathbf{X}_{S}^{\dagger}$ takes $O\left(ndk+dk^{2}+nk^{2}\right)$ time (e.g. via Equation 8), where the dominant cost $O(ndk)$ is undesired. Therefore, whenever available, information from the skeleton selection process is usually leveraged to evaluate/approximate $\mathbf{W}$ efficiently in $o(ndk)$ time (as elaborated in Section 5), which may lead to the gap $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)-\mathcal{E}_{\mathbf{X}}\left(S\right)>0$. that (approximately) minimizes the interpolation error: $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)\triangleq\left\|\mathbf{X}-\mathbf{W}\mathbf{X}_{S}\right\|_{F}^{2}$.
In this work, we explore fast and accurate ID algorithms from the perspectives of both skeleton subset selection and interpolation matrix construction, and we start by posing the main question:
How do we construct an $(r,\epsilon)$-ID efficiently with as small a skeleton $S$ as possible,
without prior knowledge of the target rank $k$?
1.1 What are Fast and Accurate ID Algorithms?
To formalize the concept of “fast and accurate” ID algorithms, we dissect the main question above into the following properties, together providing systematic performance measurements for ID algorithms from both the efficiency and accuracy perspectives.
  1. (a)
    Skeleton complexity measures the minimum size of the skeleton set (i.e. the minimum possible rank of ID) that an algorithm must select before $\mathcal{E}_{\mathbf{X}}\left(S\right)\leq(1+\epsilon)\left\|\mathbf{X}-{\mathbf{X}}_{\langle r\rangle}\right\|_{F}^{2}$ (i.e. where there exists an $(r,\epsilon)$-ID associated with $S$).
  2. (b)
    Asymptotic complexity measures the number of floating point operations (FLOPs) in the skeleton selection process.
  3. (c)
    Parallelizability (in the context of this work) refers to the implementation advantage that the dominant cost for skeleton selection in the algorithm appears as matrix-matrix, instead of matrix-vector, multiplications. On modern processors, Level-3 BLAS, or matrix-matrix, operations can leverage memory cashing much more efficiently than Level-2 or Level-1 BLAS operations [3, 23]. As a result, a sequence of matrix operations implemented using Level-3 BLAS can attain significantly higher performance than an optimal implementation using Level-2 or Level-1 BLAS, especially on parallel processors.
  4. (d)
    Error-revealing property refers to the ability of an algorithm to evaluate the skeletonization error efficiently on the fly so that the target rank $k$ does not need to be given a priori.
    Definition 1.1 (Error-revealing).
    An ID algorithm is error-revealing if after selecting a skeleton subset $S$, it can evaluate $\mathcal{E}_{\mathbf{X}}\left(S\right)$ efficiently with at most $O\left(n\right)$ operations.
    That is, given any relative error tolerance $\tau\in(0,1)$, an error-revealing ID algorithm can determine an appropriate rank $k$ with negligible additional cost (i.e. no more than $O(nk)$) while selecting a skeleton subset $S$ ($\left|S\right|=k$) such that $\mathcal{E}_{\mathbf{X}}\left(S\right)\leq\tau\left\|\mathbf{X}\right\|_{F}^{2}$. For the sake of analysis, we define a useful constant associated with $\mathbf{X}$ throughout the paper:
    (3) $\eta_{r}\triangleq{\left\|\mathbf{X}-{\mathbf{X}}_{\langle r\rangle}\right\|_{F}^{2}}/{\left\|\mathbf{X}\right\|_{F}^{2}},$
    which quantifies the relative tail weight of $\mathbf{X}$ with respect to the rank $r$ and represents the relative optimal rank-$r$ approximation error of $\mathbf{X}$. Notice that when $\tau<(1+\epsilon)\eta_{r}$, the algorithm outputs an $(r,\epsilon)$-ID.
  5. (e)
    ID-revealing property characterizes whether the skeleton selection stage of an ID algorithm generates sufficient information for efficient interpolation matrix construction.
    Definition 1.2 (Exactly, approximately, and non-ID-revealing).
    We say a skeleton selection algorithm is exactly ID-revealing if it contains sufficient information in addition to $S$ such that the optimal interpolation matrix $\mathbf{W}=\mathbf{X}\mathbf{X}_{S}^{\dagger}$ can be evaluated efficiently in $O(nk^{2})$ time. Otherwise, if a suboptimal interpolation matrix $\mathbf{W}\approx\mathbf{X}\mathbf{X}_{S}^{\dagger}$ can be constructed in $O(nk^{2})$ time, we say that the skeleton selection algorithm is approximately ID-revealing. If neither $\mathbf{X}\mathbf{X}_{S}^{\dagger}$ nor its approximations can be constructed in $O(nk^{2})$ time, the skeleton selection algorithm is non-ID-revealing.
1.2 How to Combine Adaptiveness and Randomness for ID?
Adaptiveness and randomness are two critical algorithmic properties that can be leveraged to characterize a skeleton selection algorithm. On the one hand, with adaptiveness, the skeleton selection in each step is aware of the selected skeletons from the previous steps so that redundant skeleton selection can be better avoided. On the other hand, randomness diversifies the skeleton selection, thereby improving the robustness of algorithms to scarce adversarial inputs and providing strong statistical guarantees for skeleton complexities.
To ground the notions of adaptiveness and randomness, we synopsize some representative existing skeleton selection algorithms below, as well as in Table 1, with a focus on the aforementioned properties for performance measurement.
  1. (a)
    Greedy pivoting is a classical strategy in numerical linear algebra that involves only adaptiveness. As an example of greedy pivoting, column-pivoted QR (CPQR) [22, Section 5.4.2] picks the point in the residual with the maximum norm in each step and adaptively updates the residual via projection onto the orthogonal complement of the skeletons. Despite the error-revealing and exactly ID-revealing abilities, along with the remarkable empirical success, of CPQR, deterministic greedy pivoting algorithms like CPQR are inherently sequential and vulnerable to adversarial inputs where the skeleton complexity approaches the problem size $n$ [29, 5].
  2. (b)
    Sampling comprises the well-researched body of skeleton selection methods that involve only randomness. Some common examples related to ID include squared-norm sampling [20], leverage score sampling [33], and DPP sampling [2, 31, 12]. As a trade-off between the skeleton complexity and efficiency, the fast ($O(nd)$-time) sampling methods like squared-norm sampling tend to suffer from high skeleton complexities that depend heavily on the matrix [13, 5]; whereas constructions of the more sophisticated distributions like leverage score and DPP are usually expensive [18, 27, 10]. Moreover, for ID, sampling methods generally fail to be error-revealing or ID-revealing.
  3. (c)
    Random pivoting combines adaptiveness and randomness by replacing the greedy selection of maximum norm in CPQR with random sampling according to the squared-norm distribution associated with the residual [13, 14, 5]. Closely related to the ID problem considered in this work, the idea of random pivoting has been revitalized by the inspiring recent work [5] in the context of column Nyström approximation with a nearly optimal skeleton complexity guarantee [5, Theorem 3.1]. However, analogous to CPQR, although random pivoting enjoys the desired error-revealing and exact ID-revealing properties, the sequential nature of random pivoting compromises its efficiency in practice.
  4. (d)
    Sketchy pivoting is an alternative combination of adaptiveness and randomness that randomly embeds the row space of $\mathbf{X}$ via sketching [26, 49], on top of which skeletons are selected via greedy pivoting [48, 16, 39]. In contrast to random pivoting, the cost of sketchy pivoting is dominated by the embarrassingly parallelizable sketching process. As a trade-off for superior empirical efficiency, sketchy pivoting sacrifices the error-revealing property since sketching requires prior knowledge of the target $k$ or its overestimation. Furthermore, sketchy pivoting is approximately ID-revealing due to the loss of information during sketching. As a simple effective remedy for the loss of accuracy, we observe that multiplicative oversampling can remarkably improve the gap between the interpolation and skeletonization error $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)-\mathcal{E}_{\mathbf{X}}\left(S\right)$, which we refer to as oversampled sketchy ID (OSID).
Table 1: Summary of performance for skeleton selection methods that leverage adaptiveness and/or randomness. Recall from Equation 3 that $\eta_{r}\triangleq{\left\|\mathbf{X}-{\mathbf{X}}_{\langle r\rangle}\right\|_{F}^{2}}/{\left\|\mathbf{X}\right\|_{F}^{2}}$. The skeleton complexities in $(\cdot)^{*}$ are conjectured based on extensive empirical evidence and intuitive rationale but without formal proofs. In the “Dominant Cost” column (i.e. showing both the asymptotic complexity and parallelizability), “m-v” stands for matrix-vector multiplications, which are significantly less efficient than “m-m”—matrix-matrix multiplications—due to the lack of parallelizability [23].
Skeleton Selection Skeleton Complexity Dominant Cost Error-revealing ID-revealing Greedy pivoting/CPQR ([22, Section 5.4.2]) $k\geq\left(1-\left(1+\epsilon\right)\eta_{r}\right)n$ $O(ndk)$ m-v Yes Exactly Squared-norm sampling (Proposition 2.3 [20]) $k\geq\frac{r-1}{\epsilon\eta_{r}}+\frac{1}{\epsilon}$ $O(nd)$ m-m No Non Random pivoting (Algorithm 1 [5, 13, 14]) $k\geq k_{\text{RP}}=\frac{r}{\epsilon}+$ $r\min\left\{\log\left(\frac{1}{\epsilon\eta_{r}}\right),1+\log\left(\frac{2^{r}}{\epsilon}\right)\right\}$ $O(ndk)$ m-v Yes Exactly Sketchy pivoting (Algorithm 3 [48, 16]) $\left(k\gtrsim k_{\text{RP}}\right)^{*}$ $O(ndk)$ m-m No Approximately RBRP (this paper) (Algorithm 4) $\left(k\gtrsim k_{\text{RP}}\right)^{*}$ $O(ndk)$ m-m Yes Exactly
To fill in a critical missing piece in the family of existing methods in Table 1 that leverage adaptiveness and randomness, in this work, we introduce robust blockwise random pivoting (RBRP, Algorithm 4)—an ID algorithm that is $O(ndk)$-time, parallelizable, error-revealing, and exactly ID-revealing, with comparable skeleton complexities to the nearly optimal random pivoting [5] in practice.
In particular, the plain blockwise random pivoting (e.g. an extension of the blocked RPCholesky algorithm introduced in [5, Algorithm 2.2]) tends to suffer from unnecessarily large skeleton complexity (cf. Figure 1 (left)) under adversarial inputs (e.g. Example 4.1) due to the lack of local adaptiveness within each block. As a remedy, RBRP leverages robust blockwise filtering (Remark 4.2)—applying CPQR to every small sampled block locally and discarding the potentially redundant points through a truncation on the relative residual of the CPQR. By choosing a reasonable block size, such robust blockwise filtering effectively resolves the inefficiency in skeleton complexity encountered by the plain blockwise random pivoting (cf. Figure 1 (right)), with negligible additional cost.
1.3 Notations and Roadmap
For $\mathbf{X}\in\mathbb{R}^{n\times d}$, let $\mathbf{X}=\mathbf{U}\bm{\Sigma}\mathbf{V}^{\top}=\sum_{i=1}^{\min(n,d)}\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\top}$ be the singular value decomposition of $\mathbf{X}$ with $\sigma_{1}\geq\cdots\geq\sigma_{\min(n,d)}\geq 0$. Given any $r\in[\min(n,d)]$, we denote $\mathbf{U}_{r}=\left[\mathbf{u}_{1},\cdots,\mathbf{u}_{r}\right]$, $\bm{\Sigma}_{r}=\mathop{\mathrm{diag}}\left(\sigma_{1},\cdots,\sigma_{r}\right)$, $\mathbf{V}_{r}=\left[\mathbf{v}_{1},\cdots,\mathbf{v}_{r}\right]$ such that ${\mathbf{X}}_{\langle r\rangle}=\mathbf{U}_{r}\bm{\Sigma}_{r}\mathbf{V}_{r}^{\top}$.
We follow the MATLAB notation for vector and matrix indexing throughout this work. For any $n\in\mathbb{N}$, let $[n]=\left\{1,\cdots,n\right\}$ and $\Delta_{n}=\left\{\mathbf{p}\in[0,1]^{n}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \left\|\mathbf{p}\right\|_{1}=1\right\}$ be the probability simplex of dimension $n$. Also, let $\mathfrak{S}_{n}$ be the set of all permutations of $[n]$. For any distribution $\mathbf{p}\in\Delta_{n}$ and $k\in\mathbb{N}$, $\mathbf{p}^{k}$ represents the joint distribution over $\Delta_{n}^{k}=\Delta_{n}\times\cdots\times\Delta_{n}$ such that $S=\left\{s_{j}\sim\mathbf{p}\right\}_{j\in[k]}\sim\mathbf{p}^{k}$ is a set of $k$ i.i.d. samples. For any $n\in\mathbb{N}$, let $\mathbf{e}_{i}$ ($i\in[n]$) be the $i$-th canonical base of $\mathbb{R}^{n}$. For any $m\in\mathbb{N}$, let $\textbf{1}_{m}\in\mathbb{R}^{m}$ and $\textbf{0}_{m}\in\mathbb{R}^{m}$ be the vector with all entries equal to one and zero, respectively.
As a roadmap, we review the adaptiveness-only and randomness-only skeleton selection algorithms in Section 2, along with the two different combinations of adaptiveness and randomness in Section 3. Then, we introduce the RBRP algorithm formally in Section 4. In Section 5, we discuss the efficient construction of an interpolation matrix for exactly and approximately ID-revealing algorithms. In Section 6, we demonstrate the appealing empirical accuracy and efficiency of RBRP via extensive numerical experiments. We provide the experiment code, as well as a high-performance implementation of RBRP, at https://github.com/dyjdongyijun/Robust_Blockwise_Random_Pivoting.
2 Adaptiveness vs. Randomness
2.1 Adaptiveness via Greedy Squared-norm Pivoting
Greedy pivoting is a classical way of incorporating adaptiveness in skeleton selection [24, 44, 48]. Column pivoted QR (CPQR) [22, Section 5.4.2] is one of the most commonly used greedy pivoting methods. The pivoting strategy and adaptive update are two key components that characterize a greedy pivoting method. In particular, CPQR involves greedy squared-norm pivoting and adaptive QR updates as synopsized below.
Starting with $\mathbf{X}^{(0)}\leftarrow\mathbf{X}$, given a(n) (active) data (sub)matrix $\mathbf{X}^{(t)}\in\mathbb{R}^{n\times d}$ with at most $n-t$ nonzero rows, CPQR on $\mathbf{X}^{\top}$ greedily pivots the row in $\mathbf{X}^{(t)}$ with the maximum $\ell_{2}$-norm $s_{t+1}\leftarrow\mathop{\mathrm{argmax}}_{i\in[n]}\left\|\mathbf{X}^{(t)}\left(i,:\right)\right\|_{2}^{2}$ (i.e. squared-norm pivoting) and adaptively updates the active submatrix by projecting it onto the orthogonal complement of the selected skeleton222 For illustration, we consider the exact arithmetic here and express the QR update as a Gram-Schmidt process, whereas, in practice, QR updates are usually conducted via Householder reflection [22, Section 5.1.2] for numerical stability. :
(4) $\mathbf{X}^{(t+1)}\leftarrow\mathbf{X}^{(t)}-\mathbf{X}^{(t)}\frac{\mathbf{X}^{(t)}(s_{t+1},:)^{\top}\mathbf{X}^{(t)}(s_{t+1},:)}{\left\|\mathbf{X}^{(t)}(s_{t+1},:)\right\|_{2}^{2}}.$
The rank-revealing property of CPQR [24, Theorem 7.2] implies that the first $k$ pivots $S$ consisting of $\left\{s_{1},\cdots,s_{k}\right\}$ form a reasonable skeleton subset with
(5) $\mathcal{E}_{\mathbf{X}}\left(S\right)\leq 4^{k}\left(n-k\right)\left\|\mathbf{X}-{\mathbf{X}}_{\langle k\rangle}\right\|_{F}^{2}.$
The exponential dependence on $k$ in Equation 5 is tight under adversarial inputs (e.g. Kahan matrices [29][24, Example 1]). In the worst case, CPQR can take almost all of the $n$ points as skeletons before finding an $(r,\epsilon)$-ID. Concretely, [5, Theorem 4.1] implies that for CPQR,
(6) $\mathcal{E}_{\mathbf{X}}\left(S\right)\leq\frac{1}{\eta_{r}}\cdot\left(1-\frac{k}{n}\right)\left\|\mathbf{X}-{\mathbf{X}}_{\langle r\rangle}\right\|_{F}^{2}.$
That is, CPQR provides an $(r,\epsilon)$-ID when $k\geq\left(1-(1+\epsilon)\eta_{r}\right)n$. Moreover, for any $S$ with $\left|S\right|=k\leq\min\left(1-(1+\epsilon)\eta_{r},\eta_{r}\right)n-1$, there exists an $\mathbf{X}$ where CPQR fails to form an $(r,\epsilon)$-ID.
Remark 2.1 (Greedy squared-norm pivoting is vulnerable but empirically successful).
The vulnerability to adversaries is a common and critical drawback of squared-norm pivoting in CPQR (and greedy pivoting methods in general). Nevertheless, it has been widely observed and conjectured that such adversarial inputs are extremely scarce in practice [45]. Therefore, CPQR serves as a good skeleton selection method for most data matrices empirically [48, 16], as long as orthonormality is maintained to high precision [6].
Remark 2.2 (CPQR is inherently sequential but error-revealing).
From the computational efficiency perspective, a major drawback of CPQR is that the adaptive updates are inherently sequential (only half of the operations can be cast into high-performance matrix-matrix multiplication in the best-known algorithm that underlies LAPACK’s routine geqp3 [41]). As a result, CPQR tends to be observably slower than simple alternatives with the same asymptotic complexity like Gaussian elimination with partial pivoting [22, Section 3.4]—another common greedy pivoting method that retains parallelizability at a cost of the rank-revealing guarantee.
Nevertheless, the adaptive nature grants CPQR an appealing bonus—the error-revealing ability. It is known that the skeletonization error of CPQR, i.e. $\mathcal{E}_{\mathbf{X}}\left(S\right)=\left\|\mathbf{X}^{\left(\left|S\right|\right)}\right\|_{F}^{2}$, can be downdated efficiently at the end of each update.
2.2 Randomness via Squared-norm Sampling
Sampling is a widely used approach for random skeleton selection (also referred to as column subset selection, e.g. in [11, 9]). To identify a nearly optimal (smallest possible) skeleton subset $S$ for an $(r,\epsilon)$-ID, the intuitive goal is to construct a distribution $\mathbf{p}\in\Delta_{n}$ over the $n$ data points $\left\{\mathbf{x}_{i}\right\}_{i\in[n]}$ that weights each $\mathbf{x}_{i}$ according to its relative “importance.”
Analogous to the squared-norm pivoting in CPQR (Remark 2.1), sampling according to the squared-norm distribution is a natural choice for such “importance” sampling. Concretely, adapting results from [13] and [5, Theorem 4.2] leads to the following expected skeleton complexity for squared-norm sampling.
Proposition 2.3 (Squared-norm sampling).
Consider a size-$k$ skeleton subset $S=\left\{s_{1},\cdots,s_{k}\right\}$ where each $s_{j}\in S$ is sampled from the squared-norm distribution $s_{j}\sim\mathbf{p}=\left(p_{1},\cdots,p_{n}\right)$, $p_{i}=\frac{\left\|\mathbf{x}_{i}\right\|_{2}^{2}}{\left\|\mathbf{X}\right\|_{F}^{2}}$ independently with replacement. Then
$\mathbb{E}_{S\sim\mathbf{p}^{k}}\left[\mathcal{E}_{\mathbf{X}}\left(S\right)\right]\leq\left(1+\frac{r-1}{k\eta_{r}}+\frac{1}{k}\right)\left\|\mathbf{X}-{\mathbf{X}}_{\langle r\rangle}\right\|_{F}^{2}.$
That is, squared-norm sampling yields an $(r,\epsilon)$-ID when $k\geq\frac{r-1}{\epsilon\eta_{r}}+\frac{1}{\epsilon}$.
Compared to the skeleton complexity guarantee of squared-norm pivoting (CPQR) in Equation 6, the squared-norm sampling in Proposition 2.3 provides a potentially weaker guarantee in expectation (instead of worst-case) but yields a better skeleton complexity that scales proportionally to $r$ (instead of $n$).
Despite the improved expected skeleton complexity compared to CPQR, like CPQR (Remark 2.1), squared-norm sampling can also be vulnerable to adversarial inputs by selecting redundant skeletons with higher probabilities. Such potential inefficiency can be reflected by the linear dependence of the skeleton complexity $k\geq\frac{r-1}{\epsilon\eta_{r}}+\frac{1}{\epsilon}$ on $\frac{r}{\eta_{r}}$, as instantiated in Example 2.4.
Example 2.4 (Adversarial input for squared-norm sampling).
For $n\in\mathbb{N}$ divisible by $k\in\mathbb{N}$, $r\in\mathbb{N}$ such that $k=(1+\beta)r<n$ for some $\beta\in\mathbb{N}$, letting $\alpha_{i}=\sqrt{\alpha}\gg\beta\geq 1$ for all $1\leq i\leq r$, and $\alpha_{i}=1$ for all $r+1\leq i\leq k$, we consider the following data matrix
$\mathbf{X}^{\top}=\left[\alpha_{1}\mathbf{e}_{1}\textbf{1}_{n/k}^{\top},\cdots,\alpha_{r}\mathbf{e}_{r}\textbf{1}_{n/k}^{\top},\alpha_{r+1}\mathbf{e}_{r+1}\textbf{1}_{n/k}^{\top},\cdots,\alpha_{k}\mathbf{e}_{k}\textbf{1}_{n/k}^{\top}\right]\in\mathbb{R}^{d\times n}$
with SVD
$\mathbf{X}=\underbrace{\begin{bmatrix}\sqrt{\frac{k}{n}}\textbf{1}_{n/k}&\cdots&\textbf{0}_{n/k}\\ \vdots&\ddots&\vdots\\ \textbf{0}_{n/k}&\cdots&\sqrt{\frac{k}{n}}\textbf{1}_{n/k}\end{bmatrix}}_{\mathbf{U}\in\mathbb{R}^{n\times k}}\underbrace{\begin{bmatrix}\sqrt{\frac{n}{k}}\alpha_{1}&&\\ &\ddots&\\ &&\sqrt{\frac{n}{k}}\alpha_{k}\end{bmatrix}}_{\bm{\Sigma}\in\mathbb{R}^{k\times k}}\underbrace{\begin{bmatrix}\mathbf{e}_{1}^{\top}\\ \vdots\\ \mathbf{e}_{k}^{\top}\end{bmatrix}}_{\mathbf{V}^{\top}\in\mathbb{R}^{k\times d}}.$
Observe that a skeleton subset of size $k$ with the $k$ scaled canonical bases $\left\{\alpha_{i}\mathbf{e}_{i}\middle|i\in[k]\right\}$ is sufficient to form an ID that exactly recovers $\mathbf{X}$. That is, the uniform distribution over the $k$ scaled canonical bases is preferred.
[…]
Such inefficiency is mirrored in the skeleton complexity $k\geq\frac{r-1}{\epsilon\eta_{r}}+\frac{1}{\epsilon}$ with a small $\eta_{r}$:
$\eta_{r}=\frac{k-r}{\alpha r+k-r}=\frac{\beta}{\alpha+\beta}\ll 1.$
For example, with $\beta=1$ and $\alpha\gg k$, the squared-norm sampling suffers from a skeleton complexity $\frac{k}{\epsilon}+\frac{\alpha(r-1)-r}{\epsilon}\gg k$ far greater than necessary.
It is worth highlighting that adaptiveness effectively remedies this inefficiency of squared-norm sampling. For instance, after picking $\alpha_{i}\mathbf{e}_{i}$ as the first skeleton, the adaptive update (e.g. Equation 4) eliminates the remaining $\frac{n}{k}-1$ rows of $\mathbf{X}$ in the $\mathbf{e}_{i}$ direction and therefore helps exclude the possible redundant samples.
As the well-studied alternatives to squared-norm sampling, determinantal point process (DPP sampling) [2, 12] and leverage score sampling [33] provide better skeleton complexity guarantees but with higher cost of constructing the corresponding distributions as a trade-off.
For example, the k-DPP sampling [31] is known for its nearly optimal skeleton complexity in expectation [2, 25, 5]:
(7) $k=\left|S\right|\geq\frac{r}{\epsilon}+r-1\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ \mathbb{E}_{S\sim\text{k-DPP}(\mathbf{X})}\left[\mathcal{E}_{\mathbf{X}}\left(S\right)\right]\leq(1+\epsilon)\left\|\mathbf{X}-{\mathbf{X}}_{\langle r\rangle}\right\|_{F}^{2}.$
As a trade-off, the classical SVD-based algorithm for k-DPP$(\mathbf{X})$ [27] takes $O\left(nd^{2}\right)$ to construct the distribution and $O\left(nk^{2}\right)$ to draw $k$ samples.
In contrast to sampling from sophisticated distributions, uniform sampling works well for incoherent matrices (whose singular vectors distribute evenly along all canonical bases) but easily fails on coherent ones [8]. Although an in-depth comparison of different sampling methods is beyond the scope of this work, we refer the interested reader to the following enlightening references: [13, 8, 12, 5].
3 Combining Adaptiveness and Randomness
We recall that the vulnerability of greedy pivoting (Remark 2.1) comes from the scarce adversarial inputs which can be negligible in expectation under randomization. Meanwhile, the vulnerability of squared-norm sampling (Example 2.4) can be effectively circumvented by incorporating adaptive updates. Therefore, a natural question is how to combine adaptiveness and randomness effectively for better skeleton selection. In this section, we review two existing ideas that involve both adaptiveness and randomness in different ways, while comparing their respective advantages and drawbacks.
3.1 Random Pivoting
Random pivoting [13, 14, 1, 5] is arguably the most intuitive skeleton selection method that combines adaptiveness and randomness. It can be simply described as “replacing the greedy squared-norm pivoting in CPQR with squared-norm sampling” or “adaptively updating the dataset via QR after each step of squared-norm sampling,” as formalized in Algorithm 1.
Algorithm 1 Sequential random pivoting (SRP)
1:(a) Data matrix $\mathbf{X}=\left[\mathbf{x}_{1},\cdots,\mathbf{x}_{n}\right]^{\top}\in\mathbb{R}^{n\times d}$, (b) Relative error tolerance $\tau=\left(1+\epsilon\right)\eta_{r}\in(0,1)$ (or target rank $k\in\mathbb{N}$),
2:(a) Skeleton indices $S=\left\{s_{1},\cdots,s_{k}\right\}\subset[n]$, (b) $\mathbf{L}\in\mathbb{R}^{n\times k}$, (c) $\bm{\pi}\in\mathfrak{S}_{n}$
3:$\mathbf{L}^{(0)}\leftarrow\textbf{0}_{n\times 0}$, $\mathbf{Q}^{(0)}\leftarrow\textbf{0}_{d\times 0}$, $\bm{\pi}\leftarrow[1,\cdots,n]$, $S\leftarrow\emptyset$, $t=0$
4:$\mathbf{d}^{(0)}\in\mathbb{R}_{\geq 0}^{n}$ with $\mathbf{d}^{(0)}(i)=\left\|\mathbf{x}_{i}\right\|_{2}^{2}\leavevmode\nobreak\ \forall\leavevmode\nobreak\ i\in[n]$triangleright $O(nd)$
5:while $\left\|\mathbf{d}^{(t)}\right\|_{1}<\tau\left\|\mathbf{d}^{(0)}\right\|_{1}$ (or $\left|S\right|<k$do
6:     $t\leftarrow t+1$
7:     $s_{t}\sim\mathbf{d}^{(t-1)}/\left\|\mathbf{d}^{(t-1)}\right\|_{1}$
8:     $S\leftarrow S\cup\left\{s_{t}\right\}$, $\bm{\pi}\leftarrow\mathtt{swap}\left(\bm{\pi},t,s_{t}\right)$triangleright $\bm{\pi}(1:t)$ consists of $S$
9:     $\mathbf{a}^{(t)}\leftarrow\textbf{0}_{n}$, $\bm{\pi}_{a}\leftarrow\bm{\pi}(t:n)$triangleright “a” for active
10:     $\mathbf{v}\leftarrow\mathbf{x}_{s_{t}}-\mathbf{Q}^{(t-1)}\left(\left(\mathbf{Q}^{(t-1)}\right)^{\top}\mathbf{x}_{s_{t}}\right)\in\mathbb{R}^{d}$triangleright $O(dt)$
11:     $\mathbf{Q}^{(t)}\leftarrow\left[\mathbf{Q}^{(t-1)},\mathbf{v}/\left\|\mathbf{v}\right\|_{2}\right]\in\mathbb{R}^{d\times t}$triangleright $O(d)$
12:     $\mathbf{a}^{(t)}\left(\bm{\pi}_{a}\right)\leftarrow\mathbf{X}\left(\bm{\pi}_{a},:\right)\mathbf{v}$triangleright $O\left(nd\right)$
13:     $\mathbf{L}^{(t)}\leftarrow\left[\mathbf{L}^{(t-1)},\mathbf{a}^{(t)}/\sqrt{\mathbf{a}^{(t)}\left(s_{t}\right)}\right]$triangleright $\mathbf{a}^{(t)}\left(s_{t}\right)=\mathbf{d}^{(t-1)}(s_{t})$
14:     $\mathbf{d}^{(t)}(i)\leftarrow 0\leavevmode\nobreak\ \forall\leavevmode\nobreak\ i\in S$, $\mathbf{d}^{(t)}(i)\leftarrow\mathbf{d}^{(t-1)}(i)-\frac{\left(\mathbf{a}^{(t)}(i)\right)^{2}}{\mathbf{a}^{(t)}\left(s_{t}\right)}\leavevmode\nobreak\ \forall\leavevmode\nobreak\ i\notin S$triangleright $O(n)$
15:$k\leftarrow\left|S\right|$, $\mathbf{L}\leftarrow\mathbf{L}^{(t)}$
Algorithm 1 has an asymptotic complexity of $O\left(ndk+dk^{2}\right)$, with $k$ inherently sequential passes through $\mathbf{X}$ (or its submatrices), while the storage of $\mathbf{L}^{(t)}$ and $\mathbf{Q}^{(t)}$ requires $O(nk)$ and $O(dk)$ memory, respectively. Meanwhile, analogous to greedy pivoting (Remark 2.2), Algorithm 1 is error-revealing thanks to its adaptive nature.
The idea of random pivoting in Algorithm 1 is ubiquitous in various related problems, e.g. the combination of volume sampling and adaptive sampling for low-rank approximation [14], the Randomly Pivoted Cholesky (RPCholesky) [5] for kernel column Nyström approximation [35, 19.2], and the $D^{2}$-sampling for k-means++ clustering [1].
Specifically, applying RPCholesky [5, Algorithm 2.1] to the kernel matrix $\mathbf{X}\mathbf{X}^{\top}\in\mathbb{R}^{n\times n}$ can be reformulated as Algorithm 2, which is equivalent to Algorithm 1 in the exact arithmetic (Remark 3.1). Compared to Algorithm 1, the skeleton selection stage of Algorithm 2 has an asymptotic complexity of $O\left(ndk+nk^{2}\right)$ (cf. $O\left(ndk+dk^{2}\right)$ for Algorithm 1), while only requiring $O(nk)$ memory for the storage of $\mathbf{L}^{(t)}$.
Algorithm 2 RPCholesky [5, Algorithm 2.1]
1:$\mathbf{L}^{(0)}\leftarrow\textbf{0}_{n\times 0}$, $\bm{\pi}\leftarrow[1,\cdots,n]$, $S\leftarrow\emptyset$, $t=0$
2:$\mathbf{d}^{(0)}\in\mathbb{R}_{\geq 0}^{n}$ with $\mathbf{d}^{(0)}(i)=\left\|\mathbf{x}_{i}\right\|_{2}^{2}\leavevmode\nobreak\ \forall\leavevmode\nobreak\ i\in[n]$triangleright $O(nd)$
3:while $\left\|\mathbf{d}^{(t)}\right\|_{1}<\tau\left\|\mathbf{d}^{(0)}\right\|_{1}$ (or $\left|S\right|<k$do
4:     $t\leftarrow t+1$
5:     $s_{t}\sim\mathbf{d}^{(t-1)}/\left\|\mathbf{d}^{(t-1)}\right\|_{1}$
6:     $S\leftarrow S\cup\left\{s_{t}\right\}$, $\bm{\pi}\leftarrow\mathtt{swap}\left(\bm{\pi},t,s_{t}\right)$triangleright $\bm{\pi}(1:t)$ consists of $S$
7:     $\mathbf{a}^{(t)}\leftarrow\textbf{0}_{n}$, $\bm{\pi}_{a}\leftarrow\bm{\pi}(t:n)$triangleright “a” for active
8:     $\mathbf{a}^{(t)}\left(\bm{\pi}_{a}\right)\leftarrow\mathbf{X}\left(\bm{\pi}_{a},:\right)\mathbf{x}_{s_{t}}-\mathbf{L}^{(t-1)}\left(\bm{\pi}_{a},1:t-1\right){\mathbf{L}^{(t-1)}\left(s_{t},1:t-1\right)}^{\top}$triangleright $O\left(nd+nt\right)$
9:     $\mathbf{L}^{(t)}\leftarrow\left[\mathbf{L}^{(t-1)},\mathbf{a}^{(t)}/\sqrt{\mathbf{a}^{(t)}\left(s_{t}\right)}\right]$triangleright $\mathbf{a}^{(t)}\left(s_{t}\right)=\mathbf{d}^{(t-1)}(s_{t})$
10:     $\mathbf{d}^{(t)}(i)\leftarrow 0\leavevmode\nobreak\ \forall\leavevmode\nobreak\ i\in S$, $\mathbf{d}^{(t)}(i)\leftarrow\mathbf{d}^{(t-1)}(i)-\frac{\left(\mathbf{a}^{(t)}(i)\right)^{2}}{\mathbf{a}^{(t)}\left(s_{t}\right)}\leavevmode\nobreak\ \forall\leavevmode\nobreak\ i\notin S$triangleright $O(n)$
Remark 3.1 (Relation between QR with sequential random pivoting and RPCholesky).
Sequential random pivoting (Algorithm 1) on $\mathbf{X}$ is equivalent to RPCholesky (Algorithm 2 [5, Algorithm 2.1]) on $\mathbf{X}\mathbf{X}^{\top}$ in the exact arithmetic.
Proof 3.2 (Rationale for Remark 3.1).
To show the equivalence in the exact arithmetic, it is sufficient to observe that Algorithms 1 and 2 yield the same $\mathbf{a}^{(t)}$ and $\mathbf{d}^{(t)}$ in each iteration $t=1,2,\cdots$.
By induction, when $t=1$, both Algorithms 1 and 2 lead to $\mathbf{a}^{(1)}=\mathbf{X}\mathbf{x}_{s_{1}}$ and
$\mathbf{d}^{(1)}(i)=\left\|\mathbf{x}_{i}\right\|_{2}^{2}-\frac{\left(\mathbf{x}_{i}^{\top}\mathbf{x}_{s_{1}}\right)^{2}}{\left\|\mathbf{x}_{s_{1}}\right\|_{2}^{2}}=\left\|\left(\mathbf{I}_{d}-\mathbf{Q}^{(1)}\left(\mathbf{Q}^{(1)}\right)^{\top}\right)\mathbf{x}_{i}\right\|_{2}^{2}\quad\forall\leavevmode\nobreak\ i\in[n]\setminus\left\{s_{1}\right\},$
while $\mathbf{d}^{(1)}(s_{1})=0$. We also observe that $\mathbf{L}^{(1)}=\mathbf{X}\mathbf{Q}^{(1)}$ since $\mathbf{a}^{(1)}\left(s_{1}\right)=\left\|\mathbf{x}_{s_{1}}\right\|_{2}^{2}$.
Suppose that in the $(t-1)$-th iteration, Algorithms 1 and 2 produce equivalent $\mathbf{a}^{(t-1)}$ and $\mathbf{d}^{(t-1)}$ under the exact arithmetic, while $\mathbf{d}^{(t-1)}\left(i\right)=\left\|\left(\mathbf{I}_{d}-\mathbf{Q}^{(t-1)}\left(\mathbf{Q}^{(t-1)}\right)^{\top}\right)\mathbf{x}_{i}\right\|_{2}^{2}$ for all $i\in[n]$ and $\mathbf{L}^{(t-1)}=\mathbf{X}\mathbf{Q}^{(t-1)}$.
Then in the $t$-th iteration, both Algorithms 1 and 2 compute
$\mathbf{a}^{(t)}=\mathbf{X}\mathbf{x}_{s_{t}}-\mathbf{X}\mathbf{Q}^{(t-1)}\left(\mathbf{Q}^{(t-1)}\right)^{\top}\mathbf{x}_{s_{t}}=\mathbf{X}\left(\mathbf{I}_{d}-\mathbf{Q}^{(t-1)}\left(\mathbf{Q}^{(t-1)}\right)^{\top}\right)\mathbf{x}_{s_{t}},$
$\mathbf{L}^{(t)}=\mathbf{X}\mathbf{Q}^{(t)}$, and for all $i\in[n]$,
$\mathbf{d}^{(t)}(i)=$ $\left\|\left(\mathbf{I}_{d}-\mathbf{Q}^{(t-1)}\left(\mathbf{Q}^{(t-1)}\right)^{\top}\right)\mathbf{x}_{i}\right\|_{2}^{2}-\frac{\left(\mathbf{x}_{i}^{\top}\left(\mathbf{I}_{d}-\mathbf{Q}^{(t-1)}\left(\mathbf{Q}^{(t-1)}\right)^{\top}\right)\mathbf{x}_{s_{t}}\right)^{2}}{\mathbf{x}_{s_{t}}^{\top}\left(\mathbf{I}_{d}-\mathbf{Q}^{(t-1)}\left(\mathbf{Q}^{(t-1)}\right)^{\top}\right)\mathbf{x}_{s_{t}}}$
$=$ $\left\|\left(\mathbf{I}_{d}-\mathbf{Q}^{(t)}\left(\mathbf{Q}^{(t)}\right)^{\top}\right)\mathbf{x}_{i}\right\|_{2}^{2}.$
In light of the equivalence between Algorithms 1 and 2, [5, Theorem 3.1] provides the following skeleton complexity guarantee for sequential random pivoting.
Proposition 3.3 (Sequential random pivoting).
For any $\epsilon>0$ and $r\in[n]$, the skeleton subset $S$ selected by Algorithm 1 provides $\mathbb{E}\left[\mathcal{E}_{\mathbf{X}}\left(S\right)\right]\leq\left(1+\epsilon\right)\left\|\mathbf{X}-{\mathbf{X}}_{\langle r\rangle}\right\|_{F}^{2}$ when333 Without strictly specifying, we generally assume that $\epsilon>0$ is small enough such that the $(r,\epsilon)$-ID is a non-trivial approximation, i.e. $\epsilon<\min\left\{1/\eta_{r},2^{r}\right\}$ such that both logarithmic terms are positive. Otherwise, the skeleton complexity guarantee still holds by replacing any negative logarithmic terms with zeros.
$k=\left|S\right|\geq\frac{r}{\epsilon}+r\cdot\min\left\{\log\left(\frac{1}{\epsilon\eta_{r}}\right),1+\log\left(\frac{2^{r}}{\epsilon}\right)\right\}.$
From Proposition 3.3, we notice that sequential random pivoting almost matches the nearly optimal skeleton complexity guarantee of DPP (Equation 7) up to some logarithmic factors $\log\left(\frac{1}{\epsilon\eta_{r}}\right)$ on $r$. Further, the appealing skeleton complexity of random pivoting is empirically demonstrated in the form of RPCholesky for column Nyström approximation in [5].
Although random pivoting leverages the strength of both adaptiveness and randomness and achieves an attractive skeleton complexity with affordable asymptotic complexity, the lack of parallelizability due to its sequential nature becomes a critical concern when considering empirical runtime.
Remark 3.4 (Inefficiency of sequential updates).
Given a sequence of matrix operations, it is well-known that an appropriate implementation using Level-3 BLAS, or matrix-matrix, operations will run more efficiently on modern processors than an optimal implementation using Level-2 or Level-1 BLAS [3]. This is largely due to the greater potential for the Level-3 BLAS to make more efficient use of memory caching in the processor. Notice that the computational bottleneck of Algorithm 1 with asymptotic complexity $O\left(ndk\right)$ (or $O\left(ndk\right)$) consists of $k$ matrix-vector multiplications (Level-2 BLAS) with $\mathbf{X}$. This leads to the considerable slowdown of Algorithm 1 in practice.
3.2 Sketchy Pivoting
Alternative to random pivoting, randomness can be combined with adaptiveness through randomized linear embedding (also known as sketching), which leads to the sketchy pivoting methods [34, 48, 16, 19]. As given in Algorithm 3, the general framework of sketchy pivoting [16, Algorithm 1] consists of two stages.
Algorithm 3 Sketchy pivoting [48, 16] (SkLUPP/SkCPQR)
1:(a) $\mathbf{X}=\left[\mathbf{x}_{1},\cdots,\mathbf{x}_{n}\right]^{\top}\in\mathbb{R}^{n\times d}$, (b) Target rank $k\in\mathbb{N}$, (c) Sample size $l\geq k$, (d) Pivoting strategy: LUPP/CPQR
2:(a) $S=\left\{s_{1},\cdots,s_{k}\right\}\subset[n]$, (b) $\mathbf{L}\in\mathbb{R}^{n\times k}$, (c) $\bm{\pi}\in\mathfrak{S}_{n}$
3:Draw a randomized linear embedding $\bm{\Omega}\in\mathbb{R}^{d\times l}$
4:$\mathbf{Y}\leftarrow\mathbf{X}\bm{\Omega}\in\mathbb{R}^{n\times l}$triangleright $O\left(ndk\right)$
5:if LUPP then
[…]
7:else if CPQR then
[…]
In the first stage, randomness is incorporated through sketching, which serves two purposes.
  • From the computational efficiency perspective, sketching reduces the data dimension from $d$ to $l=O(k)$444 For applications like low-rank approximations and skeleton selection, a small constant oversampling like $k+5\leq l\leq k+10$ is usually sufficient in practice. , while transferring the $O(ndk)$ computational bottleneck from the sequential greedy pivoting to the (input-sparsity-time and) parallelizable matrix-matrix multiplications.
  • From the skeleton complexity perspective, applying randomized embeddings to the row space of $\mathbf{X}$ via sketching improves the empirical robustness of the subsequent greedy pivoting [45, 16]. Intuitively, this is because the scarce adversarial inputs for greedy pivoting (Remark 2.1) are effectively negligible under randomization.
Common choices of randomized linear embeddings in the sketching stage include Gaussian embeddings [28] (with asymptotic complexity $O\left(ndk\right)$), subsampled randomized trigonometric transforms [50, 43, 46, 4] (with asymptotic complexity $O\left(nd\log(k)\right)$), and sparse embeddings [36, 37, 7, 47] (with asymptotic complexity $O\left(nd\right)$). We refer the interested readers to [26, 49, 35] for a general review of sketching.
In the second stage, adaptiveness is utilized as greedy pivoting on the sketched sample matrix $\mathbf{Y}=\mathbf{X}\bm{\Omega}\in\mathbb{R}^{n\times l}$ to select skeletons. Both LUPP [22, Section 3.4] and CPQR [22, Section 5.4.2] on $\mathbf{Y}$ cost $O(nk^{2})$ asymptotically, whereas LUPP is remarkably faster in practice due to its superior parallelizability [16]. Intuitively, applying greedy pivoting on the top of sketching can be viewed as an analog of the squared-norm sampling with a slightly different squared-norm-based distribution.
For example, for a Gaussian random matrix $\bm{\Omega}$ with i.i.d. entries from $\mathcal{N}\left(0,1/l\right)$, each row $\mathbf{y}_{i}=\bm{\Omega}^{\top}\mathbf{x}_{i}$ of $\mathbf{Y}$ is a Gaussian random vector with i.i.d. entries from $\mathcal{N}\left(\textbf{0},\left\|\mathbf{x}_{i}\right\|_{2}^{2}\mathbf{I}_{l}\right)$555 It is worth highlighting that although rows of $\mathbf{Y}$ consist of i.i.d. entries, entries in each column of $\mathbf{Y}$ are dependent. In particular, for all $j\in[l]$, $\mathrm{Cov}\left(\mathbf{Y}\left(:,j\right)\right)=\mathbf{X}\mathbf{D}^{-1}\mathbf{X}^{\top}$ where $\mathbf{D}=\mathop{\mathrm{diag}}\left(\left\|\mathbf{x}_{1}\right\|_{2}^{2},\cdots,\left\|\mathbf{x}_{n}\right\|_{2}^{2}\right)$. . When applying greedy squared-norm pivoting (CPQR) in the second stage, the first pivot is selected according to $s_{1}\leftarrow\mathop{\mathrm{argmax}}_{i\in[n]}\left\|\mathbf{x}_{i}\right\|_{2}^{2}Z_{i}$ where every $Z_{i}=\sum_{j=1}^{l}Y_{ij}^{2}$ is a $\chi^{2}_{l}$ random variable but $\left\{Z_{i}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ i\in[n]\right\}$ are dependent. The intuition behind sketchy pivoting is similar to random pivoting, but the randomness arises from the random variables $Z_{i}$, instead of sampling.
In the extreme scenarios, when the variance of $\left\{Z_{i}\right\}_{i\in[n]}$ is negligible in comparison to the squared norms $\left\{\left\|\mathbf{x}_{i}\right\|_{2}^{2}\right\}_{i\in[n]}$, sketchy pivoting behaves similarly to greedy pivoting. Meanwhile, when the variance of $\left\{Z_{i}\right\}_{i\in[n]}$ is much larger than differences among the squared norms, sketchy pivoting tends to behave like uniform sampling. Intuitively, there exists some $\left\{Z_{i}\right\}_{i\in[n]}$ with appropriate variance interpolating the two extreme cases such that sketchy pivoting mimics the behavior of squared-norm sampling and enjoys a nearly optimal skeleton complexity guarantee similar to Proposition 3.3.
In practice, sketchy pivoting (Algorithm 3) with Gaussian embedding provides high-quality skeleton selection with comparable error $\mathcal{E}_{\mathbf{X}}\left(S\right)$ to random pivoting (Algorithm 1, cf. Figure 1), but much more efficiently thanks to the parallelizability of sketching [16]. However, as a major empirical limitation, sketchy pivoting is not error-revealing and requires prior knowledge of the target rank $k$.
4 Robust Blockwise Random Pivoting
In light of the existing skeleton selection methods summarized in Table 1 that leverage/combine randomness and adaptiveness, a parallelizable, error-revealing, and exactly ID-revealing skeleton selection algorithm that attains similar skeleton and asymptotic complexities as the sequential random pivoting is a desirable missing piece.
Blockwise random pivoting is a natural extension of its sequential variation Algorithm 1 that consists of matrix-matrix multiplications and is therefore parallelizable. In the kernel formulation (with an input kernel matrix $\mathbf{X}\mathbf{X}^{\top}$), [5, Algorithm 2.2] introduced a blocked version of RPCholesky that can be generalized as blockwise random pivoting (BRP—Algorithm 4 with $\tau_{b}=0$). However, for a large block size $b\in\mathbb{N}$ that improves parallelizability, such plain BRP can pick up to $b$ times more skeletons than necessary due to the similar pitfall of squared-norm sampling instantiated in Example 2.4 (e.g. Figure 1). Moreover, in practice, the numerical instability issue tends to be exacerbated as the block size increases.
Algorithm 4 Robust blockwise random pivoting (RBRP)
1:(a) $\mathbf{X}=\left[\mathbf{x}_{1},\cdots,\mathbf{x}_{n}\right]^{\top}\in\mathbb{R}^{n\times d}$, (b) $\tau=\left(1+\epsilon\right)\eta_{r}\in(0,1)$ (or $k\in\mathbb{N}$), (c) Block size $b\in\mathbb{N}$, (d) $\tau_{b}\in[0,1)$—Tolerance for blockwise filtering (we choose $\tau_{b}=\frac{1}{b}$)
2:(a) $S=\left\{s_{1},\cdots,s_{k}\right\}\subset[n]$, (b) $\mathbf{L}\in\mathbb{R}^{n\times k}$, (c) $\bm{\pi}\in\mathfrak{S}_{n}$
3:$\mathbf{L}^{(0)}\leftarrow\textbf{0}_{n\times 0}$, $\mathbf{Q}^{(0)}\leftarrow\textbf{0}_{d\times 0}$, $\bm{\pi}\leftarrow[1,\cdots,n]$, $S\leftarrow\emptyset$, $t=0$
4:$\mathbf{d}^{(0)}\in\mathbb{R}_{\geq 0}^{n}$ with $\mathbf{d}^{(0)}(i)=\left\|\mathbf{x}_{i}\right\|_{2}^{2}\leavevmode\nobreak\ \forall\leavevmode\nobreak\ i\in[n]$triangleright $O(nd)$
5:while $\left\|\mathbf{d}^{(t)}\right\|_{1}>\tau\left\|\mathbf{d}^{(0)}\right\|_{1}$ (or $\left|S\right|<k$do
6:     $t\leftarrow t+1$, $b\leftarrow\min(b,k-\left|S\right|)$
7:     Sample $S_{t}=\left(s_{\left|S\right|+1},\cdots,s_{\left|S\right|+b}\right)\sim\mathbf{d}^{(t-1)}/\left\|\mathbf{d}^{(t-1)}\right\|_{1}$ without replacement $\triangleright$ $O(n)$
8:     $\mathbf{V}\leftarrow\mathbf{X}\left(S_{t},:\right)^{\top}-\mathbf{Q}^{(t-1)}\left(\left(\mathbf{Q}^{(t-1)}\right)^{\top}\mathbf{X}\left(S_{t},:\right)^{\top}\right)\in\mathbb{R}^{d\times b}$triangleright $O(db\left|S\right|)$
[…]
10:     $b^{\prime}\leftarrow\max\left\{i\in[b]\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \left\|\mathbf{R}_{\mathbf{V}}\left(i:b,i:b\right)\right\|_{F}^{2}\geq\tau_{b}\cdot\left\|\mathbf{R}_{\mathbf{V}}\right\|_{F}^{2}\right\}$triangleright $O\left(b^{2}\right)$
11:     $S^{\prime}_{t}\leftarrow S_{t}\left(\bm{\pi}_{\mathbf{V}}\left(1:b^{\prime}\right)\right)$, $\mathbf{Q}^{\prime}_{\mathbf{V}}\leftarrow\mathbf{Q}_{\mathbf{V}}\left(:,1:b^{\prime}\right)$
12:     $\mathbf{Q}^{(t)}\leftarrow\left[\mathbf{Q}^{(t-1)},\mathbf{Q}^{\prime}_{\mathbf{V}}\right]\in\mathbb{R}^{d\times\left(\left|S\right|+b^{\prime}\right)}$
13:     $\bm{\pi}\leftarrow\mathtt{swap}\left(\bm{\pi},\left|S\right|+1:\left|S\right|+b^{\prime},S^{\prime}_{t}\right)$, $\bm{\pi}_{a}\leftarrow\bm{\pi}(\left|S\right|+1:n)$
14:     $\widehat{\mathbf{L}}^{(t)}\leftarrow\textbf{0}_{n\times b^{\prime}}$
15:     $\widehat{\mathbf{L}}^{(t)}\left(\bm{\pi}_{a},:\right)\leftarrow\mathbf{X}\left(\bm{\pi}_{a},:\right)\mathbf{Q}^{\prime}_{\mathbf{V}}$triangleright $O\left(ndb^{\prime}\right)$
16:     $S\leftarrow S\cup S^{\prime}_{t}$
17:     $\mathbf{L}^{(t)}\leftarrow\left[\mathbf{L}^{(t-1)},\widehat{\mathbf{L}}^{(t)}\right]\in\mathbb{R}^{n\times\left|S\right|}$
18:     $\mathbf{d}^{(t)}(i)\leftarrow 0\leavevmode\nobreak\ \forall\leavevmode\nobreak\ i\in S$, $\mathbf{d}^{(t)}(i)\leftarrow\mathbf{d}^{(t-1)}(i)-\left\|\widehat{\mathbf{L}}^{(t)}\left(i,:\right)\right\|_{2}^{2}\leavevmode\nobreak\ \forall\leavevmode\nobreak\ i\notin S$triangleright $O(nb^{\prime})$
19:$k\leftarrow\left|S\right|$, $\mathbf{L}\leftarrow\mathbf{L}^{(t)}$
To exemplify the pitfall of plain BRP, we present an adversarial input in Example 4.1 based on a tailored Gaussian mixture model (GMM).
Example 4.1 (Pitfall of plain blockwise random pivoting).
For $n,d,k\in\mathbb{N}$ such that $n/k=m\in\mathbb{N}$, we draw $n$ points from a GMM with means $\mathcal{C}=\left\{\bm{\mu}_{j}\right\}_{j\in[k]}$, covariance $\mathbf{I}_{d}$, and cluster size $m$$\mathcal{X}=\left\{\mathbf{x}_{i}\in\mathbb{R}^{d}\right\}_{i\in[n]}=\bigcup_{j=1}^{k}\mathcal{X}_{j}$ where
$\mathcal{X}_{j}=\left\{\mathbf{x}_{m(j-1)+\iota}=\bm{\mu}_{j}+\bm{\xi}_{\iota}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \bm{\mu}_{j}=10j\cdot\mathbf{e}_{j},\leavevmode\nobreak\ \bm{\xi}_{\iota}\sim\mathcal{N}\left(\textbf{0}_{d},\mathbf{I}_{d}\right)\leavevmode\nobreak\ \emph{i.i.d.}\leavevmode\nobreak\ \forall\leavevmode\nobreak\ \iota\in[m]\right\}.$
Consider a GMM data matrix $\mathbf{X}=\left[\mathbf{x}_{1},\cdots,\mathbf{x}_{n}\right]^{\top}\in\mathbb{R}^{n\times d}$. Since the means in $\mathcal{C}$ have distinct norms whose discrepancies dominate the covariance $\mathbf{I}_{d}$, $\mathcal{X}$ can be partitioned into $k$ clusters $\left\{\mathcal{X}_{j}\right\}_{j\in[k]}$ each containing $m$ points with distinct norms.
The best size-$k$ skeleton subset consists of exactly one point from each cluster $\mathcal{X}_{j}$, while multiple points from the same cluster are redundant. However, the plain blockwise adaptiveness (random (BRP)/greedy pivoting (BGP)) tends to pick multiple points from the same cluster. For instance, with block size $b$, BGP can pick up to $\min\left(b,m\right)$ points in the same cluster. Similar inefficiency also appears in BRP but is generally alleviated thanks to randomness (Figure 1 (left)).
To resolve the challenges illustrated by Example 4.1, we introduce a robust blockwise random pivoting (RBRP) algorithm—Algorithm 4—that empirically achieves comparable skeleton complexities to the sequential adaptive methods (SRP and CPQR), as demonstrated in Figure 1 (right), while exploiting the parallelizability of matrix-matrix multiplications.
Remark 4.2 (Robust blockwise filtering).
The key step in Algorithm 4 that improves the robustness of BRP is the robust blockwise filtering with tolerance $\tau_{b}\in[0,1]$—applying truncated CPQR locally to the small residuals of the selected candidates $\mathbf{V}\in\mathbb{R}^{d\times b}$:
$\mathbf{V}\left(:,\bm{\pi}_{\mathbf{V}}\right)=\mathbf{Q}_{\mathbf{V}}\mathbf{R}_{\mathbf{V}}\approx\mathbf{Q}_{\mathbf{V}}\left(:,1:b^{\prime}\right)\mathbf{R}_{\mathbf{V}}\left(1:b^{\prime},:\right)$
where the relative truncation error is upper-bounded by $\tau_{b}$
$\left\|\mathbf{V}\left(:,\bm{\pi}_{\mathbf{V}}\right)-\mathbf{Q}_{\mathbf{V}}\left(:,1:b^{\prime}\right)\mathbf{R}_{\mathbf{V}}\left(1:b^{\prime},:\right)\right\|_{F}^{2}=\left\|\mathbf{R}_{\mathbf{V}}\left(b^{\prime}+1:b,b^{\prime}+1:b\right)\right\|_{F}^{2}<\tau_{b}\left\|\mathbf{V}\right\|_{F}^{2}.$
With small constant block sizes, the robust blockwise filtering based on local CPQR can be computed with negligible additional cost $O(db^{2})$, despite the sequential nature of CPQR.
In general, Algorithm 4 takes $O\left(ndk+bdk^{2}\right)$ time asymptotically (with a slightly larger lower-order term, cf. $O\left(ndk+dk^{2}\right)$ for the sequential version in Algorithm 1), where the dominant cost $O(ndk)$ consists of $k$ parallelizable passes through $\mathbf{X}$. Analogous to Algorithm 1, the storage of $\mathbf{L}^{(t)}$ and $\mathbf{Q}^{(t)}$ requires $O(nk)$ and $O(dk)$ memory, respectively. Meanwhile, the adaptiveness grants Algorithm 4 the blockwise error-revealing ability.
Table 2: Summary of abbreviations for the skeleton selection methods in the experiments.
Abbreviation Skeleton selection method Algorithm
CPQR Sequential greedy pivoting [22, Section 5.4.2]
SqNorm Squared-norm sampling Proposition 2.3 [13]
DPP DPP sampling [2, Equation 8][31, 21]
SRP Sequential random pivoting Algorithm 1 [5]
SkCPQR Sketchy pivoting with CPQR Algorithm 3 [48]
SkLUPP Sketchy pivoting with LUPP Algorithm 3 [16]
BGP Blockwise greedy pivoting Remark 4.3 with $\tau_{b}=0$
BRP Blockwise random pivoting Algorithm 4 with $\tau_{b}=0$
RBGP Robust blockwise greedy pivoting Remark 4.3 with $\tau_{b}=\frac{1}{b}$
RBRP Robust blockwise random pivoting Algorithm 4 with $\tau_{b}=\frac{1}{b}$
Remark 4.3 (Blockwise greedy pivoting).
For the completeness of comparison, in the experiments, we consider a variation of the (robust) blockwise random pivoting ((R)BRP)—(robust) blockwise greedy pivoting ((R)BGP). In Algorithm 4, (R)BGP replaces the sampling step,
sample $S_{t}=\left(s_{\left|S\right|+1},\cdots,s_{\left|S\right|+b}\right)\sim\mathbf{d}^{(t-1)}/\left\|\mathbf{d}^{(t-1)}\right\|_{1}$ without replacement,
with the task of choosing points corresponding to the top-$b$ probabilities greedily:
select $S_{t}=\left(s_{\left|S\right|+1},\cdots,s_{\left|S\right|+b}\right)\leftarrow$ indices of the top-$b$ entries in $\mathbf{d}^{(t-1)}$.
With $\tau_{b}=0$, BGP is reduced to CPQR. For adversarial inputs like the one in Example 4.1, BGP tends to suffer from far worse skeleton complexities than those of BRP (Figure 1 (left)).
Figure 1: The skeletonization error $\mathcal{E}_{\mathbf{X}}\left(S\right)$ of algorithms in Table 2 on a GMM-based adversarial input as described in Example 4.1, where the adaptiveness is critical for achieving a good skeleton complexity/accuracy. (Left): Sequential random pivoting (SRP) and greedy pivoting (CPQR) enjoy the best skeleton complexities. The plain blockwise random pivoting (BRP) and greedy pivoting (BGP) tend to suffer from observably higher skeleton complexities than the sequential adaptive methods, while greedy selection tends to severely exacerbate such inefficiency. The squared-norm sampling (SqNorm) can be viewed as plain BRP with block size $b=k$, which suffers from a similar inefficiency as BRP/BGP. DPP sampling attains relatively good accuracy even without adaptiveness, thanks to its nearly optimal skeleton complexity [2, 25, 5], but with a much higher runtime in practice as a trade-off (e.g., with [21], k-DPP takes about $10^{3}\times$ longer than other algorithms in Table 2). (Right): With robust blockwise filtering (Remark 4.2), robust blockwise random pivoting (RBRP) and greedy pivoting (RBGP), as well as the sketchy pivoting algorithms, achieves the comparable, nearly optimal skeleton complexities as the sequential adaptive methods. It is worth highlighting that despite the comparable skeleton complexities between RBRP and RBGP, RBGP is generally much slower than RBRP (cf. Figure 2). This is because, without randomness, blockwise greedy pivoting tends to pick more redundant points in each block, which are later filtered out.
By comparing Figure 1 (left) and (right), we observe that the robust blockwise filtering (Remark 4.2) improves both blockwise random and greedy pivoting. In particular, with $\tau_{b}=\frac{1}{b}$, RBRP (and RBGP) empirically attains comparable skeleton complexities to those of the sequential random pivoting on the GMM adversarial input (Example 4.1). The same observation extends to all the natural and synthetic data matrices tested in our experiments (Section 6), which leads to the following 4.4:
Conjecture 4.4 (Skeleton complexity of RBRP).
RBRP (Algorithm 4) with $\tau_{b}=\frac{1}{b}$ shares a skeleton complexity similar to sequential random pivoting (Proposition 3.3) in expectation.
Proof 4.5 (Rationale for 4.4).
Consider Algorithm 4 at the $t$-th step with $\left|S\right|$ selected skeletons from the previous steps, $\left|S_{t}\right|=b$ candidates drawn from $\frac{\mathbf{d}^{(t-1)}}{\left\|\mathbf{d}^{(t-1)}\right\|_{1}}$, and $\left|S^{\prime}_{t}\right|=b^{\prime}$ remaining skeletons in $S_{t}$ that pass through the robust blockwise filtering (Remark 4.2). Now suppose we switch to sequential random pivoting after the $(t-1)$-th step and draw $b^{\prime}$ skeletons adaptively, denoted as $S^{\prime\prime}_{t}$. Then, the key rationale for 4.4 is that $\mathcal{E}_{\mathbf{X}}\left(S\cup S^{\prime}_{t}\right)\approx\mathcal{E}_{\mathbf{X}}\left(S\cup S^{\prime\prime}_{t}\right)$ with (a reasonably) high probability when $\tau_{b}=\frac{1}{b}$.
To see this, we observe the close connection between the selection scheme of $S^{\prime}_{t}$ and that of $S^{\prime\prime}_{t}$:
  • The robust blockwise filtering for $S^{\prime}_{t}$ in RBRP adaptively selects the point with the maximum residual norm in each step (i.e. CPQR on the residual matrix $\mathbf{V}$ in Algorithm 4), within a small subset $S_{t}$ of $b$ points sampled according to the squared-norm distribution over $[n]\setminus S$.
  • The random pivoting for $S^{\prime\prime}_{t}$ in SRP adaptively selects the point according to the squared-norm distribution of the residual over the complement of the skeleton subset in each step.
Intuitively, every selection made for $S^{\prime}_{t}$ and $S^{\prime\prime}_{t}$ yields similar decay in the skeletonization error if the squared norms of the residual matrix, corresponding to remaining points in $S_{t}$, are around the maxima in the complement of the skeleton subset, so that these remaining points in $S_{t}$ can be selected by SRP with high probability666 It is worth mentioning that, in contrast to SRP, RBRP does not recompute the entire residual matrix after each skeleton selection. Instead, within each block, RBRP computes the residual matrix locally for $\mathbf{X}\left(S_{t},:\right)$ only, while updating the entire residual matrix at the end of each block via matrix-matrix multiplication. . Instead of computing the squared norms of the entire residual matrix explicitly for each selection as SRP (which leads to the undesired sequential updates), RBRP leverages the threshold $\tau_{b}=\frac{1}{b}$ to encourage the first $b^{\prime}$ points selected by the robust blockwise filtering in each block to have top squared norms in the corresponding residual matrices. Notice that with $\tau_{b}=\frac{1}{b}$, we enforce the sum of squared norms of the $b^{\prime}$ points to be larger than the average squared norm of the original block: $\left\|\mathbf{R}_{\mathbf{V}}\left(i:b,i:b\right)\right\|_{F}^{2}\geq\frac{1}{b}\left\|\mathbf{V}\right\|_{F}^{2}$. Since the squared norms of residuals are non-increasing under QR updates, $\frac{1}{b}\left\|\mathbf{V}\right\|_{F}^{2}$ serves as an overestimate for the squared norms of redundant points sampled in $S_{t}$.
As toy examples, if $\mathbf{V}$ is rank deficient (i.e. $\mathop{\mathrm{rank}}(\mathbf{V})=b^{\prime\prime}<b$), then there exists $b^{\prime}\leq b^{\prime\prime}$ such that $\left\|\mathbf{R}_{\mathbf{V}}\left(b^{\prime}:b,b^{\prime}:b\right)\right\|_{F}^{2}\geq\frac{1}{b}\left\|\mathbf{V}\right\|_{F}^{2}$, but
$0=\left\|\mathbf{R}_{\mathbf{V}}\left(b^{\prime\prime}+1:b,b^{\prime\prime}+1:b\right)\right\|_{F}^{2}\leq\left\|\mathbf{R}_{\mathbf{V}}\left(b^{\prime}+1:b,b^{\prime}+1:b\right)\right\|_{F}^{2}<\frac{1}{b}\left\|\mathbf{V}\right\|_{F}^{2}.$
Whereas if $\mathbf{V}\in\mathbb{R}^{d\times b}$ contains orthogonal columns with the same squared norm, then for all $i\in[b]$, we have $\left\|\mathbf{R}_{\mathbf{V}}\left(i:b,i:b\right)\right\|_{F}^{2}\geq\left|\mathbf{R}_{\mathbf{V}}\left(b,b\right)\right|^{2}=\frac{1}{b}\left\|\mathbf{V}\right\|_{F}^{2}$, and therefore $b^{\prime}=b$.
5 Interpolation Matrix Construction
In this section, we consider the construction of the interpolation matrix $\mathbf{W}$ after skeleton selection. In particular, we are interested in exploiting information from the $O\left(ndk\right)$ skeleton selection algorithms in Table 1, i.e. the matrix $\mathbf{L}\in\mathbb{R}^{n\times k}$ in Algorithms 1, 3, and 4, to construct $\mathbf{W}$ efficiently in $O(nk^{2})$ time.
We begin by noting the optimal interpolation matrix that solves the least square problem
(8) $\min_{\mathbf{W}\in\mathbb{R}^{n\times k}}\left\|\mathbf{X}-\mathbf{W}\mathbf{X}_{S}\right\|_{F}^{2}\quad\Rightarrow\quad\mathbf{W}=\mathbf{X}\mathbf{X}_{S}^{\dagger}=\mathbf{X}\mathbf{X}_{S}^{\top}\left(\mathbf{X}_{S}\mathbf{X}_{S}^{\top}\right)^{-1},$
so that we can compute such optimal $\mathbf{W}$ exactly in $O\left(ndk+dk^{2}+nk^{2}\right)$ time via QR decomposition on $\mathbf{X}_{S}$. However, the dominant cost $O(ndk)$ requires additional passes through $\mathbf{X}$ and can be prohibitive after the skeleton selection stage.
5.1 Exactly ID-revealing Algorithms
With the output matrix $\mathbf{L}\in\mathbb{R}^{n\times k}$ from Algorithms 1 and 4, the corresponding optimal interpolation matrix $\mathbf{X}\mathbf{X}_{S}^{\dagger}$ can be computed in $O(nk^{2})$ time following Algorithm 5.
Algorithm 5 Interpolation matrix construction (ID)
1:(a) $\bm{\pi}\in\mathfrak{S}_{n}$ (such that $S=\bm{\pi}\left(1:k\right)$), (b) $\mathbf{L}\in\mathbb{R}^{n\times k}$,
2:(a) $\mathbf{W}\in\mathbb{R}^{n\times k}$
[…]
4:$\mathbf{W}\leftarrow\textbf{0}_{n\times k}$, $\mathbf{W}\left(\bm{\pi}\left(1:k\right),:\right)=\mathbf{I}_{k}$
[…]
Proposition 5.1.
The sequential (Algorithm 1) and blockwise (Algorithm 4) random pivoting, as well as their greedy pivoting variations, are exactly ID-revealing (Definition 1.2).
Proof 5.2 (Proof of Proposition 5.1).
We first observe that the difference between sequential/blockwise random versus greedy pivoting lies only in the skeleton subset $S$. That is, assuming the same skeleton selection $S$, the resulting matrices $\mathbf{L}$ from random and greedy pivoting would be the same. Therefore, it suffices to show that Proposition 5.1 holds for random pivoting in Algorithms 1 and 4.
For both Algorithms 1 and 4, recall that $\mathbf{L}\leftarrow\mathbf{L}^{(t)}$. We observe that $\mathbf{L}\mathbf{Q}^{(t)\top}$ provides a rank-$k$ approximation for $\mathbf{X}$: $\left\|\mathbf{X}-\mathbf{L}\mathbf{Q}^{(t)\top}\right\|_{F}^{2}=\left\|\mathbf{d}^{(t)}\right\|_{1}=\mathcal{E}_{\mathbf{X}}\left(S\right)<\tau\left\|\mathbf{X}\right\|_{F}^{2}$. We will show that $\mathbf{L}\mathbf{Q}^{(t)\top}=\mathbf{X}\mathbf{X}_{S}^{\dagger}\mathbf{X}_{S}$ in exact arithmetic.
Then, with $\mathbf{L}_{1}\leftarrow\mathbf{L}\left(\bm{\pi}\left(1:k\right),:\right)$ and $\mathbf{L}_{2}\leftarrow\mathbf{L}\left(\bm{\pi}\left(k+1:n\right),:\right)$ as in Algorithm 5, by constructing, $\mathbf{L}_{2}\mathbf{Q}^{(t)\top}=\mathbf{X}\left(\bm{\pi}\left(k+1:n\right),:\right)$ and $\mathbf{L}_{1}\mathbf{Q}^{(t)\top}=\mathbf{X}_{S}=\mathbf{X}\left(\bm{\pi}\left(1:k\right),:\right)$ where $\mathbf{L}_{1}\in\mathbb{R}^{k\times k}$ is invertible as $\mathbf{X}_{S}$ consists of linearly independent rows. Therefore, $\mathbf{X}_{S}^{\dagger}=\mathbf{Q}^{(t)}\mathbf{L}_{1}^{-1}$, and the optimal interpolation matrix $\mathbf{W}=\mathbf{X}\mathbf{X}_{S}^{\dagger}$ can be expressed up to permutation as
$\mathbf{W}\left(\bm{\pi},:\right)=\mathbf{X}\left(\bm{\pi},:\right)\mathbf{X}_{S}^{\dagger}=\begin{bmatrix}\mathbf{L}_{1}\mathbf{Q}^{(t)\top}\\ \mathbf{L}_{2}\mathbf{Q}^{(t)\top}\end{bmatrix}\mathbf{Q}^{(t)}\mathbf{L}_{1}^{-1}=\begin{bmatrix}\mathbf{L}_{1}\\ \mathbf{L}_{2}\end{bmatrix}\mathbf{L}_{1}^{-1}=\begin{bmatrix}\mathbf{I}_{k}\\ \mathbf{L}_{2}\mathbf{L}_{1}^{-1}\end{bmatrix}.$
The computation of $\mathbf{L}_{2}\mathbf{L}_{1}^{-1}$ involves solving a small $k\times k$ system $(n-k)$ times, which takes $O(nk^{2})$ time in general. Overall, given $\mathbf{L}$ from Algorithms 1 and 4, we have that Algorithm 5 constructs the optimal interpolation matrix $\mathbf{W}=\mathbf{X}\mathbf{X}_{S}^{\dagger}=\mathbf{L}\mathbf{L}_{1}^{-1}$.
Moreover, when $\mathbf{L}_{1}$ is well-conditioned, the special lower triangular structure of $\mathbf{L}_{1}$ can be leveraged to further accelerate the construction of $\mathbf{W}$, i.e. $\mathbf{L}_{2}\mathbf{L}_{1}^{-1}$ can be evaluated via backward substitution in $O\left((n-k)k^{2}\right)$ time. However in practice, $\mathbf{L}_{1}$ from the blockwise algorithm Algorithm 4 with a reasonably large block size tends to be ill-conditioned, as elaborated in Remark 5.3.
Remark 5.3 (Stable evaluation of $\mathbf{L}_{2}\mathbf{L}_{1}^{-1}$ for (R)BRP).
In Algorithm 4, the resulting $\mathbf{L}_{1}=\mathbf{L}\left(\bm{\pi}\left(1:k\right),:\right)$ tends to be ill-conditioned, especially with a large block size $b$. For numerical stability, instead of using backward substitution, we evaluate $\mathbf{L}_{1}^{-1}$ via truncated SVD with small singular value clipping (i.e. truncating singular values below a given tolerance, e.g. $10^{-12}$ in our experiments, and their associated singular vectors), which takes $O\left(nk^{2}+k^{3}\right)$ time. We emphasize that compared to backward substitution, computing the pseudoinverse via truncated SVD of the small $k\times k$ matrix $\mathbf{L}_{1}$ has negligible additional cost in the common low-rank approximation setting with $k\ll n$.
5.2 Approximately ID-revealing Algorithms
In contrast to Algorithms 1 and 4, $\mathbf{L}$ from the sketchy pivoting algorithms in Algorithm 3 is insufficient for computing the optimal interpolation matrix $\mathbf{X}\mathbf{X}_{S}^{\dagger}$ exactly, intuitively due to the loss of information in sketching (i.e. randomized dimension reduction).
Proposition 5.4.
The sketchy pivoting algorithms (Algorithm 3) are approximately ID-revealing.
Proposition 5.4 follows directly from the observation that applying Algorithm 5 on $\mathbf{L}$ from Algorithm 3 (i.e. $\mathbf{W}=\mathbf{L}\mathbf{L}_{1}^{-1}$) solves a sketched version of Equation 8:
(9) $\min_{\mathbf{W}\in\mathbb{R}^{n\times k}}\left\|\mathbf{X}\widehat{\bm{\Omega}}-\mathbf{W}\mathbf{X}_{S}\widehat{\bm{\Omega}}\right\|_{F}^{2}\quad\Rightarrow\quad\mathbf{W}=\mathbf{L}\mathbf{L}_{1}^{-1}=\mathbf{X}\widehat{\bm{\Omega}}\left(\mathbf{X}_{S}\widehat{\bm{\Omega}}\right)^{\dagger},$
where $\widehat{\bm{\Omega}}=\bm{\Omega}\left(:,1:k\right)\in\mathbb{R}^{d\times k}$ consists of the first $k$ columns in the randomized linear embedding $\bm{\Omega}\in\mathbb{R}^{d\times l}$ drawn in Algorithm 3. Unfortunately, such sketched least square estimation [49, 42, 15] is known to be suboptimal as long as $l<d$ [40].
Specifically for the interpolation matrix construction, the sketched estimation with a sample size as small as $k$ (i.e. without oversampling) in Algorithm 3 leads to large interpolation error $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)$, as numerically illustrated in Section 6.2 (e.g. comparing $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)$ as SkCPQR/SkLUPP-ID in Figure 2 and $\mathcal{E}_{\mathbf{X}}\left(S\right)$ as SkCPQR/SkLUPP in Figure 1 (right)).
Algorithm 6 Oversampled sketchy interpolation matrix construction (OSID)
1:(a) $\bm{\pi}\in\mathfrak{S}_{n}$ ($S=\bm{\pi}\left(1:k\right)$), (b) Sample size $l\geq k$ (or pre-computed $\mathbf{Y}\in\mathbb{R}^{n\times l}$)
2:(a) $\mathbf{W}\in\mathbb{R}^{n\times k}$
3:(Draw a randomized linear embedding $\bm{\Omega}\in\mathbb{R}^{d\times l}$)
4:($\mathbf{Y}\leftarrow\mathbf{X}\bm{\Omega}\in\mathbb{R}^{n\times l}$)
[…]
6:$\mathbf{W}\leftarrow\textbf{0}_{n\times k}$, $\mathbf{W}\left(\bm{\pi}\left(1:k\right),:\right)=\mathbf{I}_{k}$
[…]
As an intuitive but effective remedy, the oversampled sketchy ID (OSID) in Algorithm 6 generalizes Algorithm 5 by increasing the sample size beyond $k$ (i.e. oversampling), which is shown to considerably alleviate such suboptimality of Equation 9 in interpolation error $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)$ (cf. Section 6.2, SkCPQR/SkLUPP-ID versus -OSID in Figures 2, 3, 4, 5, and 6). Concretely, with a moderate multiplicative oversampling $l=O(k)$ (e.g. $l=2k$ in Figures 2, 3, 4, 5, and 6 is usually sufficient in practice), Algorithm 6 can be expressed as
(10) $\text{OSID}:\qquad\begin{split}&\underset{k\times l}{\mathbf{Y}_{S}}=\mathbf{Y}\left(S,:\right),\qquad\underset{l\times k}{\mathbf{Y}_{S}^{\top}}=\underset{l\times k}{\mathbf{Q}}\ \underset{k\times k}{\mathbf{R}}\\ &\mathbf{W}=\mathbf{Y}\mathbf{Y}_{S}^{\dagger}=\left(\mathbf{Y}\mathbf{Q}\right)\mathbf{R}^{-\top}=\mathbf{X}\left(\bm{\Omega}\bm{\Omega}^{\dagger}\right)\mathbf{X}_{S}^{\dagger}\end{split}$
where $\mathbb{E}_{\bm{\Omega}}\left[\mathbf{W}\right]=\mathbf{X}\mathbf{X}_{S}^{\dagger}$ is an unbiased estimate for the optimal interpolation matrix when the embedding $\bm{\Omega}$ is isotropic, i.e., $\mathbb{E}_{\bm{\Omega}}\left[\bm{\Omega}\bm{\Omega}^{\dagger}\right]=\mathbf{I}_{d}$.
Leveraging the existing theories for randomized linear embedding, [49, Theorem 9 and 23] implies that with Equation 10, the gap between the interpolation and skeletonization errors converges as $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)-\mathcal{E}_{\mathbf{X}}\left(S\right)=O\left({k^{2}}/{l}\right)$ when $\bm{\Omega}$ is a sparse embedding and $l\gg k^{2}$, whereas such convergence can be further improved to $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)-\mathcal{E}_{\mathbf{X}}\left(S\right)=O\left({k}/{l}\right)$ with Gaussian embeddings (e.g. by [49, Theorem 6]).
Asymptotically, Algorithm 6 takes $O(nlk+nk^{2}+lk^{2})=O(nk^{2}+k^{3})$ time. Compared to Algorithm 5 without oversampling, the dominant cost $O(nk^{2})$ increases proportionally to the oversampling factor $l/k$. Nevertheless, such additional cost is affordable in practice (cf. Figures 2, 3, 4, 5, and 6 (right)) thanks to the parallelizability of matrix-matrix multiplications [23].
6 Experiments
6.1 Data Matrices
In the experiments, we compare the accuracy and efficiency of various ID algorithms on several types of data matrices, including the matrix forms of natural image datasets and synthetic random matrices with varied spectra, outlined as follows.
  1. 1.
    We draw a synthetic random data matrix $\mathbf{X}\in\mathbb{R}^{2000\times 500}$ from the adversarial GMM described in Example 4.1 with $k=100$ cluster centers, denoted as GMM.
  2. 2.
    Recall that the MNIST testing set [32] consists of $10,000$ images of hand-written digits from 00 to $9$. We denote MNIST as a data matrix consisting of $n=1000$ random images sampled uniformly from the MNIST testing set where each row contains a flattened and normalized $28\times 28$ image such that $d=784$. The nonzero entries take approximately $20\%$ of the matrix.
  3. 3.
    The CIFAR-10 dataset [30] consists of $60,000$ colored images of size $32\times 32\times 3$. We denote CIFAR-10 as a data matrix consisting of $n=1000$ random images sampled uniformly from the CIFAR-10 dataset where each row contains a flattened and normalized image such that $d=3072$.
  4. 4.
    Let Gaussian-exp be a random dense data matrix of size $1000\times 1000$ with exponential spectral decay. We construct Gaussian-exp from its SVD $\mathbf{X}=\mathbf{U}\bm{\Sigma}\mathbf{V}^{\top}$ where $\mathbf{U},\mathbf{V}\in\mathbb{R}^{1000\times 1000}$ are random unitary matrices drawn from the Haar measure, and $\bm{\Sigma}=\mathop{\mathrm{diag}}\left(\sigma_{1},\cdots,\sigma_{1000}\right)$ where $\sigma_{i}=1$ for all $i\leq 100$ and $\sigma_{i}=\max\left\{0.8^{i-100},10^{-5}\right\}$ for all $i>100$.
  5. 5.
    Let SNN be a random sparse non-negative (SNN) matrix [44] of size $1000\times 1000$ such that
    (11) $\mathbf{X}=\sum_{i=1}^{100}\frac{10}{i}\mathbf{u}_{i}\mathbf{v}_{i}^{T}+\sum_{i=101}^{1000}\frac{1}{i}\mathbf{u}_{i}\mathbf{v}_{i}^{T}$
    where $\mathbf{U}=\left[\mathbf{u}_{1},\cdots,\mathbf{u}_{1000}\right]$ and $\mathbf{V}=\left[\mathbf{v}_{1},\cdots,\mathbf{v}_{1000}\right]$ are $1000\times 1000$ random sparse matrices with non-negative entries and sparsity $0.1$.
6.2 Accuracy and Efficiency of ID
In Figures 2, 3, 4, 5, and 6, we compare the empirical accuracy (left) and efficiency (right) of different ID algorithms on data matrices described in Section 6.1, with a focus on those demonstrating robust skeleton complexities on the adversarial input Example 4.1 (i.e. algorithms in Figure 1 (right)).
Figure 2: Relative interpolation error and runtime of ID algorithms with robust skeleton complexities on the GMM adversarial input (Example 4.1). Recall from Figure 1 that sampling methods, especially squared-norm sampling, suffer from significantly higher skeleton complexities than sequential/robust blockwise random pivoting and sketchy pivoting on GMM. For demonstration, we omit squared-norm sampling from the plot.
Taking an integrated view of ID including skeleton selection and interpolation matrix construction, we consider the following relative ID error777 All error measurements are normalized with $\left\|\mathbf{X}\right\|_{F}^{2}$ to relative errors. and runtime measurements:
  1. (a)
    For error-reveling algorithms (i.e. except for sketchy pivoting and sampling)888 In Figures 2, 3, 4, 5, and 6, we also omit the residual errors of the sequential error-reveling algorithms SRP and CPQR (which overlap with the corresponding interpolation errors) for conciseness of demonstration. , we plot the residual errors (denoted as residual) from their on-the-fly skeletonization error estimates and the runtimes for skeleton selection (e.g. $\left\|\mathbf{d}^{(t)}\right\|_{1}$ in Algorithm 4) only.
  2. (b)
    For exactly ID-revealing algorithms (i.e. except for sketchy pivoting and sampling), we plot the runtimes and interpolation errors $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)$ with $\mathbf{W}$ from Algorithm 5 (ID), which are equal to the corresponding skeletonization errors $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)=\mathcal{E}_{\mathbf{X}}\left(S\right)$.
  3. (c)
    For approximately ID-revealing algorithms (i.e. sketchy pivoting), we plot the runtimes and interpolation errors $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)$ with $\mathbf{W}$ from both Algorithm 5 (ID) and Algorithm 6 (OSID) with small multiplicative oversampling $l=2k$, where both $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)>\mathcal{E}_{\mathbf{X}}\left(S\right)$, and Algorithm 6 tends to be more expensive.
  4. (d)
    For non-ID-revealing algorithms (i.e. sampling), without explicitly specifying in the legends, we plot the runtimes and interpolation errors $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)=\mathcal{E}_{\mathbf{X}}\left(S\right)$ with the optimal $\mathbf{W}$ from solving the exact least square problem in Equation 8.
The legends in Figures 2, 3, 4, 5, and 6 are formatted in terms of “skeleton selection algorithm”-“ID error”. The abbreviations of different skeleton selection algorithms are summarized in Table 2, whereas the ID error measurements are described as above.
Figure 3: Relative interpolation error and runtime of ID algorithms on MNIST.
Figure 4: Relative interpolation error and runtime of ID algorithms on CIFAR-10.
Remark 6.1 (Choice of block size).
The choice of block size $b$ controls the trade-off between parallelizability and error-revealing property. To be precise, an unsuitably small $b$ lessens the efficiency gained from moving flops to matrix-matrix multiplications, whereas the larger $b$ leads to higher overestimation of the target rank. In the experiments, we use $b=30$ or $40$ to exploit the parallelizability of matrix-matrix multiplications without considerably compromising the error-revealing property.
Figure 5: Relative interpolation error and runtime of ID algorithms on Gaussian-exp.
Figure 6: Relative interpolation error and runtime of ID algorithms on SNN.
For the adversarial GMM dataset in Figure 2, it is worth mentioning that the sharp transition of RBGP at $\left|S\right|\approx 100$ is because GMM contains $100$ clusters, each with distinct squared-norms. Recall that blockwise greedy pivoting picks points with the maximum norms without randomness. When $\left|S\right|\leq 100$, each block drains points from the same clusters with the maximum residual norms first, most of which are redundant points and are later excluded by the robust blockwise filtering. This leads to the inefficiency of RBGP when $\left|S\right|\leq 100$. Such inefficiency of RBGP in comparison to RBRP again illustrates the superiority of random pivoting over greedy pivoting, i.e. the merit of randomness.
Table 3: GPU runtime of the relatively efficient/accurate ID algorithms in Figure 2 on the synthetic adversarial GMM dataset (Example 4.1) of size $10^{5}\times 10^{3}$, with $k=100$ cluster centers. Experiments were performed on an Nvidia V100 GPU. Our code relied on the magma_dgeqp3_gpu routine from the MAGMA library  [17] for CPQR and the cuBLAS library [38] for other matrix operations.
rank 52 100 220 346 472
CPQR-ID 7.739e-02 1.447e-01 3.023e-01 4.439e-01 5.611e-01
SkCPQR-OSID 2.394e-02 5.033e-02 1.173e-01 2.134e-01 3.297e-01
SkLUPP-OSID 1.204e-02 2.233e-02 4.835e-02 8.254e-02 1.189e-01
RBRP-ID 4.448e-02 7.231e-02 1.073e-01 1.336e-01 1.596e-01
Further in Table 3, we compare the GPU runtimes of the high-performance implementation for RBRP on the GMM dataset with the other relatively efficient/accurate ID algorithms in Figure 2, including (i) CPQR-ID with one of the best skeleton complexities (as RBRP-ID) and (ii) SkCPQR/SkLUPP-OSID with slightly compromised skeleton complexities (compared to RBRP-ID) but the best CPU runtimes. We observe that RBRP-ID is consistently more efficient than CPQR-ID across all ranks. Moreover, as the rank increases, the efficiency of RBRP-ID approaches that of SkCPQR/SkLUPP-OSID, in addition to better accuracy.
From experiments on the (non-adversarial) natural and synthetic datasets in Figures 3, 4, 5, and 6, we observe the following (by algorithms):
  • The robust blockwise random pivoting (RBRP) enjoys (one of) the best skeleton complexities, with the skeletonization error estimates aligning perfectly with the interpolation errors (left), demonstrating the error-revealing and exactly ID-revealing abilities of RBRP. The skeleton selection process of RBRP tends to take slightly longer than that of sketchy pivoting (SkLUPP/SkCPQR) with ID (which is not error-revealing and suffers from much larger interpolation errors than RBRP). However, RBRP is generally faster than sketchy pivoting (SkLUPP/SkCPQR) with small multiplicative oversampling (OSID), while enjoying lower interpolation error. Overall, RBRP is shown to provide the best combination of accuracy and efficiency when the explicit construction of ID (in contrast to the skeleton subset only) is inquired.
  • The sketchy pivoting algorithms (SkLUPP/SkCPQR) also enjoy (one of) the best skeleton complexities, despite the lack of error-revealing ability and being approximately ID-revealing (i.e. suffering from large interpolation errors). That is, for skeleton selection only (without asking for the interpolation matrix), sketchy pivoting provides the most efficient selection of close-to-the-best skeleton subsets in practice (cf. Figure 1). Moreover, our experiments verify the conclusion in [16] that SkLUPP tends to be more efficient than SkCPQR in selecting skeletons without compromising the skeletonization error.
  • The inherently sequential algorithms, including SRP and CPQR, are highly competitive in terms of accuracy while being considerably slower than the parallelizable ones.
[…]
7 Discussion
In this work, we focus on fast and accurate ID algorithms from five perspectives that measure the empirical performance systematically: (a) skeleton complexity, (b) asymptotic complexity, (c) parallelizability, (d) error-revealing property, and (e) ID-revealing property. With a careful exploration of various existing ID algorithms through the lens of adaptiveness and/or randomness, we reemphasize the effectiveness of combining adaptiveness and randomness in solving ID problems, while unveiling a critical missing piece in the family of existing algorithms that achieves all five of the aforementioned desired properties. To close the gap, we propose robust blockwise random pivoting (RBRP)—a parallelizable, error-revealing, and exactly ID-revealing algorithm that enjoys among-the-best skeleton and asymptotic complexities in practice compared to the existing algorithms and demonstrates promising robustness to adversarial inputs. Through numerical experiments, we illustrate the empirical accuracy, efficiency, and robustness of RBRP on a broad spectrum of natural and synthetic datasets.
Acknowledgments
The authors wish to thank Yifan Chen, Ethan Epperly, Kevin Miller, Chris-topher Musco, Rachel Ward, and Robert Webber for enlightening discussions.

### v5

@@
 The interpolative decomposition (ID) aims to construct a low-rank approximation formed by a basis consisting of row/column skeletons in the original matrix and a corresponding interpolation matrix.
-This work explores fast and accurate ID algorithms from five essential perspectives for empirical performance: (a) skeleton complexity that measures the minimum possible ID rank for a given low-rank approximation error, (b) asymptotic complexity in FLOPs, (c) parallelizability of the computational bottleneck as matrix-matrix multiplications, (d) error-revealing property that enables automatic rank detection for given error tolerances without prior knowledge of target ranks, (e) ID-revealing property that ensures efficient construction of the optimal interpolation matrix after selecting the skeletons.
-While a broad spectrum of algorithms have been developed to optimize parts of the aforementioned perspectives, practical ID algorithms proficient in all perspectives remain absent.
-To fill in the gap, we introduce robust blockwise random pivoting (RBRP) that is parallelizable, error-revealing, and exactly ID-revealing, with comparable skeleton and asymptotic complexities to the best existing ID algorithms in practice.
-Through extensive numerical experiments on various synthetic and natural datasets, we demonstrate the appealing empirical performance of RBRP from the five perspectives above, as well as the robustness of RBRP to adversarial inputs.
+This work explores fast and accurate ID algorithms from comprehensive perspectives for empirical performance, including accuracy in both skeleton selection and interpolation matrix construction, efficiency in terms of asymptotic complexity and hardware efficiency, as well as rank adaptiveness.
+While many algorithms have been developed to optimize some of these aspects, practical ID algorithms proficient in all aspects remain absent.
+To fill in the gap, we introduce robust blockwise random pivoting (RBRP) that is asymptotically fast, hardware-efficient, and rank-adaptive, providing accurate skeletons and interpolation matrices comparable to the best existing ID algorithms in practice.
+Through extensive numerical experiments on various synthetic and natural datasets, we demonstrate the appealing empirical performance of RBRP from the aforementioned perspectives, as well as the robustness of RBRP to adversarial inputs.
 keywords:
-Randomized Numerical Linear Algebra, Interpolative Decomposition, Column Subset Selection, Adaptive Sampling, Parallelization
+Randomized Numerical Linear Algebra, Interpolative Decomposition, Column Subset Selection, Adaptive Sampling
 {MSCcodes}
@@
 1 Introduction
-The interpolative decomposition (ID) is a special type of low-rank approximation where a basis for the row/column space is formed by subselecting rows/columns from the target matrix.
-This allows the ID to capture the essential structures and relationships within the data.
-ID has found applications in various fields, including numerical analysis, data compression, and machine learning, where preserving the integrity of the original data is of paramount importance.
-Given a data matrix $\mathbf{X}=\left[\mathbf{x}_{1},\cdots,\mathbf{x}_{n}\right]^{T}\in\mathbb{R}^{n\times d}$ consisting of $n$ data points $\left\{\mathbf{x}_{i}\in\mathbb{R}^{d}\right\}_{i\in[n]}$, along with a constant $\epsilon>0$ and a target rank $r\in\mathbb{N}$, we aim to construct an $(r,\epsilon)$-ID of $\mathbf{X}$:
-(1) $\underset{n\times d}{\mathbf{X}}\approx\underset{n\times k}{\mathbf{W}}\ \underset{k\times d}{\mathbf{X}_{S}}\quad\emph{s.t.}\quad\left\|\mathbf{X}-\mathbf{W}\mathbf{X}_{S}\right\|_{F}^{2}\leq(1+\epsilon)\left\|\mathbf{X}-{\mathbf{X}}_{\langle r\rangle}\right\|_{F}^{2}$
-where (i) $k$is the (unknown) target rank; (ii) ${\mathbf{X}}_{\langle r\rangle}$is the optimal rank-$r$ approximation of $\mathbf{X}$ (given by the singular value decomposition (SVD)); (iii) $S=\left\{s_{1},\cdots,s_{k}\right\}\subset[n]$are the skeleton indices corresponding to the row skeleton submatrix $\mathbf{X}_{S}=[\mathbf{x}_{s_{1}},\cdots,\mathbf{x}_{s_{k}}]^{\top}\in\mathbb{R}^{k\times d}$; and (iv) $\mathbf{W}\in\mathbb{R}^{n\times k}$is an interpolation matrix with (nearly) optimal error $\left\|\mathbf{X}-\mathbf{W}\mathbf{X}_{S}\right\|_{F}^{2}$ for the given $\mathbf{X}_{S}$.
+The interpolative decomposition (ID) is a special type of low-rank approximation where a basis for the row/column space of a given target matrix is formed by subselecting rows/columns.
+This allows an ID approximation to identify structure and essential information within data.
+The ID has found a wide range of applications in numerical analysis, data compression, and machine learning.
+For example, in fast direct solvers for linear elliptic partial differential equations (PDEs), the ID is a widely used method for compressing certain off-diagonal blocks of rank-structured matrices [51, 32, 39, 40, 13, 53].
+Another emerging application of the ID lies in the low-rank adaptation (LoRA) [42] for parameter-efficient fine-tuning of large language models [41, 47] where ID of weight matrices can be leveraged as good initialization for the low-rank adapter matrices [63].
+More broadly, the idea of ID naturally connects to various related problems like CUR decomposition, kernel quadrature, optimal experimental design, data selection, active learning, and model pruning [49, 29, 57, 3, 15, 24, 4, 48].
+Given $\mathbf{X}=\left[\mathbf{x}_{1},\cdots,\mathbf{x}_{n}\right]^{T}\in\mathbb{R}^{n\times d}$ consisting of $n$ data points $\left\{\mathbf{x}_{i}\in\mathbb{R}^{d}\right\}_{i\in[n]}$, along with a constant $\epsilon>0$ and a target rank $r\in\mathbb{N}$, we aim to construct a $(r,\epsilon)$-ID of $\mathbf{X}$:
+(1) $\underset{n\times d}{\mathbf{X}}\approx\underset{n\times k}{\mathbf{W}}\ \underset{k\times d}{\mathbf{X}_{S}}\quad\emph{s.t.}\quad\left\|\mathbf{X}-\mathbf{W}\mathbf{X}_{S}\right\|_{F}^{2}\leqslant(1+\epsilon)\left\|\mathbf{X}-{\langle\mathbf{X}\rangle}_{r}\right\|_{F}^{2}$
+where $k$ is the (unknown) target rank; ${\langle\mathbf{X}\rangle}_{r}$ is the optimal rank-$r$ approximation of $\mathbf{X}$; $S=\left\{s_{1},\cdots,s_{k}\right\}\subset[n]$ are the skeleton indices corresponding to the row skeleton submatrix $\mathbf{X}_{S}=[\mathbf{x}_{s_{1}},\cdots,\mathbf{x}_{s_{k}}]^{\top}\in\mathbb{R}^{k\times d}$; and $\mathbf{W}\in\mathbb{R}^{n\times k}$ is an interpolation matrix that (nearly) minimizes the error $\left\|\mathbf{X}-\mathbf{W}\mathbf{X}_{S}\right\|_{F}^{2}$ for the given $\mathbf{X}_{S}$.
+While we focus on row-wise ID throughout this work, the column-wise ID can be defined and conducted analogously by considering the transpose of $\mathbf{X}$.
 The construction of an ID can be divided into two stages:
@@
   2. (ii)
-    Given $S$, compute an interpolation matrix $\mathbf{W}\in\mathbb{R}^{n\times\left|S\right|}$ efficiently111 It is worth clarifying that although Equation 2 directly provides $\mathbf{W}=\mathbf{X}\mathbf{X}_{S}^{\dagger}$ as the minimizer of $\min_{\mathbf{W}}\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)=\mathcal{E}_{\mathbf{X}}\left(S\right)$, the explicit computation of $\mathbf{X}\mathbf{X}_{S}^{\dagger}$ takes $O\left(ndk+dk^{2}+nk^{2}\right)$ time (e.g.
-via Equation 8), where the dominant cost $O(ndk)$ is undesired.
-Therefore, whenever available, information from the skeleton selection process is usually leveraged to evaluate/approximate $\mathbf{W}$ efficiently in $o(ndk)$ time (as elaborated in Section 5), which may lead to the gap $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)-\mathcal{E}_{\mathbf{X}}\left(S\right)>0$.
+    Given $S$, compute an interpolation matrix $\mathbf{W}\in\mathbb{R}^{n\times\left|S\right|}$ efficiently111 Although (2) directly provides $\mathbf{W}=\mathbf{X}\mathbf{X}_{S}^{\dagger}$ as the minimizer of $\min_{\mathbf{W}}\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)=\mathcal{E}_{\mathbf{X}}\left(S\right)$, the explicit computation of $\mathbf{X}\mathbf{X}_{S}^{\dagger}$ takes $O(ndk)$ time (e.g.
+via (6)), which is undesired.
+Therefore, whenever available, information from the skeleton selection process is usually leveraged to evaluate/approximate $\mathbf{W}$ efficiently in $o(ndk)$ time (as elaborated in Section 5), which may lead to the gap $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)>\mathcal{E}_{\mathbf{X}}\left(S\right)$.
 that (approximately) minimizes the interpolation error: $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)\triangleq\left\|\mathbf{X}-\mathbf{W}\mathbf{X}_{S}\right\|_{F}^{2}$.
-In this work, we explore fast and accurate ID algorithms from the perspectives of both skeleton subset selection and interpolation matrix construction, and we start by posing the main question:
-How do we construct an $(r,\epsilon)$-ID efficiently with as small a skeleton $S$ as possible,
+In this work, we explore fast and accurate ID algorithms from both the skeleton subset selection and interpolation matrix construction perspectives, starting with the question:
+How do we construct a $(r,\epsilon)$-ID efficiently with as small a skeleton $S$ as possible,
 without prior knowledge of the target rank $k$?
@@
   1. (a)
-    Skeleton complexity measures the minimum size of the skeleton set (i.e.
-the minimum possible rank of ID) that an algorithm must select before $\mathcal{E}_{\mathbf{X}}\left(S\right)\leq(1+\epsilon)\left\|\mathbf{X}-{\mathbf{X}}_{\langle r\rangle}\right\|_{F}^{2}$ (i.e.
-where there exists an $(r,\epsilon)$-ID associated with $S$).
+    Skeleton complexity measures the utility of skeletons selected by an ID algorithm in terms of the minimum possible rank of a $(r,\epsilon)$-ID for any given $\mathbf{X}$.
+Precisely, the skeleton complexity is a property of an ID algorithm defined as the minimum rank $k$ such that for any $\mathbf{X}$, $\left|S\right|=k$ is sufficient to guarantee either a $(r,\epsilon)$-ID for a deterministic algorithm or $\mathbb{E}[\mathcal{E}_{\mathbf{X}}\left(S\right)]\leqslant(1+\epsilon)\left\|\mathbf{X}-{\langle\mathbf{X}\rangle}_{r}\right\|_{F}^{2}$ for a randomized algorithm.
   2. (b)
@@
   3. (c)
-    Parallelizability (in the context of this work) refers to the implementation advantage that the dominant cost for skeleton selection in the algorithm appears as matrix-matrix, instead of matrix-vector, multiplications.
+    Hardware efficiency refers to the implementation advantage that dominant cost of skeleton selection in an ID algorithm appears as matrix-matrix, instead of matrix-vector, multiplications.
 On modern processors, Level-3 BLAS, or matrix-matrix, operations can leverage memory cashing much more efficiently than Level-2 or Level-1 BLAS operations [7, 34].
@@
   4. (d)
-    Error-revealing property refers to the ability of an algorithm to evaluate the skeletonization error efficiently on the fly so that the target rank $k$ does not need to be given a priori.
-    Definition 1.1 (Error-revealing).
-    An ID algorithm is error-revealing if after selecting a skeleton subset $S$, it can evaluate $\mathcal{E}_{\mathbf{X}}\left(S\right)$ efficiently with at most $O\left(n\right)$ operations.
-    That is, given any relative error tolerance $\tau\in(0,1)$, an error-revealing ID algorithm can determine an appropriate rank $k$ with negligible additional cost (i.e.
-no more than $O(nk)$) while selecting a skeleton subset $S$ ($\left|S\right|=k$) such that $\mathcal{E}_{\mathbf{X}}\left(S\right)\leq\tau\left\|\mathbf{X}\right\|_{F}^{2}$.
-For the sake of analysis, we define a useful constant associated with $\mathbf{X}$ throughout the paper:
-    (3) $\eta_{r}\triangleq{\left\|\mathbf{X}-{\mathbf{X}}_{\langle r\rangle}\right\|_{F}^{2}}/{\left\|\mathbf{X}\right\|_{F}^{2}},$
+    Rank-adaptiveness refers to the ability to evaluate the skeletonization error efficiently on the fly without requiring prior knowledge of the target rank $k$.
+    Remark 1.1.
+    An ID algorithm is rank-adaptive if after selecting a skeleton subset $S$, it can evaluate $\mathcal{E}_{\mathbf{X}}\left(S\right)$ efficiently with at most $O\left(n\right)$ operations.
+That is, given any relative error tolerance $\tau>0$, a rank-adaptive ID algorithm can determine an appropriate rank $k$ with negligible (i.e.
+no more than $O(nk)$) additional cost, while selecting a skeleton subset $S$ of size $k$ such that $\mathcal{E}_{\mathbf{X}}\left(S\right)\leqslant\tau\left\|\mathbf{X}\right\|_{F}^{2}$.
+    For the sake of analysis, we define a useful constant associated with $\mathbf{X}$:
+    (3) $\eta_{r}\triangleq{\left\|\mathbf{X}-{\langle\mathbf{X}\rangle}_{r}\right\|_{F}^{2}}/{\left\|\mathbf{X}\right\|_{F}^{2}},$
     which quantifies the relative tail weight of $\mathbf{X}$ with respect to the rank $r$ and represents the relative optimal rank-$r$ approximation error of $\mathbf{X}$.
-Notice that when $\tau<(1+\epsilon)\eta_{r}$, the algorithm outputs an $(r,\epsilon)$-ID.
+Notice that when $\tau<(1+\epsilon)\eta_{r}$, the algorithm outputs a $(r,\epsilon)$-ID.
   5. (e)
-    ID-revealing property characterizes whether the skeleton selection stage of an ID algorithm generates sufficient information for efficient interpolation matrix construction.
-    Definition 1.2 (Exactly, approximately, and non-ID-revealing).
-    We say a skeleton selection algorithm is exactly ID-revealing if it contains sufficient information in addition to $S$ such that the optimal interpolation matrix $\mathbf{W}=\mathbf{X}\mathbf{X}_{S}^{\dagger}$ can be evaluated efficiently in $O(nk^{2})$ time.
-Otherwise, if a suboptimal interpolation matrix $\mathbf{W}\approx\mathbf{X}\mathbf{X}_{S}^{\dagger}$ can be constructed in $O(nk^{2})$ time, we say that the skeleton selection algorithm is approximately ID-revealing.
-If neither $\mathbf{X}\mathbf{X}_{S}^{\dagger}$ nor its approximations can be constructed in $O(nk^{2})$ time, the skeleton selection algorithm is non-ID-revealing.
+    ID-revealing property characterizes if the skeleton selection stage of an ID algorithm contains sufficient information for efficient construction of the interpolation matrix.
+    Definition 1.1 (ID-revealing properties).
+    We say a skeleton selection algorithm is
+    • ID-revealing if it contains sufficient information in addition to $S$ such that the optimal interpolation matrix $\mathbf{W}=\mathbf{X}\mathbf{X}_{S}^{\dagger}$ can be evaluated efficiently in $O(nk^{2})$ time;
+    • $\gamma$-ID-revealing for some $\gamma>1$ if it can construct a suboptimal interpolation matrix $\mathbf{W}\approx\mathbf{X}\mathbf{X}_{S}^{\dagger}$ with $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)\leqslant\gamma\mathcal{E}_{\mathbf{X}}\left(S\right)$ in $O(nk^{2})$ time; and
+    • non-ID-revealing if neither the optimal interpolation matrix nor its approximations can be constructed in $O(nk^{2})$ time, i.e., the interpolation matrix can only be constructed by solving the least-squares problem (2) in $O(ndk)$ time.
 1.2 How to Combine Adaptiveness and Randomness for ID?
-Adaptiveness and randomness are two critical algorithmic properties that can be leveraged to characterize a skeleton selection algorithm.
-On the one hand, with adaptiveness, the skeleton selection in each step is aware of the selected skeletons from the previous steps so that redundant skeleton selection can be better avoided.
-On the other hand, randomness diversifies the skeleton selection, thereby improving the robustness of algorithms to scarce adversarial inputs and providing strong statistical guarantees for skeleton complexities.
+Adaptiveness and randomness are two key algorithmic properties for skeleton selection.
+On the one hand, with adaptiveness, skeleton selection in each step is aware of the previous selections so that redundant skeleton selection can be better avoided.
+On the other hand, randomness balances the exploitation with exploration, thereby improving the robustness to adversarial inputs.
 To ground the notions of adaptiveness and randomness, we synopsize some representative existing skeleton selection algorithms below, as well as in Table 1, with a focus on the aforementioned properties for performance measurement.
@@
     Greedy pivoting is a classical strategy in numerical linear algebra that involves only adaptiveness.
-As an example of greedy pivoting, column-pivoted QR (CPQR) [22, Section 5.4.2] picks the point in the residual with the maximum norm in each step and adaptively updates the residual via projection onto the orthogonal complement of the skeletons.
-Despite the error-revealing and exactly ID-revealing abilities, along with the remarkable empirical success, of CPQR, deterministic greedy pivoting algorithms like CPQR are inherently sequential and vulnerable to adversarial inputs where the skeleton complexity approaches the problem size $n$ [29, 5].
+As an example of greedy pivoting, column-pivoted QR (CPQR) [33, Section 5.4.2] starts with the input matrix $\mathbf{X}$ as the active submatrix, picks the row with the maximum norm in each step, and adaptively updates the active submatrix via projection onto the orthogonal complement of the skeletons.
+CPQR is rank-adaptive and ID-revealing, and usually provides low skeleton complexities in practice.
+However, deterministic greedy pivoting algorithms like CPQR are inherently sequential and vulnerable to adversarial inputs where the skeleton complexity can approach the problem size $n$ [43, 8].
   2. (b)
@@
 As a trade-off between the skeleton complexity and efficiency, the fast ($O(nd)$-time) sampling methods like squared-norm sampling tend to suffer from high skeleton complexities that depend heavily on the matrix [20, 8]; whereas constructions of the more sophisticated distributions like leverage score and DPP are usually expensive [27, 6, 16].
-Moreover, for ID, sampling methods generally fail to be error-revealing or ID-revealing.
+Moreover, for ID, sampling methods generally fail to be rank-adaptive or ID-revealing.
   3. (c)
     Random pivoting combines adaptiveness and randomness by replacing the greedy selection of maximum norm in CPQR with random sampling according to the squared-norm distribution associated with the residual [20, 21, 8].
-Closely related to the ID problem considered in this work, the idea of random pivoting has been revitalized by the inspiring recent work [5] in the context of column Nyström approximation with a nearly optimal skeleton complexity guarantee [5, Theorem 3.1].
-However, analogous to CPQR, although random pivoting enjoys the desired error-revealing and exact ID-revealing properties, the sequential nature of random pivoting compromises its efficiency in practice.
+Closely related to the ID problem considered in this work, the idea of random pivoting has been revitalized by the inspiring recent work [8] in the context of column Nyström approximation with a nearly optimal skeleton complexity guarantee [8, Theorem 5.1].
+However, analogous to CPQR, although random pivoting enjoys the desired rank-adaptiveness and ID-revealing properties, the sequential nature of random pivoting compromises its efficiency in practice.
   4. (d)
-    Sketchy pivoting is an alternative combination of adaptiveness and randomness that randomly embeds the row space of $\mathbf{X}$ via sketching [26, 49], on top of which skeletons are selected via greedy pivoting [48, 16, 39].
+    Sketchy pivoting combines adaptiveness and randomness alternatively by first sketching $\mathbf{X}$ from right and then selecting skeletons through greedy pivoting on the sketched matrix [62, 23].
 In contrast to random pivoting, the cost of sketchy pivoting is dominated by the embarrassingly parallelizable sketching process.
-As a trade-off for superior empirical efficiency, sketchy pivoting sacrifices the error-revealing property since sketching requires prior knowledge of the target $k$ or its overestimation.
-Furthermore, sketchy pivoting is approximately ID-revealing due to the loss of information during sketching.
-As a simple effective remedy for the loss of accuracy, we observe that multiplicative oversampling can remarkably improve the gap between the interpolation and skeletonization error $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)-\mathcal{E}_{\mathbf{X}}\left(S\right)$, which we refer to as oversampled sketchy ID (OSID).
+As a trade-off for superior empirical efficiency, sketchy pivoting sacrifices the rank-adaptiveness since sketching requires prior knowledge of the target $k$ or its overestimation.
+Furthermore, sketchy pivoting is not ID-revealing due to the loss of information during sketching.
+As a simple but effective remedy for the loss of accuracy, we observe that multiplicative oversampling can remarkably improve the gap between the interpolation and skeletonization error $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)-\mathcal{E}_{\mathbf{X}}\left(S\right)$, which we refer to as oversampled sketchy ID (OSID).
+In Section 5.2, we show that with sufficiently large oversampling, OSID is $\gamma$-ID-revealing for some small constant $\gamma>1$.
 Table 1: Summary of performance for skeleton selection methods that leverage adaptiveness and/or randomness.
-Recall from Equation 3 that $\eta_{r}\triangleq{\left\|\mathbf{X}-{\mathbf{X}}_{\langle r\rangle}\right\|_{F}^{2}}/{\left\|\mathbf{X}\right\|_{F}^{2}}$.
-The skeleton complexities in $(\cdot)^{*}$ are conjectured based on extensive empirical evidence and intuitive rationale but without formal proofs.
-In the “Dominant Cost” column (i.e.
-showing both the asymptotic complexity and parallelizability), “m-v” stands for matrix-vector multiplications, which are significantly less efficient than “m-m”—matrix-matrix multiplications—due to the lack of parallelizability [23].
-Skeleton Selection Skeleton Complexity Dominant Cost Error-revealing ID-revealing Greedy pivoting/CPQR ([22, Section 5.4.2]) $k\geq\left(1-\left(1+\epsilon\right)\eta_{r}\right)n$ $O(ndk)$ m-v Yes Exactly Squared-norm sampling (Proposition 2.3 [20]) $k\geq\frac{r-1}{\epsilon\eta_{r}}+\frac{1}{\epsilon}$ $O(nd)$ m-m No Non Random pivoting (Algorithm 1 [5, 13, 14]) $k\geq k_{\text{RP}}=\frac{r}{\epsilon}+$ $r\min\left\{\log\left(\frac{1}{\epsilon\eta_{r}}\right),1+\log\left(\frac{2^{r}}{\epsilon}\right)\right\}$ $O(ndk)$ m-v Yes Exactly Sketchy pivoting (Algorithm 3 [48, 16]) $\left(k\gtrsim k_{\text{RP}}\right)^{*}$ $O(ndk)$ m-m No Approximately RBRP (this paper) (Algorithm 4) $\left(k\gtrsim k_{\text{RP}}\right)^{*}$ $O(ndk)$ m-m Yes Exactly
-To fill in a critical missing piece in the family of existing methods in Table 1 that leverage adaptiveness and randomness, in this work, we introduce robust blockwise random pivoting (RBRP, Algorithm 4)—an ID algorithm that is $O(ndk)$-time, parallelizable, error-revealing, and exactly ID-revealing, with comparable skeleton complexities to the nearly optimal random pivoting [5] in practice.
+Recall from (3) that $\eta_{r}\triangleq{\left\|\mathbf{X}-{\langle\mathbf{X}\rangle}_{r}\right\|_{F}^{2}}/{\left\|\mathbf{X}\right\|_{F}^{2}}$.
+For randomized methods, ($\mathbb{E}$) denotes that $\mathbb{E}[\mathcal{E}_{\mathbf{X}}\left(S\right)]\leqslant(1+\epsilon)\left\|\mathbf{X}-{\langle\mathbf{X}\rangle}_{r}\right\|_{F}^{2}$ holds in expectation.
+The skeleton complexities in $(\cdot)^{*}$ are conjectured based on extensive empirical evidence without formal proofs.
+In the “Dominant Cost” column (showing both the asymptotic complexity and hardware efficiency), “m-v” stands for matrix-vector multiplications, which are significantly less efficient than “m-m”—matrix-matrix multiplications [34].
+Skeleton Selection Skeleton Complexity Dominant Cost Rank-adaptive ID-revealing Greedy pivoting/CPQR ([33, Section 5.4.2]) $k\geqslant\left(1-\left(1+\epsilon\right)\eta_{r}\right)n$ $O(ndk)$ m-v Yes Yes Squared-norm sampling ((5[30]) ($\mathbb{E}$) $k\geqslant\frac{r-1}{\epsilon\eta_{r}}+\frac{1}{\epsilon}$ $O(nd)$ m-m No No Random pivoting (Algorithm 1 [8, 20, 21]) ($\mathbb{E}$) $k\geqslant k_{\text{RP}}=\frac{r}{\epsilon}+$ $r\min\left\{\log\left(\frac{1}{\epsilon\eta_{r}}\right),1+\log\left(\frac{2^{r}}{\epsilon}\right)\right\}$ $O(ndk)$ m-v Yes Yes Sketchy pivoting ([23, Algorithm 1][62]) ($\mathbb{E}$) $\left(k\gtrsim k_{\text{RP}}\right)^{*}$ $O(ndk)$ m-m No No RBRP (this paper) (Algorithm 2) ($\mathbb{E}$) $\left(k\gtrsim k_{\text{RP}}\right)^{*}$ $O(ndk)$ m-m Yes Yes
+To fill in a critical missing piece in the family of existing methods in Table 1 that leverage adaptiveness and randomness, in this work, we introduce robust blockwise random pivoting (RBRP, Algorithm 2)—an ID algorithm that is $O(ndk)$-time, hardware-efficient, rank-adaptive, and ID-revealing, with comparable skeleton complexities to the nearly optimal random pivoting [8] in practice.
 In particular, the plain blockwise random pivoting (e.g.
-an extension of the blocked RPCholesky algorithm introduced in [5, Algorithm 2.2]) tends to suffer from unnecessarily large skeleton complexity (cf.
+an extension of the blocked RPCholesky algorithm introduced in [8, Algorithm 3]) tends to suffer from unnecessarily large skeleton complexity (cf.
 Figure 1 (left)) under adversarial inputs (e.g.
@@
 1.3 Notations and Roadmap
-For $\mathbf{X}\in\mathbb{R}^{n\times d}$, let $\mathbf{X}=\mathbf{U}\bm{\Sigma}\mathbf{V}^{\top}=\sum_{i=1}^{\min(n,d)}\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\top}$ be the singular value decomposition of $\mathbf{X}$ with $\sigma_{1}\geq\cdots\geq\sigma_{\min(n,d)}\geq 0$.
-Given any $r\in[\min(n,d)]$, we denote $\mathbf{U}_{r}=\left[\mathbf{u}_{1},\cdots,\mathbf{u}_{r}\right]$, $\bm{\Sigma}_{r}=\mathop{\mathrm{diag}}\left(\sigma_{1},\cdots,\sigma_{r}\right)$, $\mathbf{V}_{r}=\left[\mathbf{v}_{1},\cdots,\mathbf{v}_{r}\right]$ such that ${\mathbf{X}}_{\langle r\rangle}=\mathbf{U}_{r}\bm{\Sigma}_{r}\mathbf{V}_{r}^{\top}$.
+For $\mathbf{X}\in\mathbb{R}^{n\times d}$, let $\mathbf{X}=\mathbf{U}\bm{\Sigma}\mathbf{V}^{\top}=\sum_{i=1}^{\min(n,d)}\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\top}$ be the singular value decomposition of $\mathbf{X}$ with $\sigma_{1}\geqslant\cdots\geqslant\sigma_{\min(n,d)}\geqslant 0$.
+Given any $r\in[\min(n,d)]$, we denote $\mathbf{U}_{r}=\left[\mathbf{u}_{1},\cdots,\mathbf{u}_{r}\right]$, $\bm{\Sigma}_{r}=\mathop{\mathrm{diag}}\left(\sigma_{1},\cdots,\sigma_{r}\right)$, $\mathbf{V}_{r}=\left[\mathbf{v}_{1},\cdots,\mathbf{v}_{r}\right]$ such that ${\langle\mathbf{X}\rangle}_{r}=\mathbf{U}_{r}\bm{\Sigma}_{r}\mathbf{V}_{r}^{\top}$.
 We follow the MATLAB notation for vector and matrix indexing throughout this work.
-For any $n\in\mathbb{N}$, let $[n]=\left\{1,\cdots,n\right\}$ and $\Delta_{n}=\left\{\mathbf{p}\in[0,1]^{n}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \left\|\mathbf{p}\right\|_{1}=1\right\}$ be the probability simplex of dimension $n$.
+For any $n\in\mathbb{N}$, let $[n]=\left\{1,\cdots,n\right\}$ and $\Delta_{n}\in[0,1]^{n}$ be the probability simplex of dimension $n$.
 Also, let $\mathfrak{S}_{n}$ be the set of all permutations of $[n]$.
@@
 For any $m\in\mathbb{N}$, let $\textbf{1}_{m}\in\mathbb{R}^{m}$ and $\textbf{0}_{m}\in\mathbb{R}^{m}$ be the vector with all entries equal to one and zero, respectively.
+We adapt the standard asymptotic notations: for any functions $f,g:\mathbb{R}_{+}\to\mathbb{R}_{+}$, we write $f=O\left(g\right)$ or $f\lesssim g$ if there exists some constant $C>0$ such that $f(x)\leqslant Cg(x)$ for all $x\in\mathbb{R}_{+}$; $f=\Omega\left(g\right)$ or $f\gtrsim g$ if $g=O\left(f\right)$; $f=\Theta(g)$ or $f\asymp g$ if $f=O\left(g\right)$ and $f=\Omega\left(g\right)$.
 As a roadmap, we review the adaptiveness-only and randomness-only skeleton selection algorithms in Section 2, along with the two different combinations of adaptiveness and randomness in Section 3.
 Then, we introduce the RBRP algorithm formally in Section 4.
-In Section 5, we discuss the efficient construction of an interpolation matrix for exactly and approximately ID-revealing algorithms.
+In Section 5, we discuss the efficient construction of interpolation matrices after skeleton selection.
 In Section 6, we demonstrate the appealing empirical accuracy and efficiency of RBRP via extensive numerical experiments.
-We provide the experiment code, as well as a high-performance implementation of RBRP, at https://github.com/dyjdongyijun/Robust_Blockwise_Random_Pivoting.
+The code, as well as a high-performance implementation of RBRP, is available at https://github.com/dyjdongyijun/Robust_Blockwise_Random_Pivoting.
 2 Adaptiveness vs. Randomness
@@
 Greedy pivoting is a classical way of incorporating adaptiveness in skeleton selection [35, 60, 62].
-Column pivoted QR (CPQR) [22, Section 5.4.2] is one of the most commonly used greedy pivoting methods.
+Column pivoted QR (CPQR) [33, Section 5.4.2] is one of the most commonly used greedy pivoting methods.
 The pivoting strategy and adaptive update are two key components that characterize a greedy pivoting method.
-In particular, CPQR involves greedy squared-norm pivoting and adaptive QR updates as synopsized below.
-Starting with $\mathbf{X}^{(0)}\leftarrow\mathbf{X}$, given a(n) (active) data (sub)matrix $\mathbf{X}^{(t)}\in\mathbb{R}^{n\times d}$ with at most $n-t$ nonzero rows, CPQR on $\mathbf{X}^{\top}$ greedily pivots the row in $\mathbf{X}^{(t)}$ with the maximum $\ell_{2}$-norm $s_{t+1}\leftarrow\mathop{\mathrm{argmax}}_{i\in[n]}\left\|\mathbf{X}^{(t)}\left(i,:\right)\right\|_{2}^{2}$ (i.e.
-squared-norm pivoting) and adaptively updates the active submatrix by projecting it onto the orthogonal complement of the selected skeleton222 For illustration, we consider the exact arithmetic here and express the QR update as a Gram-Schmidt process, whereas, in practice, QR updates are usually conducted via Householder reflection [22, Section 5.1.2] for numerical stability.
+In particular, given an active submatrix $\mathbf{X}^{(t)}\in\mathbb{R}^{n\times d}$ with at most $n-t$ nonzero rows, CPQR involves greedy squared-norm pivoting and adaptive QR updates222 Notice that (4) can be viewed as the $(t+1)$th-step of either CPQR on $\mathbf{X}^{\top}$ or row pivoted LQ (RPLQ) on $\mathbf{X}$.
+Throughout this work, we refer to column pivoting on $\mathbf{X}^{\top}$ as CPQR without ambiguity.
+In (4), we consider the exact arithmetic and express the QR update as a Gram-Schmidt process for illustration purposes.
+In practice, QR updates are usually conducted via Householder reflection [33, Section 5.1.2] for numerical stability.
 :
-(4) $\mathbf{X}^{(t+1)}\leftarrow\mathbf{X}^{(t)}-\mathbf{X}^{(t)}\frac{\mathbf{X}^{(t)}(s_{t+1},:)^{\top}\mathbf{X}^{(t)}(s_{t+1},:)}{\left\|\mathbf{X}^{(t)}(s_{t+1},:)\right\|_{2}^{2}}.$
-The rank-revealing property of CPQR [24, Theorem 7.2] implies that the first $k$ pivots $S$ consisting of $\left\{s_{1},\cdots,s_{k}\right\}$ form a reasonable skeleton subset with
-(5) $\mathcal{E}_{\mathbf{X}}\left(S\right)\leq 4^{k}\left(n-k\right)\left\|\mathbf{X}-{\mathbf{X}}_{\langle k\rangle}\right\|_{F}^{2}.$
-The exponential dependence on $k$ in Equation 5 is tight under adversarial inputs (e.g.
-Kahan matrices [29][24, Example 1]).
-In the worst case, CPQR can take almost all of the $n$ points as skeletons before finding an $(r,\epsilon)$-ID.
-Concretely, [5, Theorem 4.1] implies that for CPQR,
-(6) $\mathcal{E}_{\mathbf{X}}\left(S\right)\leq\frac{1}{\eta_{r}}\cdot\left(1-\frac{k}{n}\right)\left\|\mathbf{X}-{\mathbf{X}}_{\langle r\rangle}\right\|_{F}^{2}.$
-That is, CPQR provides an $(r,\epsilon)$-ID when $k\geq\left(1-(1+\epsilon)\eta_{r}\right)n$.
-Moreover, for any $S$ with $\left|S\right|=k\leq\min\left(1-(1+\epsilon)\eta_{r},\eta_{r}\right)n-1$, there exists an $\mathbf{X}$ where CPQR fails to form an $(r,\epsilon)$-ID.
+(4) $s_{t+1}\leftarrow\mathop{\mathrm{argmax}}_{i\in[n]}\left\|\mathbf{X}^{(t)}\left(i,:\right)\right\|_{2}^{2},\ \mathbf{X}^{(t+1)}\leftarrow\mathbf{X}^{(t)}-\mathbf{X}^{(t)}\frac{\mathbf{X}^{(t)}(s_{t+1},:)^{\top}\mathbf{X}^{(t)}(s_{t+1},:)}{\left\|\mathbf{X}^{(t)}(s_{t+1},:)\right\|_{2}^{2}}.$
+The rank-revealing property of CPQR [35, Theorem 7.2] implies that the first $k$ pivots $S=\left\{s_{1},\cdots,s_{k}\right\}$ form a reasonable skeleton subset with $\mathcal{E}_{\mathbf{X}}\left(S\right)\leqslant 4^{k}\left(n-k\right)\left\|\mathbf{X}-{\langle\mathbf{X}\rangle}_{k}\right\|_{F}^{2}$, where the exponential dependence on $k$ is tight under adversarial inputs (e.g.
+Kahan matrices [43][35, Example 1]).
+In the worst case, [8, Theorem C.2] shows that CPQR requires almost all the $n$ points ($k\geqslant\left(1-(1+\epsilon)\eta_{r}\right)n$) before finding a $(r,\epsilon)$-ID.
 Remark 2.1 (Greedy squared-norm pivoting is vulnerable but empirically successful).
@@
 Therefore, CPQR serves as a good skeleton selection method for most data matrices empirically [62, 23], as long as orthonormality is maintained to high precision [10].
-Remark 2.2 (CPQR is inherently sequential but error-revealing).
-From the computational efficiency perspective, a major drawback of CPQR is that the adaptive updates are inherently sequential (only half of the operations can be cast into high-performance matrix-matrix multiplication in the best-known algorithm that underlies LAPACK’s routine geqp3 [41]).
-As a result, CPQR tends to be observably slower than simple alternatives with the same asymptotic complexity like Gaussian elimination with partial pivoting [22, Section 3.4]—another common greedy pivoting method that retains parallelizability at a cost of the rank-revealing guarantee.
-Nevertheless, the adaptive nature grants CPQR an appealing bonus—the error-revealing ability.
+Remark 2.2 (CPQR is inherently sequential but rank-adaptive).
+From the efficiency perspective, a major drawback of CPQR is that the adaptive updates are inherently sequential (only half of the operations can be cast into high-performance matrix-matrix multiplication in the best-known algorithm that underlies LAPACK’s routine geqp3 [58]).
+As a result, CPQR tends to be observably slower than simple alternatives with the same asymptotic complexity like Gaussian elimination with partial pivoting [33, Section 3.4]—another common greedy pivoting method that retains parallelizability at a cost of the rank-revealing guarantee.
+Nevertheless, the adaptive nature grants CPQR an appealing bonus—the rank-adaptiveness.
 It is known that the skeletonization error of CPQR, i.e.
@@
 Sampling is a widely used approach for random skeleton selection (also referred to as column subset selection, e.g.
-in [11, 9]).
-To identify a nearly optimal (smallest possible) skeleton subset $S$ for an $(r,\epsilon)$-ID, the intuitive goal is to construct a distribution $\mathbf{p}\in\Delta_{n}$ over the $n$ data points $\left\{\mathbf{x}_{i}\right\}_{i\in[n]}$ that weights each $\mathbf{x}_{i}$ according to its relative “importance.”
-Analogous to the squared-norm pivoting in CPQR (Remark 2.1), sampling according to the squared-norm distribution is a natural choice for such “importance” sampling.
-Concretely, adapting results from [13] and [5, Theorem 4.2] leads to the following expected skeleton complexity for squared-norm sampling.
-Proposition 2.3 (Squared-norm sampling).
-Consider a size-$k$ skeleton subset $S=\left\{s_{1},\cdots,s_{k}\right\}$ where each $s_{j}\in S$ is sampled from the squared-norm distribution $s_{j}\sim\mathbf{p}=\left(p_{1},\cdots,p_{n}\right)$, $p_{i}=\frac{\left\|\mathbf{x}_{i}\right\|_{2}^{2}}{\left\|\mathbf{X}\right\|_{F}^{2}}$ independently with replacement.
-Then
-$\mathbb{E}_{S\sim\mathbf{p}^{k}}\left[\mathcal{E}_{\mathbf{X}}\left(S\right)\right]\leq\left(1+\frac{r-1}{k\eta_{r}}+\frac{1}{k}\right)\left\|\mathbf{X}-{\mathbf{X}}_{\langle r\rangle}\right\|_{F}^{2}.$
-That is, squared-norm sampling yields an $(r,\epsilon)$-ID when $k\geq\frac{r-1}{\epsilon\eta_{r}}+\frac{1}{\epsilon}$.
-Compared to the skeleton complexity guarantee of squared-norm pivoting (CPQR) in Equation 6, the squared-norm sampling in Proposition 2.3 provides a potentially weaker guarantee in expectation (instead of worst-case) but yields a better skeleton complexity that scales proportionally to $r$ (instead of $n$).
-Despite the improved expected skeleton complexity compared to CPQR, like CPQR (Remark 2.1), squared-norm sampling can also be vulnerable to adversarial inputs by selecting redundant skeletons with higher probabilities.
-Such potential inefficiency can be reflected by the linear dependence of the skeleton complexity $k\geq\frac{r-1}{\epsilon\eta_{r}}+\frac{1}{\epsilon}$ on $\frac{r}{\eta_{r}}$, as instantiated in Example 2.4.
+in [17, 14]), where the intuitive goal is to construct a distribution $\mathbf{p}\in\Delta_{n}$ over the $n$ data points $\left\{\mathbf{x}_{i}\right\}_{i\in[n]}$ according to their relative “importance.” Analogous to the squared-norm pivoting in CPQR ((4)), sampling according to the squared-norm distribution:
+(5) $p_{i}={\left\|\mathbf{x}_{i}\right\|_{2}^{2}}/{\left\|\mathbf{X}\right\|_{F}^{2}}\quad\forall\leavevmode\nobreak\ i\in[n]$
+is a natural choice for such “importance” sampling.
+In particular, [20] and [8, Theorem C.3] shows that squared-norm sampling requires $k\geqslant\frac{r-1}{\epsilon\eta_{r}}+\frac{1}{\epsilon}$ skeleton points to form a $(r,\epsilon)$-ID in expectation.
+Compared to the skeleton complexity of pivoting (CPQR), the squared-norm sampling provides a lower skeleton complexity in expectation that scales proportionally to $r$ instead of $n$.
+Despite the improved expected skeleton complexity compared to CPQR, like CPQR, squared-norm sampling can also be vulnerable to adversarial inputs by selecting redundant skeletons with higher probabilities.
+Such potential inefficiency can be reflected by the linear dependence of the skeleton complexity $k\geqslant\frac{r-1}{\epsilon\eta_{r}}+\frac{1}{\epsilon}$ on $\frac{r}{\eta_{r}}$, as instantiated in Example 2.3.
 Example 2.3 (Adversarial input for squared-norm sampling).
-For $n\in\mathbb{N}$ divisible by $k\in\mathbb{N}$, $r\in\mathbb{N}$ such that $k=(1+\beta)r<n$ for some $\beta\in\mathbb{N}$, letting $\alpha_{i}=\sqrt{\alpha}\gg\beta\geq 1$ for all $1\leq i\leq r$, and $\alpha_{i}=1$ for all $r+1\leq i\leq k$, we consider the following data matrix
+For $n\in\mathbb{N}$ divisible by $k\in\mathbb{N}$, $r\in\mathbb{N}$ such that $k=(1+\beta)r<n$ for some $\beta\in\mathbb{N}$, letting $\alpha_{i}=\sqrt{\alpha}\gg\beta\geqslant 1$ for all $1\leqslant i\leqslant r$, and $\alpha_{i}=1$ for all $r+1\leqslant i\leqslant k$, we consider the following data matrix
 $\mathbf{X}^{\top}=\left[\alpha_{1}\mathbf{e}_{1}\textbf{1}_{n/k}^{\top},\cdots,\alpha_{r}\mathbf{e}_{r}\textbf{1}_{n/k}^{\top},\alpha_{r+1}\mathbf{e}_{r+1}\textbf{1}_{n/k}^{\top},\cdots,\alpha_{k}\mathbf{e}_{k}\textbf{1}_{n/k}^{\top}\right]\in\mathbb{R}^{d\times n}$
@@
 […]
-Such inefficiency is mirrored in the skeleton complexity $k\geq\frac{r-1}{\epsilon\eta_{r}}+\frac{1}{\epsilon}$ with a small $\eta_{r}$:
+Such inefficiency is mirrored in the skeleton complexity $k\geqslant\frac{r-1}{\epsilon\eta_{r}}+\frac{1}{\epsilon}$ with a small $\eta_{r}$:
 $\eta_{r}=\frac{k-r}{\alpha r+k-r}=\frac{\beta}{\alpha+\beta}\ll 1.$
@@
 For instance, after picking $\alpha_{i}\mathbf{e}_{i}$ as the first skeleton, the adaptive update (e.g.
-Equation 4) eliminates the remaining $\frac{n}{k}-1$ rows of $\mathbf{X}$ in the $\mathbf{e}_{i}$ direction and therefore helps exclude the possible redundant samples.
-As the well-studied alternatives to squared-norm sampling, determinantal point process (DPP sampling) [2, 12] and leverage score sampling [33] provide better skeleton complexity guarantees but with higher cost of constructing the corresponding distributions as a trade-off.
-For example, the k-DPP sampling [31] is known for its nearly optimal skeleton complexity in expectation [2, 25, 5]:
-(7) $k=\left|S\right|\geq\frac{r}{\epsilon}+r-1\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ \mathbb{E}_{S\sim\text{k-DPP}(\mathbf{X})}\left[\mathcal{E}_{\mathbf{X}}\left(S\right)\right]\leq(1+\epsilon)\left\|\mathbf{X}-{\mathbf{X}}_{\langle r\rangle}\right\|_{F}^{2}.$
+(4)) eliminates the remaining $\frac{n}{k}-1$ rows of $\mathbf{X}$ in the $\mathbf{e}_{i}$ direction and therefore helps exclude the possible redundant samples.
+Determinantal point process (DPP) [5, 19] and leverage score sampling [49] are well-studied alternatives to squared-norm sampling that provide better skeleton complexities but with much higher cost of constructing the corresponding distributions as a trade-off.
+For example, the k-DPP sampling [45] is known for its nearly optimal skeleton complexity: as few as $k\geqslant\frac{r}{\epsilon}+r-1$ skeletons are sufficient to provide a $(r,\epsilon)$-ID in expectation [5, 36, 8].
 As a trade-off, the classical SVD-based algorithm for k-DPP$(\mathbf{X})$ [6] takes $O\left(nd^{2}\right)$ to construct the distribution and $O\left(nk^{2}\right)$ to draw $k$ samples.
@@
 3 Combining Adaptiveness and Randomness
-We recall that the vulnerability of greedy pivoting (Remark 2.1) comes from the scarce adversarial inputs which can be negligible in expectation under randomization.
+Recall that the vulnerability of greedy pivoting (Remark 2.1) comes from the scarce adversarial inputs.
+Under randomization, the probability that greedy pivoting keeps making adversarial skeleton selections becomes negligible.
 Meanwhile, the vulnerability of squared-norm sampling (Example 2.3) can be effectively circumvented by incorporating adaptive updates.
@@
 Random pivoting [20, 21, 2, 8] is arguably the most intuitive skeleton selection method that combines adaptiveness and randomness.
-It can be simply described as “replacing the greedy squared-norm pivoting in CPQR with squared-norm sampling” or “adaptively updating the dataset via QR after each step of squared-norm sampling,” as formalized in Algorithm 1.
-Algorithm 1 Sequential random pivoting (SRP)
+It can be simply described as “replacing the greedy squared-norm pivoting in CPQR with squared-norm sampling” as formalized in Algorithm 1, with an asymptotic complexity of $O(ndk)$, consisting of $k$ inherently sequential passes through $\mathbf{X}$ (or its submatrices).
+The storage of $\mathbf{L}^{(t)}$ and $\mathbf{Q}^{(t)}$ requires $O(nk)$ and $O(dk)$ memory.
+Analogous to greedy pivoting (Remark 2.2), Algorithm 1 is rank-adaptive thanks to its adaptive nature.
+Algorithm 1 Sequential random pivoting (SRP) [8, Algorithm 7]
 1:(a) Data matrix $\mathbf{X}=\left[\mathbf{x}_{1},\cdots,\mathbf{x}_{n}\right]^{\top}\in\mathbb{R}^{n\times d}$, (b) Relative error tolerance $\tau=\left(1+\epsilon\right)\eta_{r}\in(0,1)$ (or target rank $k\in\mathbb{N}$),
@@
 3:$\mathbf{L}^{(0)}\leftarrow\textbf{0}_{n\times 0}$, $\mathbf{Q}^{(0)}\leftarrow\textbf{0}_{d\times 0}$, $\bm{\pi}\leftarrow[1,\cdots,n]$, $S\leftarrow\emptyset$, $t=0$
-4:$\mathbf{d}^{(0)}\in\mathbb{R}_{\geq 0}^{n}$ with $\mathbf{d}^{(0)}(i)=\left\|\mathbf{x}_{i}\right\|_{2}^{2}\leavevmode\nobreak\ \forall\leavevmode\nobreak\ i\in[n]$triangleright $O(nd)$
+4:$\mathbf{d}^{(0)}\in\mathbb{R}_{\geqslant 0}^{n}$ with $\mathbf{d}^{(0)}(i)=\left\|\mathbf{x}_{i}\right\|_{2}^{2}\leavevmode\nobreak\ \forall\leavevmode\nobreak\ i\in[n]$triangleright $O(nd)$
 5:while $\left\|\mathbf{d}^{(t)}\right\|_{1}<\tau\left\|\mathbf{d}^{(0)}\right\|_{1}$ (or $\left|S\right|<k$do
@@
 15:$k\leftarrow\left|S\right|$, $\mathbf{L}\leftarrow\mathbf{L}^{(t)}$
-Algorithm 1 has an asymptotic complexity of $O\left(ndk+dk^{2}\right)$, with $k$ inherently sequential passes through $\mathbf{X}$ (or its submatrices), while the storage of $\mathbf{L}^{(t)}$ and $\mathbf{Q}^{(t)}$ requires $O(nk)$ and $O(dk)$ memory, respectively.
-Meanwhile, analogous to greedy pivoting (Remark 2.2), Algorithm 1 is error-revealing thanks to its adaptive nature.
-The idea of random pivoting in Algorithm 1 is ubiquitous in various related problems, e.g.
-the combination of volume sampling and adaptive sampling for low-rank approximation [14], the Randomly Pivoted Cholesky (RPCholesky) [5] for kernel column Nyström approximation [35, 19.2], and the $D^{2}$-sampling for k-means++ clustering [1].
-Specifically, applying RPCholesky [5, Algorithm 2.1] to the kernel matrix $\mathbf{X}\mathbf{X}^{\top}\in\mathbb{R}^{n\times n}$ can be reformulated as Algorithm 2, which is equivalent to Algorithm 1 in the exact arithmetic (Remark 3.1).
-Compared to Algorithm 1, the skeleton selection stage of Algorithm 2 has an asymptotic complexity of $O\left(ndk+nk^{2}\right)$ (cf.
-$O\left(ndk+dk^{2}\right)$ for Algorithm 1), while only requiring $O(nk)$ memory for the storage of $\mathbf{L}^{(t)}$.
-Algorithm 2 RPCholesky [5, Algorithm 2.1]
-1:$\mathbf{L}^{(0)}\leftarrow\textbf{0}_{n\times 0}$, $\bm{\pi}\leftarrow[1,\cdots,n]$, $S\leftarrow\emptyset$, $t=0$
-2:$\mathbf{d}^{(0)}\in\mathbb{R}_{\geq 0}^{n}$ with $\mathbf{d}^{(0)}(i)=\left\|\mathbf{x}_{i}\right\|_{2}^{2}\leavevmode\nobreak\ \forall\leavevmode\nobreak\ i\in[n]$triangleright $O(nd)$
-3:while $\left\|\mathbf{d}^{(t)}\right\|_{1}<\tau\left\|\mathbf{d}^{(0)}\right\|_{1}$ (or $\left|S\right|<k$do
-4:     $t\leftarrow t+1$
-5:     $s_{t}\sim\mathbf{d}^{(t-1)}/\left\|\mathbf{d}^{(t-1)}\right\|_{1}$
-6:     $S\leftarrow S\cup\left\{s_{t}\right\}$, $\bm{\pi}\leftarrow\mathtt{swap}\left(\bm{\pi},t,s_{t}\right)$triangleright $\bm{\pi}(1:t)$ consists of $S$
-7:     $\mathbf{a}^{(t)}\leftarrow\textbf{0}_{n}$, $\bm{\pi}_{a}\leftarrow\bm{\pi}(t:n)$triangleright “a” for active
-8:     $\mathbf{a}^{(t)}\left(\bm{\pi}_{a}\right)\leftarrow\mathbf{X}\left(\bm{\pi}_{a},:\right)\mathbf{x}_{s_{t}}-\mathbf{L}^{(t-1)}\left(\bm{\pi}_{a},1:t-1\right){\mathbf{L}^{(t-1)}\left(s_{t},1:t-1\right)}^{\top}$triangleright $O\left(nd+nt\right)$
-9:     $\mathbf{L}^{(t)}\leftarrow\left[\mathbf{L}^{(t-1)},\mathbf{a}^{(t)}/\sqrt{\mathbf{a}^{(t)}\left(s_{t}\right)}\right]$triangleright $\mathbf{a}^{(t)}\left(s_{t}\right)=\mathbf{d}^{(t-1)}(s_{t})$
-10:     $\mathbf{d}^{(t)}(i)\leftarrow 0\leavevmode\nobreak\ \forall\leavevmode\nobreak\ i\in S$, $\mathbf{d}^{(t)}(i)\leftarrow\mathbf{d}^{(t-1)}(i)-\frac{\left(\mathbf{a}^{(t)}(i)\right)^{2}}{\mathbf{a}^{(t)}\left(s_{t}\right)}\leavevmode\nobreak\ \forall\leavevmode\nobreak\ i\notin S$triangleright $O(n)$
-Remark 3.1 (Relation between QR with sequential random pivoting and RPCholesky).
-Sequential random pivoting (Algorithm 1) on $\mathbf{X}$ is equivalent to RPCholesky (Algorithm 2 [5, Algorithm 2.1]) on $\mathbf{X}\mathbf{X}^{\top}$ in the exact arithmetic.
-Proof 3.2 (Rationale for Remark 3.1).
-To show the equivalence in the exact arithmetic, it is sufficient to observe that Algorithms 1 and 2 yield the same $\mathbf{a}^{(t)}$ and $\mathbf{d}^{(t)}$ in each iteration $t=1,2,\cdots$.
-By induction, when $t=1$, both Algorithms 1 and 2 lead to $\mathbf{a}^{(1)}=\mathbf{X}\mathbf{x}_{s_{1}}$ and
-$\mathbf{d}^{(1)}(i)=\left\|\mathbf{x}_{i}\right\|_{2}^{2}-\frac{\left(\mathbf{x}_{i}^{\top}\mathbf{x}_{s_{1}}\right)^{2}}{\left\|\mathbf{x}_{s_{1}}\right\|_{2}^{2}}=\left\|\left(\mathbf{I}_{d}-\mathbf{Q}^{(1)}\left(\mathbf{Q}^{(1)}\right)^{\top}\right)\mathbf{x}_{i}\right\|_{2}^{2}\quad\forall\leavevmode\nobreak\ i\in[n]\setminus\left\{s_{1}\right\},$
-while $\mathbf{d}^{(1)}(s_{1})=0$. We also observe that $\mathbf{L}^{(1)}=\mathbf{X}\mathbf{Q}^{(1)}$ since $\mathbf{a}^{(1)}\left(s_{1}\right)=\left\|\mathbf{x}_{s_{1}}\right\|_{2}^{2}$.
-Suppose that in the $(t-1)$-th iteration, Algorithms 1 and 2 produce equivalent $\mathbf{a}^{(t-1)}$ and $\mathbf{d}^{(t-1)}$ under the exact arithmetic, while $\mathbf{d}^{(t-1)}\left(i\right)=\left\|\left(\mathbf{I}_{d}-\mathbf{Q}^{(t-1)}\left(\mathbf{Q}^{(t-1)}\right)^{\top}\right)\mathbf{x}_{i}\right\|_{2}^{2}$ for all $i\in[n]$ and $\mathbf{L}^{(t-1)}=\mathbf{X}\mathbf{Q}^{(t-1)}$.
-Then in the $t$-th iteration, both Algorithms 1 and 2 compute
-$\mathbf{a}^{(t)}=\mathbf{X}\mathbf{x}_{s_{t}}-\mathbf{X}\mathbf{Q}^{(t-1)}\left(\mathbf{Q}^{(t-1)}\right)^{\top}\mathbf{x}_{s_{t}}=\mathbf{X}\left(\mathbf{I}_{d}-\mathbf{Q}^{(t-1)}\left(\mathbf{Q}^{(t-1)}\right)^{\top}\right)\mathbf{x}_{s_{t}},$
-$\mathbf{L}^{(t)}=\mathbf{X}\mathbf{Q}^{(t)}$, and for all $i\in[n]$,
-$\mathbf{d}^{(t)}(i)=$ $\left\|\left(\mathbf{I}_{d}-\mathbf{Q}^{(t-1)}\left(\mathbf{Q}^{(t-1)}\right)^{\top}\right)\mathbf{x}_{i}\right\|_{2}^{2}-\frac{\left(\mathbf{x}_{i}^{\top}\left(\mathbf{I}_{d}-\mathbf{Q}^{(t-1)}\left(\mathbf{Q}^{(t-1)}\right)^{\top}\right)\mathbf{x}_{s_{t}}\right)^{2}}{\mathbf{x}_{s_{t}}^{\top}\left(\mathbf{I}_{d}-\mathbf{Q}^{(t-1)}\left(\mathbf{Q}^{(t-1)}\right)^{\top}\right)\mathbf{x}_{s_{t}}}$
-$=$ $\left\|\left(\mathbf{I}_{d}-\mathbf{Q}^{(t)}\left(\mathbf{Q}^{(t)}\right)^{\top}\right)\mathbf{x}_{i}\right\|_{2}^{2}.$
-In light of the equivalence between Algorithms 1 and 2, [5, Theorem 3.1] provides the following skeleton complexity guarantee for sequential random pivoting.
-Proposition 3.3 (Sequential random pivoting).
-For any $\epsilon>0$ and $r\in[n]$, the skeleton subset $S$ selected by Algorithm 1 provides $\mathbb{E}\left[\mathcal{E}_{\mathbf{X}}\left(S\right)\right]\leq\left(1+\epsilon\right)\left\|\mathbf{X}-{\mathbf{X}}_{\langle r\rangle}\right\|_{F}^{2}$ when333 Without strictly specifying, we generally assume that $\epsilon>0$ is small enough such that the $(r,\epsilon)$-ID is a non-trivial approximation, i.e.
+The idea of random pivoting is ubiquitous in various related problems, e.g.
+the combination of volume sampling and adaptive sampling for low-rank approximation [21], the Randomly Pivoted Cholesky (RPCholesky) [8] for kernel column Nyström approximation [52, 19.2], and the $D^{2}$-sampling for k-means++ clustering [2].
+Specifically, in light of the close connection between partially pivoted QR and pivoted Cholesky factorization (see e.g.
+[38, §5.2] and [54, 8]), applying RPCholesky to $\mathbf{X}\mathbf{X}^{\top}\in\mathbb{R}^{n\times n}$ is equivalent to Algorithm 1 in exact arithmetic.
+In light of such equivalence, [8, Theorem 5.1] provides a nearly optimal skeleton complexity guarantee for sequential random pivoting:
+Proposition 3.1 (Sequential random pivoting [8]).
+For any $\epsilon>0$ and $r\in[n]$, the skeleton subset $S$ selected by Algorithm 1 provides $\mathbb{E}\left[\mathcal{E}_{\mathbf{X}}\left(S\right)\right]\leqslant\left(1+\epsilon\right)\left\|\mathbf{X}-{\langle\mathbf{X}\rangle}_{r}\right\|_{F}^{2}$ when333 Without strictly specifying, we generally assume that $\epsilon>0$ is small enough such that the $(r,\epsilon)$-ID is a non-trivial approximation, i.e.
 $\epsilon<\min\left\{1/\eta_{r},2^{r}\right\}$ such that both logarithmic terms are positive.
 Otherwise, the skeleton complexity guarantee still holds by replacing any negative logarithmic terms with zeros.
-$k=\left|S\right|\geq\frac{r}{\epsilon}+r\cdot\min\left\{\log\left(\frac{1}{\epsilon\eta_{r}}\right),1+\log\left(\frac{2^{r}}{\epsilon}\right)\right\}.$
-From Proposition 3.3, we notice that sequential random pivoting almost matches the nearly optimal skeleton complexity guarantee of DPP (Equation 7) up to some logarithmic factors $\log\left(\frac{1}{\epsilon\eta_{r}}\right)$ on $r$.
-Further, the appealing skeleton complexity of random pivoting is empirically demonstrated in the form of RPCholesky for column Nyström approximation in [5].
-Although random pivoting leverages the strength of both adaptiveness and randomness and achieves an attractive skeleton complexity with affordable asymptotic complexity, the lack of parallelizability due to its sequential nature becomes a critical concern when considering empirical runtime.
+$k=\left|S\right|\geqslant\frac{r}{\epsilon}+r\cdot\min\left\{\log\left(\frac{1}{\epsilon\eta_{r}}\right),1+\log\left(\frac{2^{r}}{\epsilon}\right)\right\}.$
+Although random pivoting combines the strength of both adaptiveness and randomness and achieves an attractive skeleton complexity, the lack of hardware efficiency due to its sequential nature becomes a critical concern when considering empirical runtime.
 Remark 3.2 (Inefficiency of sequential updates).
@@
 This is largely due to the greater potential for the Level-3 BLAS to make more efficient use of memory caching in the processor.
-Notice that the computational bottleneck of Algorithm 1 with asymptotic complexity $O\left(ndk\right)$ (or $O\left(ndk\right)$) consists of $k$ matrix-vector multiplications (Level-2 BLAS) with $\mathbf{X}$.
+Notice that the computational bottleneck of Algorithm 1 with asymptotic complexity $O\left(ndk\right)$ consists of $k$ matrix-vector multiplications (Level-2 BLAS) with $\mathbf{X}$.
 This leads to the considerable slowdown of Algorithm 1 in practice.
@@
 Alternative to random pivoting, randomness can be combined with adaptiveness through randomized linear embedding (also known as sketching), which leads to the sketchy pivoting methods [50, 62, 23, 28].
-As given in Algorithm 3, the general framework of sketchy pivoting [16, Algorithm 1] consists of two stages.
-Algorithm 3 Sketchy pivoting [48, 16] (SkLUPP/SkCPQR)
-1:(a) $\mathbf{X}=\left[\mathbf{x}_{1},\cdots,\mathbf{x}_{n}\right]^{\top}\in\mathbb{R}^{n\times d}$, (b) Target rank $k\in\mathbb{N}$, (c) Sample size $l\geq k$, (d) Pivoting strategy: LUPP/CPQR
-2:(a) $S=\left\{s_{1},\cdots,s_{k}\right\}\subset[n]$, (b) $\mathbf{L}\in\mathbb{R}^{n\times k}$, (c) $\bm{\pi}\in\mathfrak{S}_{n}$
-3:Draw a randomized linear embedding $\bm{\Omega}\in\mathbb{R}^{d\times l}$
-4:$\mathbf{Y}\leftarrow\mathbf{X}\bm{\Omega}\in\mathbb{R}^{n\times l}$triangleright $O\left(ndk\right)$
-5:if LUPP then
-[…]
-7:else if CPQR then
-[…]
-[…]
+The general framework of sketchy pivoting [23, Algorithm 1] consists of two stages.
 In the first stage, randomness is incorporated through sketching, which serves two purposes.
-  • From the computational efficiency perspective, sketching reduces the data dimension from $d$ to $l=O(k)$444 For applications like low-rank approximations and skeleton selection, a small constant oversampling like $k+5\leq l\leq k+10$ is usually sufficient in practice.
-, while transferring the $O(ndk)$ computational bottleneck from the sequential greedy pivoting to the (input-sparsity-time and) parallelizable matrix-matrix multiplications.
+  • From the computational efficiency perspective, sketching reduces the data dimension444 For applications like low-rank approximations and skeleton selection, constant oversampling like $l\geqslant k+10$ is usually sufficient in practice.
+We refer the interested readers to [37, 64, 52] for a general review of sketching.
+from $d$ to $l=O(k)$ and transfers the $O(ndk)$ computational bottleneck from sequential greedy pivoting to input-sparsity-time matrix-matrix multiplications.
   • From the skeleton complexity perspective, applying randomized embeddings to the row space of $\mathbf{X}$ via sketching improves the empirical robustness of the subsequent greedy pivoting [61, 23].
 Intuitively, this is because the scarce adversarial inputs for greedy pivoting (Remark 2.1) are effectively negligible under randomization.
-Common choices of randomized linear embeddings in the sketching stage include Gaussian embeddings [28] (with asymptotic complexity $O\left(ndk\right)$), subsampled randomized trigonometric transforms [50, 43, 46, 4] (with asymptotic complexity $O\left(nd\log(k)\right)$), and sparse embeddings [36, 37, 7, 47] (with asymptotic complexity $O\left(nd\right)$).
-We refer the interested readers to [26, 49, 35] for a general review of sketching.
 In the second stage, adaptiveness is utilized as greedy pivoting on the sketched sample matrix $\mathbf{Y}=\mathbf{X}\bm{\Omega}\in\mathbb{R}^{n\times l}$ to select skeletons.
-Both LUPP [22, Section 3.4] and CPQR [22, Section 5.4.2] on $\mathbf{Y}$ cost $O(nk^{2})$ asymptotically, whereas LUPP is remarkably faster in practice due to its superior parallelizability [16].
+Both LUPP [33, Section 3.4] and CPQR [33, Section 5.4.2] on $\mathbf{Y}$ cost $O(nk^{2})$ asymptotically, whereas LUPP is remarkably faster in practice due to its superior parallelizability [23].
 Intuitively, applying greedy pivoting on the top of sketching can be viewed as an analog of the squared-norm sampling with a slightly different squared-norm-based distribution.
-For example, for a Gaussian random matrix $\bm{\Omega}$ with i.i.d.
-entries from $\mathcal{N}\left(0,1/l\right)$, each row $\mathbf{y}_{i}=\bm{\Omega}^{\top}\mathbf{x}_{i}$ of $\mathbf{Y}$ is a Gaussian random vector with i.i.d.
-entries from $\mathcal{N}\left(\textbf{0},\left\|\mathbf{x}_{i}\right\|_{2}^{2}\mathbf{I}_{l}\right)$555 It is worth highlighting that although rows of $\mathbf{Y}$ consist of i.i.d.
-entries, entries in each column of $\mathbf{Y}$ are dependent.
-In particular, for all $j\in[l]$, $\mathrm{Cov}\left(\mathbf{Y}\left(:,j\right)\right)=\mathbf{X}\mathbf{D}^{-1}\mathbf{X}^{\top}$ where $\mathbf{D}=\mathop{\mathrm{diag}}\left(\left\|\mathbf{x}_{1}\right\|_{2}^{2},\cdots,\left\|\mathbf{x}_{n}\right\|_{2}^{2}\right)$.
-.
+Remark 3.3 (Sketchy pivoting v.s. random pivoting).
+For example, consider a Gaussian random matrix $\bm{\Omega}$ with i.i.d.
+entries from $\mathcal{N}\left(0,1/l\right)$.
+Each row $\mathbf{y}_{i}=\bm{\Omega}^{\top}\mathbf{x}_{i}$ of $\mathbf{Y}$ is a Gaussian random vector with i.i.d.
+entries from $\mathcal{N}(\textbf{0},\left\|\mathbf{x}_{i}\right\|_{2}^{2}\mathbf{I}_{l})$.
 When applying greedy squared-norm pivoting (CPQR) in the second stage, the first pivot is selected according to $s_{1}\leftarrow\mathop{\mathrm{argmax}}_{i\in[n]}\left\|\mathbf{x}_{i}\right\|_{2}^{2}Z_{i}$ where every $Z_{i}=\sum_{j=1}^{l}Y_{ij}^{2}$ is a $\chi^{2}_{l}$ random variable but $\left\{Z_{i}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ i\in[n]\right\}$ are dependent.
 The intuition behind sketchy pivoting is similar to random pivoting, but the randomness arises from the random variables $Z_{i}$, instead of sampling.
-In the extreme scenarios, when the variance of $\left\{Z_{i}\right\}_{i\in[n]}$ is negligible in comparison to the squared norms $\left\{\left\|\mathbf{x}_{i}\right\|_{2}^{2}\right\}_{i\in[n]}$, sketchy pivoting behaves similarly to greedy pivoting.
+In the extreme scenarios, when the variance of $\left\{Z_{i}\right\}_{i\in[n]}$ is negligible in comparison to the squared norms $\{\left\|\mathbf{x}_{i}\right\|_{2}^{2}\}_{i\in[n]}$, sketchy pivoting behaves similarly to greedy pivoting.
 Meanwhile, when the variance of $\left\{Z_{i}\right\}_{i\in[n]}$ is much larger than differences among the squared norms, sketchy pivoting tends to behave like uniform sampling.
 Intuitively, there exists some $\left\{Z_{i}\right\}_{i\in[n]}$ with appropriate variance interpolating the two extreme cases such that sketchy pivoting mimics the behavior of squared-norm sampling and enjoys a nearly optimal skeleton complexity guarantee similar to Proposition 3.1.
-In practice, sketchy pivoting (Algorithm 3) with Gaussian embedding provides high-quality skeleton selection with comparable error $\mathcal{E}_{\mathbf{X}}\left(S\right)$ to random pivoting (Algorithm 1, cf.
+In practice, sketchy pivoting provides high-quality skeleton selection with comparable error $\mathcal{E}_{\mathbf{X}}\left(S\right)$ to random pivoting (Algorithm 1, cf.
 Figure 1), but much more efficiently thanks to the parallelizability of sketching [23].
-However, as a major empirical limitation, sketchy pivoting is not error-revealing and requires prior knowledge of the target rank $k$.
+However, as a major empirical limitation, sketchy pivoting is not rank-adaptive and requires prior knowledge of the target rank $k$.
 4 Robust Blockwise Random Pivoting
-In light of the existing skeleton selection methods summarized in Table 1 that leverage/combine randomness and adaptiveness, a parallelizable, error-revealing, and exactly ID-revealing skeleton selection algorithm that attains similar skeleton and asymptotic complexities as the sequential random pivoting is a desirable missing piece.
-Blockwise random pivoting is a natural extension of its sequential variation Algorithm 1 that consists of matrix-matrix multiplications and is therefore parallelizable.
-In the kernel formulation (with an input kernel matrix $\mathbf{X}\mathbf{X}^{\top}$), [5, Algorithm 2.2] introduced a blocked version of RPCholesky that can be generalized as blockwise random pivoting (BRP—Algorithm 4 with $\tau_{b}=0$).
-However, for a large block size $b\in\mathbb{N}$ that improves parallelizability, such plain BRP can pick up to $b$ times more skeletons than necessary due to the similar pitfall of squared-norm sampling instantiated in Example 2.4 (e.g.
+In light of the existing skeleton selection methods summarized in Table 1, a hardware-efficient, rank-adaptive, and ID-revealing skeleton selection algorithm that attains similar skeleton and asymptotic complexities as the sequential random pivoting raises as a desirable missing piece.
+Blockwise random pivoting is a natural extension of its sequential variation Algorithm 1 that consists of matrix-matrix multiplications and is, therefore, hardware-efficient.
+In the kernel formulation (with an input kernel matrix $\mathbf{X}\mathbf{X}^{\top}$), [8, Algorithm 3] introduced a blocked version of RPCholesky that can be generalized as blockwise random pivoting (BRP—Algorithm 2 with $\tau_{b}=0$).
+However, when the block size $b\in\mathbf{N}$ is taken to be large enough to fully realize the advantages of hardware efficiency, such plain BRP can have up to a $b$ times larger skeleton complexity than necessary due to the similar pitfall of squared-norm sampling instantiated in Example 2.3 (e.g.
 Figure 1).
-Moreover, in practice, the numerical instability issue tends to be exacerbated as the block size increases.
 Algorithm 2 Robust blockwise random pivoting (RBRP)
@@
 3:$\mathbf{L}^{(0)}\leftarrow\textbf{0}_{n\times 0}$, $\mathbf{Q}^{(0)}\leftarrow\textbf{0}_{d\times 0}$, $\bm{\pi}\leftarrow[1,\cdots,n]$, $S\leftarrow\emptyset$, $t=0$
-4:$\mathbf{d}^{(0)}\in\mathbb{R}_{\geq 0}^{n}$ with $\mathbf{d}^{(0)}(i)=\left\|\mathbf{x}_{i}\right\|_{2}^{2}\leavevmode\nobreak\ \forall\leavevmode\nobreak\ i\in[n]$triangleright $O(nd)$
+4:$\mathbf{d}^{(0)}\in\mathbb{R}_{\geqslant 0}^{n}$ with $\mathbf{d}^{(0)}(i)=\left\|\mathbf{x}_{i}\right\|_{2}^{2}\leavevmode\nobreak\ \forall\leavevmode\nobreak\ i\in[n]$triangleright $O(nd)$
 5:while $\left\|\mathbf{d}^{(t)}\right\|_{1}>\tau\left\|\mathbf{d}^{(0)}\right\|_{1}$ (or $\left|S\right|<k$do
@@
 […]
-10:     $b^{\prime}\leftarrow\max\left\{i\in[b]\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \left\|\mathbf{R}_{\mathbf{V}}\left(i:b,i:b\right)\right\|_{F}^{2}\geq\tau_{b}\cdot\left\|\mathbf{R}_{\mathbf{V}}\right\|_{F}^{2}\right\}$triangleright $O\left(b^{2}\right)$
+10:     $b^{\prime}\leftarrow\max\left\{i\in[b]\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \left\|\mathbf{R}_{\mathbf{V}}\left(i:b,i:b\right)\right\|_{F}^{2}\geqslant\tau_{b}\cdot\left\|\mathbf{R}_{\mathbf{V}}\right\|_{F}^{2}\right\}$triangleright $O\left(b^{2}\right)$
 11:     $S^{\prime}_{t}\leftarrow S_{t}\left(\bm{\pi}_{\mathbf{V}}\left(1:b^{\prime}\right)\right)$, $\mathbf{Q}^{\prime}_{\mathbf{V}}\leftarrow\mathbf{Q}_{\mathbf{V}}\left(:,1:b^{\prime}\right)$
@@
 Similar inefficiency also appears in BRP but is generally alleviated thanks to randomness (Figure 1 (left)).
-To resolve the challenges illustrated by Example 4.1, we introduce a robust blockwise random pivoting (RBRP) algorithm—Algorithm 4—that empirically achieves comparable skeleton complexities to the sequential adaptive methods (SRP and CPQR), as demonstrated in Figure 1 (right), while exploiting the parallelizability of matrix-matrix multiplications.
+To resolve the challenges illustrated by Example 4.1, we introduce a robust blockwise random pivoting (RBRP) algorithm—Algorithm 2—that empirically achieves comparable skeleton complexities to the sequential adaptive methods (SRP and CPQR), as demonstrated in Figure 1 (right), while exploiting the hardware efficiency of matrix-matrix multiplications.
 Remark 4.2 (Robust blockwise filtering).
@@
 With small constant block sizes, the robust blockwise filtering based on local CPQR can be computed with negligible additional cost $O(db^{2})$, despite the sequential nature of CPQR.
-In general, Algorithm 4 takes $O\left(ndk+bdk^{2}\right)$ time asymptotically (with a slightly larger lower-order term, cf.
-$O\left(ndk+dk^{2}\right)$ for the sequential version in Algorithm 1), where the dominant cost $O(ndk)$ consists of $k$ parallelizable passes through $\mathbf{X}$.
-Analogous to Algorithm 1, the storage of $\mathbf{L}^{(t)}$ and $\mathbf{Q}^{(t)}$ requires $O(nk)$ and $O(dk)$ memory, respectively.
-Meanwhile, the adaptiveness grants Algorithm 4 the blockwise error-revealing ability.
+The worst-case asymptotic complexity of Algorithm 2 is $O\left(ndk+bdk^{2}\right)$ (with a slightly larger lower-order term, cf.
+$O\left(ndk+dk^{2}\right)$ for SRP in Algorithm 1) when each step of robust blockwise filtering in Algorithm 2 only keeps a single skeleton.
+In practice, robust blockwise filtering typically retains a reasonable portion of skeletons in each block, leading to a time complexity close to $O\left(ndk+dk^{2}\right)$ (see Section 6.2).
+The dominant cost $O(ndk)$ consists of $k$ hardware-efficient passes through $\mathbf{X}$ as matrix-matrix multiplications.
+Analogous to Algorithm 1, the storage of $\mathbf{L}^{(t)}$ and $\mathbf{Q}^{(t)}$ requires $O(nk)$ and $O(dk)$ memory.
+The adaptive nature of Algorithm 2 further facilitates the rank-adaptiveness.
 Table 2: Summary of abbreviations for the skeleton selection methods in the experiments.
 Abbreviation Skeleton selection method Algorithm
-CPQR Sequential greedy pivoting [22, Section 5.4.2]
-SqNorm Squared-norm sampling Proposition 2.3 [13]
-DPP DPP sampling [2, Equation 8][31, 21]
+CPQR Sequential greedy pivoting [33, Section 5.4.2]
+SqNorm Squared-norm sampling (5[20]
+DPP DPP sampling [5, Equation 8][45, 31]
 SRP Sequential random pivoting Algorithm 1 [8]
-SkCPQR Sketchy pivoting with CPQR Algorithm 3 [48]
-SkLUPP Sketchy pivoting with LUPP Algorithm 3 [16]
+SkCPQR Sketchy pivoting with CPQR [62, Algorithm 4]
+SkLUPP Sketchy pivoting with LUPP [23, Algorithm 1]
 BGP Blockwise greedy pivoting Remark 4.3 with $\tau_{b}=0$
@@
 With $\tau_{b}=0$, BGP is reduced to CPQR. For adversarial inputs like the one in Example 4.1, BGP tends to suffer from far worse skeleton complexities than those of BRP (Figure 1 (left)).
-Figure 1: The skeletonization error $\mathcal{E}_{\mathbf{X}}\left(S\right)$ of algorithms in Table 2 on a GMM-based adversarial input as described in Example 4.1, where the adaptiveness is critical for achieving a good skeleton complexity/accuracy.
-(Left): Sequential random pivoting (SRP) and greedy pivoting (CPQR) enjoy the best skeleton complexities.
-The plain blockwise random pivoting (BRP) and greedy pivoting (BGP) tend to suffer from observably higher skeleton complexities than the sequential adaptive methods, while greedy selection tends to severely exacerbate such inefficiency.
-The squared-norm sampling (SqNorm) can be viewed as plain BRP with block size $b=k$, which suffers from a similar inefficiency as BRP/BGP.
-DPP sampling attains relatively good accuracy even without adaptiveness, thanks to its nearly optimal skeleton complexity [2, 25, 5], but with a much higher runtime in practice as a trade-off (e.g., with [21], k-DPP takes about $10^{3}\times$ longer than other algorithms in Table 2).
-(Right): With robust blockwise filtering (Remark 4.2), robust blockwise random pivoting (RBRP) and greedy pivoting (RBGP), as well as the sketchy pivoting algorithms, achieves the comparable, nearly optimal skeleton complexities as the sequential adaptive methods.
-It is worth highlighting that despite the comparable skeleton complexities between RBRP and RBGP, RBGP is generally much slower than RBRP (cf.
-Figure 2).
-This is because, without randomness, blockwise greedy pivoting tends to pick more redundant points in each block, which are later filtered out.
-By comparing Figure 1 (left) and (right), we observe that the robust blockwise filtering (Remark 4.2) improves both blockwise random and greedy pivoting.
-In particular, with $\tau_{b}=\frac{1}{b}$, RBRP (and RBGP) empirically attains comparable skeleton complexities to those of the sequential random pivoting on the GMM adversarial input (Example 4.1).
-The same observation extends to all the natural and synthetic data matrices tested in our experiments (Section 6), which leads to the following 4.4:
+Figure 1: The relative skeletonization error $\mathcal{E}_{\mathbf{X}}\left(S\right)/\|\mathbf{X}\|_{F}^{2}$ and skeleton selection time (excluding interpolation matrix construction) of algorithms in Table 2 on the adversarial GMM matrix described in Section 6.1 constructed according to Example 4.1.
+For the sketchy pivoting methods, we only show the results of SkLUPP, as SkCPQR is known to have similar skeleton complexities as SkLUPP but with higher runtimes [23].
+Figure 1 concretizes the adversarial input in Example 4.1 and demonstrates the empirical effectiveness of RBRP for skeleton selection.
+  • In particular, the skeletonization errors on the left unveil the failure of plain BRP/BGP in terms of large skeleton complexities due to the lack of adaptiveness.
+Moreover, greedy selection in BGP tends to severely exacerbate this problem.
+SqNorm can be viewed as plain BRP with block size $b=k$ and therefore suffers from similar large skeleton complexities as BRP/BGP, despite its low asymptotic complexity of $O(nd)$.
+  • DPP sampling attains relatively good accuracy even without adaptiveness, thanks to its nearly optimal skeleton complexity [5, 36, 8], but with much higher runtimes as a trade-off (e.g., with [31], k-DPP takes about $10^{3}\times$ longer than other algorithms in Table 2).
+  • Providing even better skeleton complexities than DPP, SRP and CPQR achieve one of the best skeletonization errors thanks to their adaptiveness.
+However, the sequential nature of SRP and CPQR leads to higher runtimes than the blockwise methods.
+  • With robust blockwise filtering (Remark 4.2), RBRP achieves comparable skeleton complexities to SRP and CPQR, while maintaining the hardware efficiency of blockwise methods.
+Notice from the runtimes on the right that despite the comparable skeleton complexities to RBRP, RBGP is slower than RBRP on the adversarial GMM matrix.
+This is because RBGP with greedy selection tends to pick more redundant points in each block, which are later removed by the robust blockwise filtering.
+  • For skeleton selection only, SkLUPP provides the state-of-the-art performance — one of the best skeletonization errors with the most competitive runtimes.
+However, the lack of rank-adaptiveness and ID-revealing property becomes a major drawback for SkLUPP as an ID algorithm (see Section 5.2).
+Beyond the GMM adversarial input, extensive numerical experiments in Section 6 show that RBRP with $\tau_{b}=\frac{1}{b}$ empirically attains comparable skeleton complexities as sequential random pivoting, which leads to the following conjecture.
 Conjecture 4.4 (Skeleton complexity of RBRP).
-RBRP (Algorithm 4) with $\tau_{b}=\frac{1}{b}$ shares a skeleton complexity similar to sequential random pivoting (Proposition 3.3) in expectation.
+RBRP (Algorithm 2) with $\tau_{b}=\frac{1}{b}$ has similar skeleton complexity as sequential random pivoting (Proposition 3.1) in expectation.
 Proof 4.5 (Rationale for 4.4).
-Consider Algorithm 4 at the $t$-th step with $\left|S\right|$ selected skeletons from the previous steps, $\left|S_{t}\right|=b$ candidates drawn from $\frac{\mathbf{d}^{(t-1)}}{\left\|\mathbf{d}^{(t-1)}\right\|_{1}}$, and $\left|S^{\prime}_{t}\right|=b^{\prime}$ remaining skeletons in $S_{t}$ that pass through the robust blockwise filtering (Remark 4.2).
+Consider Algorithm 2 at the $t$-th step with $\left|S\right|$ selected skeletons from the previous steps, $\left|S_{t}\right|=b$ candidates drawn from ${\mathbf{d}^{(t-1)}}/{\left\|\mathbf{d}^{(t-1)}\right\|_{1}}$, and $\left|S^{\prime}_{t}\right|=b^{\prime}$ remaining skeletons in $S_{t}$ that pass through the robust blockwise filtering (Remark 4.2).
 Now suppose we switch to sequential random pivoting after the $(t-1)$-th step and draw $b^{\prime}$ skeletons adaptively, denoted as $S^{\prime\prime}_{t}$.
 Then, the key rationale for 4.4 is that $\mathcal{E}_{\mathbf{X}}\left(S\cup S^{\prime}_{t}\right)\approx\mathcal{E}_{\mathbf{X}}\left(S\cup S^{\prime\prime}_{t}\right)$ with (a reasonably) high probability when $\tau_{b}=\frac{1}{b}$.
-To see this, we observe the close connection between the selection scheme of $S^{\prime}_{t}$ and that of $S^{\prime\prime}_{t}$:
-  • The robust blockwise filtering for $S^{\prime}_{t}$ in RBRP adaptively selects the point with the maximum residual norm in each step (i.e.
-CPQR on the residual matrix $\mathbf{V}$ in Algorithm 4), within a small subset $S_{t}$ of $b$ points sampled according to the squared-norm distribution over $[n]\setminus S$.
+On this, we observe the close connection between the selection scheme of $S^{\prime}_{t}$ and that of $S^{\prime\prime}_{t}$:
+  • The robust blockwise filtering for $S^{\prime}_{t}$ in RBRP adaptively selects the point with the maximum residual norm in each step (via CPQR on $\mathbf{V}$ in Algorithm 2), within a small subset $S_{t}$ of $b$ points sampled according to the squared-norm distribution over $[n]\setminus S$.
   • The random pivoting for $S^{\prime\prime}_{t}$ in SRP adaptively selects the point according to the squared-norm distribution of the residual over the complement of the skeleton subset in each step.
-Intuitively, every selection made for $S^{\prime}_{t}$ and $S^{\prime\prime}_{t}$ yields similar decay in the skeletonization error if the squared norms of the residual matrix, corresponding to remaining points in $S_{t}$, are around the maxima in the complement of the skeleton subset, so that these remaining points in $S_{t}$ can be selected by SRP with high probability666 It is worth mentioning that, in contrast to SRP, RBRP does not recompute the entire residual matrix after each skeleton selection.
+Intuitively, every selection made for $S^{\prime}_{t}$ and $S^{\prime\prime}_{t}$ yields similar decay in the skeletonization error if the squared norms of the residual matrix, corresponding to remaining points in $S_{t}$, are around the maxima in the complement of the skeleton subset so that these remaining points in $S_{t}$ can be selected by SRP with high probability555 It is worth mentioning that, in contrast to SRP, RBRP does not recompute the entire residual matrix after each skeleton selection.
 Instead, within each block, RBRP computes the residual matrix locally for $\mathbf{X}\left(S_{t},:\right)$ only, while updating the entire residual matrix at the end of each block via matrix-matrix multiplication.
@@
 Instead of computing the squared norms of the entire residual matrix explicitly for each selection as SRP (which leads to the undesired sequential updates), RBRP leverages the threshold $\tau_{b}=\frac{1}{b}$ to encourage the first $b^{\prime}$ points selected by the robust blockwise filtering in each block to have top squared norms in the corresponding residual matrices.
-Notice that with $\tau_{b}=\frac{1}{b}$, we enforce the sum of squared norms of the $b^{\prime}$ points to be larger than the average squared norm of the original block: $\left\|\mathbf{R}_{\mathbf{V}}\left(i:b,i:b\right)\right\|_{F}^{2}\geq\frac{1}{b}\left\|\mathbf{V}\right\|_{F}^{2}$.
+Notice that with $\tau_{b}=\frac{1}{b}$, we enforce the sum of squared norms of the $b^{\prime}$ points to be larger than the average squared norm of the original block: $\left\|\mathbf{R}_{\mathbf{V}}\left(i:b,i:b\right)\right\|_{F}^{2}\geqslant\frac{1}{b}\left\|\mathbf{V}\right\|_{F}^{2}$.
 Since the squared norms of residuals are non-increasing under QR updates, $\frac{1}{b}\left\|\mathbf{V}\right\|_{F}^{2}$ serves as an overestimate for the squared norms of redundant points sampled in $S_{t}$.
 As toy examples, if $\mathbf{V}$ is rank deficient (i.e.
-$\mathop{\mathrm{rank}}(\mathbf{V})=b^{\prime\prime}<b$), then there exists $b^{\prime}\leq b^{\prime\prime}$ such that $\left\|\mathbf{R}_{\mathbf{V}}\left(b^{\prime}:b,b^{\prime}:b\right)\right\|_{F}^{2}\geq\frac{1}{b}\left\|\mathbf{V}\right\|_{F}^{2}$, but
-$0=\left\|\mathbf{R}_{\mathbf{V}}\left(b^{\prime\prime}+1:b,b^{\prime\prime}+1:b\right)\right\|_{F}^{2}\leq\left\|\mathbf{R}_{\mathbf{V}}\left(b^{\prime}+1:b,b^{\prime}+1:b\right)\right\|_{F}^{2}<\frac{1}{b}\left\|\mathbf{V}\right\|_{F}^{2}.$
-Whereas if $\mathbf{V}\in\mathbb{R}^{d\times b}$ contains orthogonal columns with the same squared norm, then for all $i\in[b]$, we have $\left\|\mathbf{R}_{\mathbf{V}}\left(i:b,i:b\right)\right\|_{F}^{2}\geq\left|\mathbf{R}_{\mathbf{V}}\left(b,b\right)\right|^{2}=\frac{1}{b}\left\|\mathbf{V}\right\|_{F}^{2}$, and therefore $b^{\prime}=b$.
+$\mathop{\mathrm{rank}}(\mathbf{V})=b^{\prime\prime}<b$), then there exists $b^{\prime}\leqslant b^{\prime\prime}$ such that $\left\|\mathbf{R}_{\mathbf{V}}\left(b^{\prime}:b,b^{\prime}:b\right)\right\|_{F}^{2}\geqslant\frac{1}{b}\left\|\mathbf{V}\right\|_{F}^{2}$, but
+$0=\left\|\mathbf{R}_{\mathbf{V}}\left(b^{\prime\prime}+1:b,b^{\prime\prime}+1:b\right)\right\|_{F}^{2}\leqslant\left\|\mathbf{R}_{\mathbf{V}}\left(b^{\prime}+1:b,b^{\prime}+1:b\right)\right\|_{F}^{2}<\frac{1}{b}\left\|\mathbf{V}\right\|_{F}^{2}.$
+Whereas if $\mathbf{V}\in\mathbb{R}^{d\times b}$ contains orthogonal columns with the same squared norm, then for all $i\in[b]$, we have $\left\|\mathbf{R}_{\mathbf{V}}\left(i:b,i:b\right)\right\|_{F}^{2}\geqslant\left|\mathbf{R}_{\mathbf{V}}\left(b,b\right)\right|^{2}=\frac{1}{b}\left\|\mathbf{V}\right\|_{F}^{2}$, and therefore $b^{\prime}=b$.
 5 Interpolation Matrix Construction
 In this section, we consider the construction of the interpolation matrix $\mathbf{W}$ after skeleton selection.
-In particular, we are interested in exploiting information from the $O\left(ndk\right)$ skeleton selection algorithms in Table 1, i.e.
-the matrix $\mathbf{L}\in\mathbb{R}^{n\times k}$ in Algorithms 1, 3, and 4, to construct $\mathbf{W}$ efficiently in $O(nk^{2})$ time.
-We begin by noting the optimal interpolation matrix that solves the least square problem
+In particular, we are interested in leveraging information from the skeleton selection processes to construct $\mathbf{W}$ efficiently.
+Since the optimal interpolation matrix solves the following least-square problem:
 (6) $\min_{\mathbf{W}\in\mathbb{R}^{n\times k}}\left\|\mathbf{X}-\mathbf{W}\mathbf{X}_{S}\right\|_{F}^{2}\quad\Rightarrow\quad\mathbf{W}=\mathbf{X}\mathbf{X}_{S}^{\dagger}=\mathbf{X}\mathbf{X}_{S}^{\top}\left(\mathbf{X}_{S}\mathbf{X}_{S}^{\top}\right)^{-1},$
-so that we can compute such optimal $\mathbf{W}$ exactly in $O\left(ndk+dk^{2}+nk^{2}\right)$ time via QR decomposition on $\mathbf{X}_{S}$.
+we can compute such optimal $\mathbf{W}$ exactly in $O\left(ndk+dk^{2}+nk^{2}\right)$ time via QR decomposition on $\mathbf{X}_{S}$.
 However, the dominant cost $O(ndk)$ requires additional passes through $\mathbf{X}$ and can be prohibitive after the skeleton selection stage.
-5.1 Exactly ID-revealing Algorithms
-With the output matrix $\mathbf{L}\in\mathbb{R}^{n\times k}$ from Algorithms 1 and 4, the corresponding optimal interpolation matrix $\mathbf{X}\mathbf{X}_{S}^{\dagger}$ can be computed in $O(nk^{2})$ time following Algorithm 5.
+5.1 ID-revealing Algorithms
+With the output matrix $\mathbf{L}\in\mathbb{R}^{n\times k}$ from Algorithms 1 and 2, the corresponding optimal interpolation matrix $\mathbf{X}\mathbf{X}_{S}^{\dagger}$ can be computed in $O(nk^{2})$ time following Algorithm 3.
 Algorithm 3 Interpolation matrix construction (ID)
-1:(a) $\bm{\pi}\in\mathfrak{S}_{n}$ (such that $S=\bm{\pi}\left(1:k\right)$), (b) $\mathbf{L}\in\mathbb{R}^{n\times k}$,
-2:(a) $\mathbf{W}\in\mathbb{R}^{n\times k}$
 […]
-4:$\mathbf{W}\leftarrow\textbf{0}_{n\times k}$, $\mathbf{W}\left(\bm{\pi}\left(1:k\right),:\right)=\mathbf{I}_{k}$
+2:$\mathbf{W}\leftarrow\textbf{0}_{n\times k}$, $\mathbf{W}\left(\bm{\pi}\left(1:k\right),:\right)=\mathbf{I}_{k}$
 […]
 Proposition 5.1.
-The sequential (Algorithm 1) and blockwise (Algorithm 4) random pivoting, as well as their greedy pivoting variations, are exactly ID-revealing (Definition 1.2).
+The sequential (Algorithm 1) and blockwise (Algorithm 2) random pivoting, as well as their greedy pivoting variations, are ID-revealing (Definition 1.1).
 Proof 5.2 (Proof of Proposition 5.1).
-We first observe that the difference between sequential/blockwise random versus greedy pivoting lies only in the skeleton subset $S$.
+We first observe that the difference between sequential/blockwise random v.s.
+greedy pivoting lies only in the skeleton subset $S$.
 That is, assuming the same skeleton selection $S$, the resulting matrices $\mathbf{L}$ from random and greedy pivoting would be the same.
-Therefore, it suffices to show that Proposition 5.1 holds for random pivoting in Algorithms 1 and 4.
-For both Algorithms 1 and 4, recall that $\mathbf{L}\leftarrow\mathbf{L}^{(t)}$.
+Therefore, it suffices to show that Proposition 5.1 holds for random pivoting in Algorithms 1 and 2.
+For both Algorithms 1 and 2, recall that $\mathbf{L}\leftarrow\mathbf{L}^{(t)}$.
 We observe that $\mathbf{L}\mathbf{Q}^{(t)\top}$ provides a rank-$k$ approximation for $\mathbf{X}$: $\left\|\mathbf{X}-\mathbf{L}\mathbf{Q}^{(t)\top}\right\|_{F}^{2}=\left\|\mathbf{d}^{(t)}\right\|_{1}=\mathcal{E}_{\mathbf{X}}\left(S\right)<\tau\left\|\mathbf{X}\right\|_{F}^{2}$.
@@
 The computation of $\mathbf{L}_{2}\mathbf{L}_{1}^{-1}$ involves solving a small $k\times k$ system $(n-k)$ times, which takes $O(nk^{2})$ time in general.
-Overall, given $\mathbf{L}$ from Algorithms 1 and 4, we have that Algorithm 5 constructs the optimal interpolation matrix $\mathbf{W}=\mathbf{X}\mathbf{X}_{S}^{\dagger}=\mathbf{L}\mathbf{L}_{1}^{-1}$.
+Overall, given $\mathbf{L}$ from Algorithms 1 and 2, we have that Algorithm 3 constructs the optimal interpolation matrix $\mathbf{W}=\mathbf{X}\mathbf{X}_{S}^{\dagger}=\mathbf{L}\mathbf{L}_{1}^{-1}$.
 Moreover, when $\mathbf{L}_{1}$ is well-conditioned, the special lower triangular structure of $\mathbf{L}_{1}$ can be leveraged to further accelerate the construction of $\mathbf{W}$, i.e.
@@
 We emphasize that compared to backward substitution, computing the pseudoinverse via truncated SVD of the small $k\times k$ matrix $\mathbf{L}_{1}$ has negligible additional cost in the common low-rank approximation setting with $k\ll n$.
-5.2 Approximately ID-revealing Algorithms
-In contrast to Algorithms 1 and 4, $\mathbf{L}$ from the sketchy pivoting algorithms in Algorithm 3 is insufficient for computing the optimal interpolation matrix $\mathbf{X}\mathbf{X}_{S}^{\dagger}$ exactly, intuitively due to the loss of information in sketching (i.e.
-randomized dimension reduction).
-Proposition 5.4.
-The sketchy pivoting algorithms (Algorithm 3) are approximately ID-revealing.
-Proposition 5.4 follows directly from the observation that applying Algorithm 5 on $\mathbf{L}$ from Algorithm 3 (i.e.
-$\mathbf{W}=\mathbf{L}\mathbf{L}_{1}^{-1}$) solves a sketched version of Equation 8:
-(9) $\min_{\mathbf{W}\in\mathbb{R}^{n\times k}}\left\|\mathbf{X}\widehat{\bm{\Omega}}-\mathbf{W}\mathbf{X}_{S}\widehat{\bm{\Omega}}\right\|_{F}^{2}\quad\Rightarrow\quad\mathbf{W}=\mathbf{L}\mathbf{L}_{1}^{-1}=\mathbf{X}\widehat{\bm{\Omega}}\left(\mathbf{X}_{S}\widehat{\bm{\Omega}}\right)^{\dagger},$
-where $\widehat{\bm{\Omega}}=\bm{\Omega}\left(:,1:k\right)\in\mathbb{R}^{d\times k}$ consists of the first $k$ columns in the randomized linear embedding $\bm{\Omega}\in\mathbb{R}^{d\times l}$ drawn in Algorithm 3.
-Unfortunately, such sketched least square estimation [49, 42, 15] is known to be suboptimal as long as $l<d$ [40].
-Specifically for the interpolation matrix construction, the sketched estimation with a sample size as small as $k$ (i.e.
-without oversampling) in Algorithm 3 leads to large interpolation error $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)$, as numerically illustrated in Section 6.2 (e.g.
-comparing $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)$ as SkCPQR/SkLUPP-ID in Figure 2 and $\mathcal{E}_{\mathbf{X}}\left(S\right)$ as SkCPQR/SkLUPP in Figure 1 (right)).
+5.2 Oversampled Sketchy ID for Sketchy Pivoting
+In contrast to random pivoting, sketchy pivoting [23, Algorithm 1] fails to be ID-revealing, intuitively due to the loss of information in the sketching process.
+Concretely, applying Algorithm 3 on sketchy pivoting leads to a sketched version of (6):
+(7) $\min_{\mathbf{W}\in\mathbb{R}^{n\times k}}\left\|\mathbf{X}\widehat{\bm{\Omega}}-\mathbf{W}\mathbf{X}_{S}\widehat{\bm{\Omega}}\right\|_{F}^{2}\quad\Rightarrow\quad\mathbf{W}=\mathbf{X}\widehat{\bm{\Omega}}\left(\mathbf{X}_{S}\widehat{\bm{\Omega}}\right)^{\dagger},$
+where $\widehat{\bm{\Omega}}=\bm{\Omega}\left(:,1:k\right)\in\mathbb{R}^{d\times k}$ consists of the first $k$ columns in the randomized linear embedding $\bm{\Omega}\in\mathbb{R}^{d\times l}$ drawn in sketchy pivoting.
+Unfortunately, such sketched least-square estimation [64, 59, 22] is known to be suboptimal as long as $l<d$ [56].
+Specifically for interpolation matrix construction, (7) without oversampling leads to large interpolation error $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)\gg\mathcal{E}_{\mathbf{X}}\left(S\right)$, as numerically illustrated in Section 6.2 (cf.
+SkLUPP-ID in Figure 2 for $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)$ v.s.
+SkLUPP in Figure 1 for $\mathcal{E}_{\mathbf{X}}\left(S\right)$ (right)).
 Algorithm 4 Oversampled sketchy interpolation matrix construction (OSID)
-1:(a) $\bm{\pi}\in\mathfrak{S}_{n}$ ($S=\bm{\pi}\left(1:k\right)$), (b) Sample size $l\geq k$ (or pre-computed $\mathbf{Y}\in\mathbb{R}^{n\times l}$)
-2:(a) $\mathbf{W}\in\mathbb{R}^{n\times k}$
-3:(Draw a randomized linear embedding $\bm{\Omega}\in\mathbb{R}^{d\times l}$)
-4:($\mathbf{Y}\leftarrow\mathbf{X}\bm{\Omega}\in\mathbb{R}^{n\times l}$)
+1:(Draw a randomized linear embedding $\bm{\Omega}\in\mathbb{R}^{d\times l}$ and compute $\mathbf{Y}\leftarrow\mathbf{X}\bm{\Omega}\in\mathbb{R}^{n\times l}$)
 […]
-6:$\mathbf{W}\leftarrow\textbf{0}_{n\times k}$, $\mathbf{W}\left(\bm{\pi}\left(1:k\right),:\right)=\mathbf{I}_{k}$
+3:$\mathbf{W}\leftarrow\textbf{0}_{n\times k}$, $\mathbf{W}\left(\bm{\pi}\left(1:k\right),:\right)=\mathbf{I}_{k}$
 […]
-As an intuitive but effective remedy, the oversampled sketchy ID (OSID) in Algorithm 6 generalizes Algorithm 5 by increasing the sample size beyond $k$ (i.e.
-oversampling), which is shown to considerably alleviate such suboptimality of Equation 9 in interpolation error $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)$ (cf.
-Section 6.2, SkCPQR/SkLUPP-ID versus -OSID in Figures 2, 3, 4, 5, and 6).
-Concretely, with a moderate multiplicative oversampling $l=O(k)$ (e.g.
-$l=2k$ in Figures 2, 3, 4, 5, and 6 is usually sufficient in practice), Algorithm 6 can be expressed as
+As an effective remedy, the oversampled sketchy ID (OSID) in Algorithm 4 generalizes Algorithm 3 by increasing the sample size beyond $k$ (i.e.
+oversampling), which is shown to considerably alleviate such suboptimality of (7) in interpolation error $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)$ (cf.
+Section 6.2, SkLUPP-ID v.s.
+SkLUPP-OSID in Figures 2, 3, 4, 5, 6, and 7).
+Concretely, with a moderate multiplicative oversampling $l=\Theta(k)>k$ (e.g.
+$l=3k$ is usually sufficient in practice, see Section 6.2), Algorithm 4 can be expressed as
 (8) $\text{OSID}:\qquad\begin{split}&\underset{k\times l}{\mathbf{Y}_{S}}=\mathbf{Y}\left(S,:\right),\qquad\underset{l\times k}{\mathbf{Y}_{S}^{\top}}=\underset{l\times k}{\mathbf{Q}}\ \underset{k\times k}{\mathbf{R}}\\ &\mathbf{W}=\mathbf{Y}\mathbf{Y}_{S}^{\dagger}=\left(\mathbf{Y}\mathbf{Q}\right)\mathbf{R}^{-\top}=\mathbf{X}\left(\bm{\Omega}\bm{\Omega}^{\dagger}\right)\mathbf{X}_{S}^{\dagger}\end{split}$
 where $\mathbb{E}_{\bm{\Omega}}\left[\mathbf{W}\right]=\mathbf{X}\mathbf{X}_{S}^{\dagger}$ is an unbiased estimate for the optimal interpolation matrix when the embedding $\bm{\Omega}$ is isotropic, i.e., $\mathbb{E}_{\bm{\Omega}}\left[\bm{\Omega}\bm{\Omega}^{\dagger}\right]=\mathbf{I}_{d}$.
-Leveraging the existing theories for randomized linear embedding, [49, Theorem 9 and 23] implies that with Equation 10, the gap between the interpolation and skeletonization errors converges as $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)-\mathcal{E}_{\mathbf{X}}\left(S\right)=O\left({k^{2}}/{l}\right)$ when $\bm{\Omega}$ is a sparse embedding and $l\gg k^{2}$, whereas such convergence can be further improved to $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)-\mathcal{E}_{\mathbf{X}}\left(S\right)=O\left({k}/{l}\right)$ with Gaussian embeddings (e.g.
-by [49, Theorem 6]).
-Asymptotically, Algorithm 6 takes $O(nlk+nk^{2}+lk^{2})=O(nk^{2}+k^{3})$ time.
-Compared to Algorithm 5 without oversampling, the dominant cost $O(nk^{2})$ increases proportionally to the oversampling factor $l/k$.
+Algorithm 4 takes $O(nlk+nk^{2}+lk^{2})=O(nk^{2})$ time.
+Compared to Algorithm 3 without oversampling, the dominant cost $O(nlk)$ increases proportionally to the oversampling factor $l/k$.
 Nevertheless, such additional cost is affordable in practice (cf.
-Figures 2, 3, 4, 5, and 6 (right)) thanks to the parallelizability of matrix-matrix multiplications [23].
+Figures 2, 3, 4, 5, 6, and 7) thanks to the hardware efficiency of matrix-matrix multiplications [34].
+While sketchy pivoting without oversampling in (7) is not ID-revealing, the existing theories for sketch-and-solve (see e.g.
+[64, §2.5]) implies that OSID in (8) with Gaussian embedding and a multiplicative oversampling $l=\Theta(k)$ is $\gamma$-ID-revealing (recall Definition 1.1) for $\gamma=(1+O(\sqrt{k/l}))^{2}$ with a constant probability:
+Proposition 5.4 (Sketchy pivoting + OSID is $\gamma$-ID-revealing).
+Selecting a skeleton subset $S$ of size $k$ via sketchy pivoting [23, Algorithm 1] outputs (i) a sketched sample matrix $\mathbf{Y}=\mathbf{X}\bm{\Omega}\in\mathbb{R}^{n\times l}$ constructed by a Gaussian embedding $\bm{\Omega}\in\mathbb{R}^{d\times l}$ with $l>k$ and i.i.d.
+entries from $\mathcal{N}(0,1/l)$, along with (ii) a permutation $\bm{\pi}\in\mathfrak{S}_{n}$ that characterizes the skeleton subset $S=\bm{\pi}\left(1:k\right)$.
+For any $\delta\in(0,1)$, when $l\gtrsim k+\log(1/\delta)$, with probability at least $1-\delta$, Algorithm 4 constructs an interpolation matrix $\mathbf{W}$ that satisfies $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)\leqslant(1+O(\sqrt{\delta^{-1}k/l}))^{2}\mathcal{E}_{\mathbf{X}}\left(S\right)$ in $O(nkl)$ time.
+Proof 5.5 (Proof of Proposition 5.4).
+The proof of Proposition 5.4 is a direct generalization of [64, Theorem 23].
+The key fact is that for a $(1/2,\delta/2,k)$-JLT [64, Definition 3] $\bm{\Omega}\in\mathbb{R}^{d\times l}$ with $(\Theta(\epsilon/\sqrt{k}),\delta/2)$-JL second moment property [64, Definition 12] for some $\epsilon>0$, (8) provides an interpolation matrix $\mathbf{W}$ that satisfies $\|\mathbf{W}\mathbf{X}_{S}-\mathbf{X}\mathbf{X}_{S}^{\dagger}\mathbf{X}_{S}\|_{F}\leqslant\epsilon\|\mathbf{X}-\mathbf{X}\mathbf{X}_{S}^{\dagger}\mathbf{X}_{S}\|_{F}$ (see e.g.
+the proofs of [64, Theorem 23] or [25, Lemma C.5]), and therefore, $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)\leqslant(1+\epsilon)^{2}\mathcal{E}_{\mathbf{X}}\left(S\right)$.
+Then, it is sufficient to recall that a Gaussian embedding $\bm{\Omega}\in\mathbb{R}^{d\times l}$ with $l\gtrsim k+\log(1/\delta)$ is a $(1/2,\delta/2,k)$-Johnson-Lindenstrauss transform [64, Theorem 6] and satisfies the $(\Theta(\epsilon/\sqrt{k}),\delta/2)$-JL second moment property when $l\gtrsim\delta^{-1}k/\epsilon^{2}$.
+Notice that Proposition 5.4 can be easily extended to various structured randomized linear embeddings like the sparse embedding in [11] and the LESS embedding in [18, 9] that offer better efficiency for constructing the sample matrix $\mathbf{Y}$.
 6 Experiments
+The experiments in Figures 1, 2, 3, 4, 5, 6, and 7 are conducted on a 14-core Apple M3 processor, while experiments in Table 3 are conducted on an Nvidia V100 GPU.
 6.1 Data Matrices
-In the experiments, we compare the accuracy and efficiency of various ID algorithms on several types of data matrices, including the matrix forms of natural image datasets and synthetic random matrices with varied spectra, outlined as follows.
+We compare accuracy and efficiency of various ID algorithms on different types of data matrices, including natural data matrices from image datasets and integral operators, as well as synthetic matrices with varied spectra, outlined as follows.
   1. 1.
-    We draw a synthetic random data matrix $\mathbf{X}\in\mathbb{R}^{2000\times 500}$ from the adversarial GMM described in Example 4.1 with $k=100$ cluster centers, denoted as GMM.
+    We construct a synthetic data matrix $\mathbf{X}\in\mathbb{R}^{2000\times 500}$ based on the adversarial GMM example described in Example 4.1 with $k=100$ cluster centers, denoted as GMM.
   2. 2.
@@
     Let Gaussian-exp be a random dense data matrix of size $1000\times 1000$ with exponential spectral decay.
-We construct Gaussian-exp from its SVD $\mathbf{X}=\mathbf{U}\bm{\Sigma}\mathbf{V}^{\top}$ where $\mathbf{U},\mathbf{V}\in\mathbb{R}^{1000\times 1000}$ are random unitary matrices drawn from the Haar measure, and $\bm{\Sigma}=\mathop{\mathrm{diag}}\left(\sigma_{1},\cdots,\sigma_{1000}\right)$ where $\sigma_{i}=1$ for all $i\leq 100$ and $\sigma_{i}=\max\left\{0.8^{i-100},10^{-5}\right\}$ for all $i>100$.
+We construct Gaussian-exp from its SVD $\mathbf{X}=\mathbf{U}\bm{\Sigma}\mathbf{V}^{\top}$ where $\mathbf{U},\mathbf{V}\in\mathbb{R}^{1000\times 1000}$ are random unitary matrices drawn from the Haar measure, and $\bm{\Sigma}=\mathop{\mathrm{diag}}\left(\sigma_{1},\cdots,\sigma_{1000}\right)$ where $\sigma_{i}=1$ for all $i\leqslant 100$ and $\sigma_{i}=\max\left\{0.8^{i-100},10^{-5}\right\}$ for all $i>100$.
   5. 5.
-    Let SNN be a random sparse non-negative (SNN) matrix [44] of size $1000\times 1000$ such that
+    SNN is a sparse non-negative (SNN) matrix [60] of size $1000\times 1000$ generated by
     (9) $\mathbf{X}=\sum_{i=1}^{100}\frac{10}{i}\mathbf{u}_{i}\mathbf{v}_{i}^{T}+\sum_{i=101}^{1000}\frac{1}{i}\mathbf{u}_{i}\mathbf{v}_{i}^{T}$
     where $\mathbf{U}=\left[\mathbf{u}_{1},\cdots,\mathbf{u}_{1000}\right]$ and $\mathbf{V}=\left[\mathbf{v}_{1},\cdots,\mathbf{v}_{1000}\right]$ are $1000\times 1000$ random sparse matrices with non-negative entries and sparsity $0.1$.
+  6. 6.
+    Let Helmholtz be a matrix $\mathbf{X}$ induced by the Helmholtz kernel:
+    $G(\mathbf{x},\mathbf{y})=\exp\left(i\kappa\left\|\mathbf{x}-\mathbf{y}\right\|_{2}\right)/\left(4\pi\left\|\mathbf{x}-\mathbf{y}\right\|_{2}\right)\quad\text{with}\quad\kappa=5.5.$
+    Consider a set of source points ${\mathcal{S}}$ consisting of $15\times 15\times 15$ Clenshaw-Curtis quadrature nodes in a 3D cube of size $2\times 2\times 2$ centered at the origin, along with a target set ${\mathcal{T}}$ containing $2000$ uniformly distributed points on the surface of a 3D sphere with radius $3$, also centered at the origin.
+We construct $\mathbf{X}$ of size $3375\times 2000$ by evaluating the Helmholtz kernel on ${\mathcal{S}}\times{\mathcal{T}}$ such that $\mathbf{X}_{ij}=G(\mathbf{x}_{i},\mathbf{y}_{j})$ for every $\mathbf{x}_{i}\in{\mathcal{S}}$ and $\mathbf{y}_{j}\in{\mathcal{T}}$.
+It is worth highlighting that ID on natural data like MNIST, CIFAR-10, and Helmholtz is closely related to practical applications like data selection and fast direct solvers.
+Remark 6.1 (ID for data selection).
+ID on a training dataset like MNIST and CIFAR-10 can be viewed as an idealized data selection problem [1, 25].
+For example, consider data selection for an overparametrized ridge regression problem without noise: given a full dataset $(\mathbf{X},\mathbf{y})\in\mathbb{R}^{n\times d}\times\mathbb{R}^{n}$ satisfying $\mathbf{y}=\mathbf{X}\bm{\theta}_{*}$ for some unknown ground truth $\bm{\theta}_{*}\in\mathbb{S}^{d-1}$, we aim to approximate $\bm{\theta}_{*}$ via $\widehat{\bm{\theta}}=\mathop{\mathrm{argmin}}_{\bm{\theta}\in\mathbb{R}^{d}}\|\mathbf{X}_{S}\bm{\theta}-\mathbf{y}_{S}\|_{2}^{2}+\alpha\|\bm{\theta}\|_{2}^{2}$ with some $\alpha>0$ on a selected data subset $(\mathbf{X}_{S},\mathbf{y}_{S})\in\mathbb{R}^{k\times d}\times\mathbb{R}^{k}$ of size $k=\left|S\right|\leqslant d$.
+In this setting, the skeletonization error of ID, $\mathcal{E}_{\mathbf{X}}\left(S\right)$, is closely related to the generalization error of data selection:
+$\text{as}\ \alpha\to 0,\quad\|\mathbf{X}(\widehat{\bm{\theta}}-\bm{\theta}_{*})\|_{2}^{2}=\|(\mathbf{X}\mathbf{X}_{S}^{\dagger}\mathbf{X}_{S}-\mathbf{X})\bm{\theta}_{*}\|_{2}^{2}\leqslant\mathcal{E}_{\mathbf{X}}\left(S\right).$
+Beyond linear models like ridge regression, this connection can be extended to the more general finetuning problems for nonlinear models (like neural networks) by replacing $\mathbf{X}$ with the high-dimensional representations or model gradients evaluated at the data [25].
+Remark 6.2 (ID for fast direct solvers).
+In fast direct solvers for linear elliptic PDEs, ID is a widely used tool for compressing certain off-diagonal blocks of a rank-structured matrix [51, 32, 39, 40, 13, 53].
+In this setting, both a high-quality skeleton selection and an accurate interpolation matrix construction are crucial for attaining a nearly minimum compression error.
+The ID experiments on the Helmholtz matrix are designed to model a compression step in the state-of-the-art fast direct solvers for Lippmann-Schwinger equation.
 6.2 Accuracy and Efficiency of ID
-In Figures 2, 3, 4, 5, and 6, we compare the empirical accuracy (left) and efficiency (right) of different ID algorithms on data matrices described in Section 6.1, with a focus on those demonstrating robust skeleton complexities on the adversarial input Example 4.1 (i.e.
-algorithms in Figure 1 (right)).
-Figure 2: Relative interpolation error and runtime of ID algorithms with robust skeleton complexities on the GMM adversarial input (Example 4.1).
-Recall from Figure 1 that sampling methods, especially squared-norm sampling, suffer from significantly higher skeleton complexities than sequential/robust blockwise random pivoting and sketchy pivoting on GMM.
-For demonstration, we omit squared-norm sampling from the plot.
-Taking an integrated view of ID including skeleton selection and interpolation matrix construction, we consider the following relative ID error777 All error measurements are normalized with $\left\|\mathbf{X}\right\|_{F}^{2}$ to relative errors.
+In Figures 2, 3, 4, 5, 6, and 7, we compare accuracy (left) and efficiency (right) of different ID algorithms on data matrices in Section 6.1.
+Figure 2: Relative interpolation error and total runtime of ID algorithms on the GMM adversarial input (Example 4.1).
+Recall from Figure 1 that in lack of adaptiveness, methods like SqNorm and BRP tend to suffer from suboptimal skeleton complexities on GMM.
+Taking an integrated view of ID including skeleton selection and interpolation matrix construction, we consider the following relative ID error666 All error measurements are normalized with $\left\|\mathbf{X}\right\|_{F}^{2}$ to relative errors.
 and runtime measurements:
   1. (a)
-    For error-reveling algorithms (i.e.
-except for sketchy pivoting and sampling)888 In Figures 2, 3, 4, 5, and 6, we also omit the residual errors of the sequential error-reveling algorithms SRP and CPQR (which overlap with the corresponding interpolation errors) for conciseness of demonstration.
-, we plot the residual errors (denoted as residual) from their on-the-fly skeletonization error estimates and the runtimes for skeleton selection (e.g.
-$\left\|\mathbf{d}^{(t)}\right\|_{1}$ in Algorithm 4) only.
+    For ID-revealing algorithms (i.e.
+except for sketchy pivoting and sampling), we plot the runtimes and interpolation errors $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)$ with $\mathbf{W}$ from Algorithm 3 (ID), which are equal to the corresponding skeletonization errors $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)=\mathcal{E}_{\mathbf{X}}\left(S\right)$.
   2. (b)
-    For exactly ID-revealing algorithms (i.e.
-except for sketchy pivoting and sampling), we plot the runtimes and interpolation errors $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)$ with $\mathbf{W}$ from Algorithm 5 (ID), which are equal to the corresponding skeletonization errors $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)=\mathcal{E}_{\mathbf{X}}\left(S\right)$.
+    For sketchy pivoting, we plot the runtimes and interpolation errors $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)$ with $\mathbf{W}$ from both Algorithm 3 (ID) and Algorithm 4 (OSID) with small multiplicative oversampling $l=3k$.
+Notice that both ID and OSID have $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)>\mathcal{E}_{\mathbf{X}}\left(S\right)$, while OSID in Algorithm 4 tends to be more expensive because of oversampling.
   3. (c)
-    For approximately ID-revealing algorithms (i.e.
-sketchy pivoting), we plot the runtimes and interpolation errors $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)$ with $\mathbf{W}$ from both Algorithm 5 (ID) and Algorithm 6 (OSID) with small multiplicative oversampling $l=2k$, where both $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)>\mathcal{E}_{\mathbf{X}}\left(S\right)$, and Algorithm 6 tends to be more expensive.
-  4. (d)
-    For non-ID-revealing algorithms (i.e.
-sampling), without explicitly specifying in the legends, we plot the runtimes and interpolation errors $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)=\mathcal{E}_{\mathbf{X}}\left(S\right)$ with the optimal $\mathbf{W}$ from solving the exact least square problem in Equation 8.
-The legends in Figures 2, 3, 4, 5, and 6 are formatted in terms of “skeleton selection algorithm”-“ID error”.
-The abbreviations of different skeleton selection algorithms are summarized in Table 2, whereas the ID error measurements are described as above.
+    For squared-norm sampling, without explicitly specifying in the legends, we plot the runtimes and interpolation errors $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)=\mathcal{E}_{\mathbf{X}}\left(S\right)$ with the optimal $\mathbf{W}$ from solving the exact least-square problem in (6).
+The legends in Figures 2, 3, 4, 5, 6, and 7 are formatted in terms of “skeleton selection algorithm”-“ID error”.
+The abbreviations of different skeleton selection algorithms are summarized in Table 2, whereas the ID error measurements are described above.
 Figure 3: Relative interpolation error and runtime of ID algorithms on MNIST.
 Figure 4: Relative interpolation error and runtime of ID algorithms on CIFAR-10.
-Remark 6.1 (Choice of block size).
-The choice of block size $b$ controls the trade-off between parallelizability and error-revealing property.
-To be precise, an unsuitably small $b$ lessens the efficiency gained from moving flops to matrix-matrix multiplications, whereas the larger $b$ leads to higher overestimation of the target rank.
-In the experiments, we use $b=30$ or $40$ to exploit the parallelizability of matrix-matrix multiplications without considerably compromising the error-revealing property.
 Figure 5: Relative interpolation error and runtime of ID algorithms on Gaussian-exp.
 Figure 6: Relative interpolation error and runtime of ID algorithms on SNN.
+Figure 7: Relative interpolation error and runtime of ID algorithms on Helmholtz.
+Remark 6.3 (Choice of block size).
+The choice of block size $b$ controls the trade-off between hardware efficiency and rank-adaptiveness.
+To be precise, an unsuitably small $b$ lessens the efficiency gained from moving flops to matrix-matrix multiplications, whereas the larger $b$ leads to a higher overestimation of the target rank.
+In the experiments, we use $b=30$ or $40$ to exploit the hardware efficiency of matrix-matrix multiplications without considerably compromising the rank-adaptiveness.
 For the adversarial GMM dataset in Figure 2, it is worth mentioning that the sharp transition of RBGP at $\left|S\right|\approx 100$ is because GMM contains $100$ clusters, each with distinct squared-norms.
 Recall that blockwise greedy pivoting picks points with the maximum norms without randomness.
-When $\left|S\right|\leq 100$, each block drains points from the same clusters with the maximum residual norms first, most of which are redundant points and are later excluded by the robust blockwise filtering.
-This leads to the inefficiency of RBGP when $\left|S\right|\leq 100$.
+When $\left|S\right|\leqslant 100$, each block drains points from the same clusters with the maximum residual norms first, most of which are redundant points and are later excluded by the robust blockwise filtering.
+This leads to the inefficiency of RBGP when $\left|S\right|\leqslant 100$.
 Such inefficiency of RBGP in comparison to RBRP again illustrates the superiority of random pivoting over greedy pivoting, i.e.
 the merit of randomness.
-Table 3: GPU runtime of the relatively efficient/accurate ID algorithms in Figure 2 on the synthetic adversarial GMM dataset (Example 4.1) of size $10^{5}\times 10^{3}$, with $k=100$ cluster centers.
+Table 3: GPU runtime of the relatively efficient/accurate ID algorithms in Figure 2 on an adversarial GMM matrix (Example 4.1) of size $10^{5}\times 10^{3}$, with $k=100$ cluster centers.
 Experiments were performed on an Nvidia V100 GPU.
@@
 Moreover, as the rank increases, the efficiency of RBRP-ID approaches that of SkCPQR/SkLUPP-OSID, in addition to better accuracy.
-From experiments on the (non-adversarial) natural and synthetic datasets in Figures 3, 4, 5, and 6, we observe the following (by algorithms):
-  • The robust blockwise random pivoting (RBRP) enjoys (one of) the best skeleton complexities, with the skeletonization error estimates aligning perfectly with the interpolation errors (left), demonstrating the error-revealing and exactly ID-revealing abilities of RBRP.
-The skeleton selection process of RBRP tends to take slightly longer than that of sketchy pivoting (SkLUPP/SkCPQR) with ID (which is not error-revealing and suffers from much larger interpolation errors than RBRP).
-However, RBRP is generally faster than sketchy pivoting (SkLUPP/SkCPQR) with small multiplicative oversampling (OSID), while enjoying lower interpolation error.
+From the (non-adversarial) datasets in Figures 3, 4, 5, 6, and 7, we observe the following:
+  • RBRP enjoys (one of) the best skeleton complexities.
+In terms of efficiency, RBRP tends to be slower than sketchy pivoting (SkLUPP) with ID (which is neither rank-adaptive nor ID-revealing).
+However, RBRP is usually faster than sketchy pivoting with OSID, while enjoying lower interpolation error.
 Overall, RBRP is shown to provide the best combination of accuracy and efficiency when the explicit construction of ID (in contrast to the skeleton subset only) is inquired.
-  • The sketchy pivoting algorithms (SkLUPP/SkCPQR) also enjoy (one of) the best skeleton complexities, despite the lack of error-revealing ability and being approximately ID-revealing (i.e.
+  • Sketchy pivoting (SkLUPP) also enjoy (one of) the best skeleton complexities, despite the lack of rank-adaptiveness and ID-revealing property (i.e.
 suffering from large interpolation errors).
@@
 Figure 1).
-Moreover, our experiments verify the conclusion in [16] that SkLUPP tends to be more efficient than SkCPQR in selecting skeletons without compromising the skeletonization error.
-  • The inherently sequential algorithms, including SRP and CPQR, are highly competitive in terms of accuracy while being considerably slower than the parallelizable ones.
+  • The sequential algorithms SRP and CPQR are highly competitive in terms of accuracy while being considerably slower than the hardware-efficient blockwise algorithms.
 […]
+  • When using SqNorm as an ID algorithm, constructing the interpolation matrix by solving (6) explicitly can slow down the overall ID algorithm significantly to the extent that the runtimes are similar to those of RBRP.
+Moreover, sampling methods like SqNorm and DPP lack rank-adaptiveness and ID-revealing ability.
 7 Discussion
-In this work, we focus on fast and accurate ID algorithms from five perspectives that measure the empirical performance systematically: (a) skeleton complexity, (b) asymptotic complexity, (c) parallelizability, (d) error-revealing property, and (e) ID-revealing property.
+In this work, we focus on fast and accurate ID algorithms from five perspectives that measure the empirical performance systematically: (a) skeleton complexity, (b) asymptotic complexity, (c) hardware efficiency, (d) rank-adaptiveness, and (e) ID-revealing property.
 With a careful exploration of various existing ID algorithms through the lens of adaptiveness and/or randomness, we reemphasize the effectiveness of combining adaptiveness and randomness in solving ID problems, while unveiling a critical missing piece in the family of existing algorithms that achieves all five of the aforementioned desired properties.
-To close the gap, we propose robust blockwise random pivoting (RBRP)—a parallelizable, error-revealing, and exactly ID-revealing algorithm that enjoys among-the-best skeleton and asymptotic complexities in practice compared to the existing algorithms and demonstrates promising robustness to adversarial inputs.
+To close the gap, we propose robust blockwise random pivoting (RBRP)—a hardware-efficient, rank-adaptive, and ID-revealing algorithm that enjoys among-the-best skeleton and asymptotic complexities in practice compared to the existing algorithms and demonstrates promising robustness to adversarial inputs.
 Through numerical experiments, we illustrate the empirical accuracy, efficiency, and robustness of RBRP on a broad spectrum of natural and synthetic datasets.
 Acknowledgments
-The authors wish to thank Yifan Chen, Ethan Epperly, Kevin Miller, Chris-topher Musco, Rachel Ward, and Robert Webber for enlightening discussions.
+The authors wish to thank Yifan Chen, Ethan Epperly, Kevin Miller, Christopher Musco, Rachel Ward, and Robert Webber, as well as the anonymous referees and editors, for enlightening discussions and constructive feedback.